### v4

# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs

Abstract
LLM-as-a-Judge presents a promising alternative to human evaluators across various tasks, but inherent biases, especially position bias — a tendency to favor solutions based on their position in the prompt — have compromised its effectiveness. Our study introduces a systematic framework to examine position bias in pairwise comparisons, focusing on repetition stability, position consistency, and preference fairness. This research significantly contributes to the field by introducing new concepts for understanding position bias and providing a multi-dimensional framework for evaluations. We conducted experiments with 12 LLM judges across MTBench and DevBench, covering 22 tasks and approximately 40 solution-generating models — candidates, resulting in over 100,000 evaluation instances. Our findings confirm that position bias in capable LLM judges is not due to random chances, along with notable variations observed across judges and tasks. Moreover, position bias is weakly influenced by the length of prompt components but significantly impacted by the quality gap between solutions. These insights can help optimize judge model selections, improve benchmark design, and inform future research on debiasing strategies, ultimately enhancing the reliability of LLM judges.
1 Introduction
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations (Zeng et al., 2023; Karpinska et al., 2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments(Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1. This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations (Qin et al., 2024), for any score-based and relation-based settings (Li et al., 2023d). Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias. However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks (Khan et al., 2024; Chua et al., 2024). Moreover, the limited understanding of the bias itself has exacerbated the problem.
Prior studies on position bias implicitly assume that this bias is not due to random variations. In other words, these studies assume that LLM judges produce consistent results across repeated trials. However, without validating this assumption, evaluations may become less reliable. As shown in Fig. 1, the occurrence of position bias can be observed by chance even for the same set of tasks and candidate models. Thus, it is unclear whether the observed position bias actually arises from the position of the prompt components or is merely due to random chance. Therefore, when exploring the position bias of LLM judges, it is necessary for the judges to be stable across repetitions to yield nontrivial results.
Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge. When the positions of candidate model solutions are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same candidate model solution. We observe repetition instability if the same LLM judge gives different judgments across repetitions. Only when an LLM judge has high repetition stability can we validate that its position bias is not due to random variations.
Another key aspect of position bias, preference fairness, has also been under-explored by previous works. Most studies have primarily focused on position consistency to measure how frequently position bias occurs across a set of evaluation instances. However, the direction of biased preference (i.e., primacy or recency) is also insightful. Fig. 2 illustrates cases where a pair of judgments with swapped prompts become position consistent, inconsistent, primacy-preferred, or recency-preferred. We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations. Specifically, judgments are considered to have a fair preference if they are position consistent or if the preference directions are evenly distributed. In contrast, an LLM judge is considered to have an unfair preference if it consistently favors candidate model solution that appear either first or last in the prompt. Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
Figure 2: Demonstration of position consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better candidate solution. A judgment is consistent if it selects the same winner across different prompt positions. Primacy/recency indicates a preference for the first/last candidate solution, respectively. Position inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment. We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring (Zheng et al., 2024b; Liusie et al., 2024), making it a practically preferable choice for investigations.
Our contribution can be summarized as following: (1) introducing a systematic framework for evaluating LLM-as-a-Judge’s position bias in terms of Repetition Stability, Position Consistency, and Preference Fairness; (2) identifying Judge-level, Candidate-level, and Task-level factors that affect position bias; and (3) providing insights into future benchmark design and paving the way for effective mitigation strategies for position bias.
2 Methods & Defintions
LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments. Our study focuses on relation-based assessment, specifically under the pairwise comparison setting, for several reasons. First, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields (Qin et al., 2024). Second, as noted by Liu et al. (2024), LLM and human evaluations are more aligned in the context of pairwise comparison than any of the score-based assessments. Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of position consistency(Zheng et al., 2024b; Liusie et al., 2024). Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, through some advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024). Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
In the context of pairwise comparative assessments, LLM judges are prompted to choose the better candidate solution to the task question. To accommodate the possibility of a tie, various option modes are introduced. Two-Option mode restricts judges to choosing the better solution between two options, labeled {A} for the first and {B} for the second as shown in Fig. 1. Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither solution is preferable. Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie. In our experiments, we employed Three-Option mode for MTBench (Zheng et al., 2024b) and Two-Option mode for DevBench (Li et al., 2024), aligning with their respective original frameworks. These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM judges.
To set up our experiment, as shown in Fig. 1, the system prompt, option mode, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the winning solution. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question. Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair. As shown in Fig. 2, if the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent with a fair preference. Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
To systematically investigate position bias in pairwise comparative judging scenarios, we evaluate LLM judges using three key metrics: Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$). While $PC$ has been widely adopted in prior research, we introduce $RS$ to validate the reliability of LLM judges and $PF$ to provide a more comprehensive assessment. Further analysis examines the impact of judge-level, candidate-level, and task-level factors on position bias. Therefore, for each judge, candidate, and task, we calculate $RS$, $PC$, and $PF$ to thoroughly explore the variance in position bias and its impacting factors.
2.1 Evaluation Metrics
We validate the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows.
Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times. It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations. We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset. This metric is formalized as
$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,...\}$ (1)
where $S$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries. The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is swapped. It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the judge model prefers the same winning solution. Formally, it is calculated as
$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}},V=\{(A,B),(B,A),(C,C),...\}$ (2)
where $(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})$ denotes the judgment pair for the $j^{th}$ query, $n$ represents the number of prompt pairs, and $V$ is the set of choice pairs that correspond to position consistency. An example of such pairs of choices can be found in Fig. 2. This formula provides a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain solution positions. In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second position” used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 2.
Position consistent judgments are fair in preference. The measurements become more complex for position inconsistent cases. Previous studies proposed two common ways to measure the preference fairness. A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023b). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach is problematic because it overlooks the fact that “position consistent judgments are also position fair,” which can lead to overly penalizing highly consistent LLM-judges. For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred. However, this outcome should be distinguished from cases where all one hundred pairs are inconsistent and biased toward one position, which would also result in 100% preference unfairness.
To overcome these limitations, we introduce a more granular and scalable measurement that combines the strengths of both methods, to assess preference fairness. The $PF$ score is formally calculated by
$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$ (3)
where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively. This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero. Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the position consistent judgments. Another advantage of this approach is that it quantifies preference fairness with a single score, rather than the two separate scores used in previous studies. The $PF$ score is interpreted as follows:
$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
To summarize, our proposed preference fairness score is superior to prior measurements due to its ability to provide a single, comprehensive score that accounts for all evaluation instances, while maintaining its sensitivity across datasets, even when the number of questions and instances varies for each task.
2.2 Factors affecting position bias
Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose solutions to the task-specific questions are evaluated by the judges. Factors that have a significant impact on position bias are highlighted with an asterisk (*) by bidirectional stepwise regression tests and marked red by empirical findings. ‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s solutions to the question, making it both a Candidate-level and Task-level factor. ‘Prompt’ encompasses the complete query presented to the judges, including Task Input, Task Output, and system prompts.
Factor Judge-level Candidate-level Task-level
Familial Property $\checkmark$* $\times$ $\times$
Answer Quality Gap $\times$ $\checkmark$* $\times$
Task Input Length $\times$ $\times$ $\checkmark$
Task Output Length $\times$ $\checkmark$* $\checkmark$*
Prompt Length $\times$ $\checkmark$ $\checkmark$
To investigate the factors influencing position bias in LLM judges, we categorized these factors into three groups: Judge-level, Candidate-level, and Task-level factors. Each group includes specific factors, that we hypothesize, may impact position bias, which we explore through a series of experiments. Table 1 lists the five factors analyzed in this study. By design, our framework allows for the integration and assessment of additional influencing factors.
Among the influencing factors we identified, we chose “familial property” for Judge-level factors instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments. The model family grouping reflects similar model sizes and training specifics of the model members. The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version. However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable” models sharing similar judging capabilities. More details and discussions about the familial property can be found in Appendix Sec. C, where we examined the LLM mutual agreements to study the similarities and differences across families.
Answer quality gap, based on our findings, is highly impactful yet under-explored in prior research. We define the quality of a candidate’s answer/solution by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between the solutions from two candidate models to the same question and is considered the Candidate-level (i.e., solution provider) factor. Ideally, when a reliable LLM judge is presented with a question and corresponding solution pairs, it would prefer the higher-quality solution, where the corresponding candidate is denoted as the winner selected by the judge. Following this assumption, we measure the answer quality gap by the win rates between two candidates on a set of tasks and questions. We regard a tie judgment as ‘half-win’ for both candidates when calculating its win rate over another, depending on option mode. If position bias occurs, the winner may be inconsistent when the order of candidate solutions is swapped in the query. Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen in both the original and swapped queries (termed “consistent wins”), cases where both responses are deemed equal (termed “consistent ties”), and cases where different winners are selected after the solutions are swapped. We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment pairs ($C_{I}$), respectively. Following Zheng et al. (2024b), we count inconsistent judgment pairs also as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
Inspired by Zheng et al. (2024b), we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs. Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, where 0.5 is the overall win rate when we only have ties. In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook. More discussions can be found in Appendix. Sec. D.
3 Experiments
3.1 Experiment Settings
In this study, we evaluated twelve up-to-date commercial models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework. Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks. However, due to their limited performance on our tasks, as shown in Table 2, we focused our investigation solely on more powerful closed-source models. We included the results of open-source models for completeness.
We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks. We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges. The prompt templates we used are identical to those used in the benchmarks. More details about the models, tasks, and prompts can be found in Appendix. Sec. H.
To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials. This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench. The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results. To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets. In total, approximately 100,000 evaluation instances were analyzed in this study.
To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks. We used variables such as average lengths of input, output, and prompt; answer quality gap; LLM judge series; candidate identities; and task categories to predict $PC$ and $PF$, respectively. Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score. This process involves both forward selection and backward elimination, with each ”step” testing whether including or excluding a variable improves the model’s AIC value. Further details about the process can be found in Appendix. Sec. F.
3.2 Empirical Results and Interpretations
The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not merely random variations. Fig. 3 and Fig. 7 demonstrate that position bias varies by judges and tasks significantly. Fig. 4 further investigates the impact of answer quality gap and lengths of prompt components on position bias. These analyses were conducted by considering all judges together on MTBench. The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E, and Sec. G.
Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables. Similarly, these factors also contribute significantly to Preference Fairness. Additionally, we found that average output length is a statistically significant predictor of $PF$. This finding is not surprising, as longer outputs are generally perceived as higher quality and more preferred. Quantitative results and more discussions can be found in Appendix. Sec. F.
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Positional Fairness ($PF$). Top 5 performances are marked in bold. “Error” is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format. High error rates and low $RS$ are marked red, implying further evaluations (i.e., $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data. Therefore, open-source models (e.g., Llama) are not investigated in the paper because of their high error rates under the difficulty and length of our evaluation instances. In contrast, up-to-date commercial models generally demonstrated minimal error and high repetition stability, making them more practical and meaningful for investigations. Chatbot Arean Leaderboard’s overall ranking of the models is listed.
Judge MTBench DevBench Arena
$RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error
Caude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 7
Claude-3-Opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00 17
Claude-3-Sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00 37
Claude-3-Haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00 46
Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 12
Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96 25
Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00 66
o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00 2
GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 6
GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00 17
GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00 52
GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 94
Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 99
Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 109
Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94 57
Figure 3: Baseline comparisons of judges across tasks on MTBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be GPT-4 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
Figure 4: Impact of answer quality gap and lengths on position bias. All figures are based on the integration results of all judge models on MTBench. Figures (a) and (b) investigate $PC$ and $PF$ vs. overall win rate, hence the answer quality gap, respectively. Generally, judgments become more consistent and fair as the quality gap decreases. Figure (c) demonstrates the relationship between $PC$ and $PF$, where higher consistency is likely to indicate better fairness. Figure (d)(e)(f) plots the relationship between lengths of prompt components and $PC$. The irregular and inconclusive patterns imply that position bias is not significantly impacted by the length of prompts and their components.
4 Main Findings
Position Bias of Capable Judges are not Mere Random Variations As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal ”Error” rates, generally exhibit $RS$ values above 0.85. The most up-to-date and capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks. These results confirm that LLM judgments, and the resulting position bias, are not merely random variations. This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
Position Bias Varies by Judge & Task As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig. 3 for $PC$ comparisons. However, certain models achieve comparable or superior performances than GPT-4 on certain tasks. For example, according to Fig. 3, for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges. Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it may be employed as a cost-effective alternative to coding evaluations.
Additionally, significant variations of $PF$ across tasks are observed. The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on the others. Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, role play, and writing tasks. Even for uniformly recency-preferred judges such as Claude-3’s and Gemini-pro’s, the extent of biased preference, as reflected by $PF$ values, varies by task.
Moreover, high position consistency does not guarantee fairness. For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively. In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
Therefore, the position bias of LLM judges is judge-dependent and task-dependent. This observation is confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$. In practice, to achieve the most reliable evaluation results, one needs to select task-specific LLMs whose consistency and fairness are in balance. Moreover, as shown in Table 2, a high ranking in the overall reasoning performance of a model does not guarantee its judging reliability. Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec. G.
Position Bias Correlates to Answer Quality Gap Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section 2.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$). Since an absolute tie has 0.5 $owr$, $\delta_{p}$ increases as $owr$ extends from 0.5 to 0 or 1. Fig. 4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answer pairs with larger quality disparities are easier to achieve consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec. E.
Similarly, as shown in Fig. 1(b), judgments generally become more preference fair as $\delta_{q}$ increases. However, the extent is not as significant as for $PC$. Also, the relationship varies by judge, as some LLMs maintain preference fairness regardless of $\delta_{q}$. For example, as shown in Fig. 6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
These observations align with the right-arrow shape as demonstrated in Fig. 4(c), where there is a general trend that judgments become preference fairer as position consistency increases. It also justifies the reasonableness of our quantification of preference fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
Together, we conclude that as the answer quality gap enlarges, judges generally become more position consistent and preference fair according to the regression curves. However, exceptions are common, as shown by the individual scatter points of these figures. This indicates that though the answer quality gap significantly influences the position bias of LLMs, other factors also play important roles. Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
Our investigation also helps to explain the observed verbosity/length bias (Liu et al., 2024; Kim et al., 2023; Zheng et al., 2024b) and self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b) of LLM judges. From our analysis, LLMs may prefer longer answers (verbosity/length bias) since more capable models usually generate longer solutions with higher qualities. Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the solutions generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4). In both cases, the observed biased preference may be fundamentally due to a prominent answer quality gap instead of length or self-identity.
Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length). Since the tokenization methods vary by models, we apply the length of the prompt component strings for uniformity and scalability. The analysis from Fig. 4(d)(e)(f) shows that there is very weak relationship between the lengths of prompt components and position bias. By stepwise regression, the average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table. 4. Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
Practical Implications for Benchmark Design We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appendix.Sec. C to gain further insights. In particular, the disagreement analysis helps identify instances where it is easy or difficult for judges to reach a consensus, reflecting the complexity of the judgment task. Based on our findings regarding the answer quality gap and LLM agreement/disagreement, this study offers practical implications for designing evaluator benchmarks that control the difficulty of judging tasks. Our findings show that the hardest-to-judge instances are those where (1) LLMs frequently disagree with each other, (2) comparable win rates and minimal quality gaps exist between the two candidate models, and (3) significant position bias is exhibited by the majority of judges.
One qualitative analysis example of a difficult-to-judge instance comes from MTBench question 120, which involved a math problem where both candidate models, Vicuna-13b-v1.3 and WizardLM-30b, provided incorrect solutions. The problem was: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find $x$ such that $f(x)=0$. Vicuna-13b-v1.3 incorrectly applied the quadratic formula to solve the cubic equation, while WizardLM-30b used an enumeration method but failed to identify the correct answer. This resulted in a 4{A}-3{B}-5{C} choice pattern among the twelve judges, illustrating a typical case of significant disagreement.
We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer. Others preferred either correctness (choosing {C} as a tie) or methodological soundness (choosing {B}, favoring WizardLM-30b). In the swapped order setting, 9 out of 12 judges displayed position bias. This highlights the challenge of judging solutions with minimal quality gaps, leading to greater position bias and subjectivity.
The key takeaway is that reducing hard-to-evaluate, highly subjective instances can improve benchmark design. Likewise, minimizing trivial cases where judges easily agree would enhance the benchmark’s robustness. Clearer and more specific evaluation criteria could also help resolve ambiguity in difficult instances.
5 Conclusion
In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of LLM judges in pairwise comparative scenarios in terms of repetition stability, position consistency, and preference fairness. Through comprehensive evaluations of 12 judges on two benchmarks across 22 tasks with over 100,000 evaluation instances, we observe significant variations of position bias across judges and tasks. We also discover that position bias is weakly length-dependent but significantly impacted by the answer quality gap. Our findings enhance the understanding of position bias, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
Appendix Table of Contents
[…]
[appendices]0
Appendix
Appendix A Contribution & Limitation
Contributions
Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge. Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through position consistency and preference fairness. We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components. Particularly, our study is the first to establish a direct connection between quantified answer quality disparities and position bias, highlighting the urgent need for a comprehensive understanding of how position bias affects LLM judges. Additionally, the mutual agreement and disagreement analysis of LLM judges are insightful for identifying the difficulty of evaluation instances and beneficial for benchmark assessment and designs. All findings from our study enhance the understanding of the position bias of LLM-as-a-Judge, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
Social Benefits Our comprehensive understanding of position bias in LLM-as-a-Judge provides multiple benefits to the community. First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation. Our focus and investigation on preference fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability. Moreover, our emphasized trade-offs between position consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work. Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results. Besides, the LLM agreement/disagreement analysis in Appendix.Sec.C 5 provides an applicable tool to identify the difficulty of judging certain instances. Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias. Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare Chen et al. (2024c), instruction following and prompting Zeng et al. (2023); Chua et al. (2024), multimodal assessment Chen et al. (2024a), and recommender systems Hou et al. (2024).
Strength
Our framework of systematically understanding position bias has the following strengths: (1) scalability and ease of use - our proposed framework is scalable to various judges, models, tasks, and judging types. It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works. (2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, position consistency, and preference fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias. (3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
Limitations
Due to computational limits, our study can be potentially extended to provide more comprehensive insights. (1) We only study the pairwise comparative assessment of LLM-as-a-Judge for its best consistency, ease of use, and scalability, but more types of judging can be explored, such as pairwise scoring and listwise ranking. (2) We only study 12 up-to-date commercial LLM judges for systematic investigations. Llama 7b-13b models that we choose as representatives of open-source models of similar sizes, face extreme difficulty in judging the instances explored in this study. Due to limited resources and considering practicality, we did not experiment with open-source models with larger sizes. However, further exploration may include analysis of these state-of-the-art open-source and fine-tuned judge models using our proposed framework. (3) We only study position bias on two benchmarks, MTBench and DevBench, which is still limited despite 22 tasks and 40 answer-generating models. Also, the number of questions for each task is limited. Further studies can be extended to more benchmarks to have a broader understanding of position bias across benchmarks, tasks, and questions. (4) Due to data accessibility, we are unable to investigate Judge-level factors in terms of model parameter sizes and architectures, which we expect will have a significant impact on position bias. Alternatively, we consider the family property to account for Judge-level factors by grouping models with similar properties (e.g., approximate parameter size) together. We also investigate variances of position bias across individual judges. However, future exploration on these model properties, especially those available on open-source models, will provide more insights.
Future Work
Besides addressing the limitations, the following future work, built on findings from our study, could provide more insights: (1)prompt setting - in this study, we apply the exact default prompt settings of MTBench and DevBench. However, the positional order and prompt style of not just the model-generated responses but also system prompt components (e.g., agent role assignment, mode of judging, direct mention of not making biases) may also have an impact on the extent to which LLM judges make position bias. (2) open-source insights: Further investigations of open-source and fine-tuned models on easier-to-evaluate benchmark may provide more insights into the training stages or model structures that are affecting or causing the position bias. Then our framework offers better control on the other impactful factors such as the answer quality gap to make these further investigations more accurate.
Appendix B Related Work
B.1 LLM-as-a-Judge
In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications. For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature (Zeng et al., 2023). However, it lacks scalability and reproducibility (Karpinska et al., 2021).
As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks (Zheng et al., 2024b; Li et al., 2024; Chiang & Lee, 2023b; Liusie et al., 2024; Chen et al., 2023; Fu et al., 2023; Wang et al., 2023a; Kocmi & Federmann, 2023b; Ji et al., 2023; Liu et al., 2024; Zeng et al., 2023; Liu et al., 2023; Shen et al., 2023; Kocmi & Federmann, 2023a; Chiang & Lee, 2023a; Karpinska et al., 2021; Dubois et al., 2024; Kim et al., 2023; Liang et al., 2023; Chang et al., 2023; Zhong et al., 2023; Chia et al., 2023; Deb & Das, 2023; Ye et al., 2024; Min et al., 2023; Li et al., 2023d; c; b; Lu et al., 2024), including open-ended story generation (Chiang & Lee, 2023a), adversarial attacks (Chiang & Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi & Federmann, 2023a), and instruction following (Zeng et al., 2023).
These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles (Zheng et al., 2024b; a; Pezeshkpour & Hruschka, 2023; Chiang & Lee, 2023b; Shen et al., 2023; Liusie et al., 2024; Qin et al., 2024; Chen et al., 2023; Huang et al., 2024; Zhu et al., 2023b; Raina et al., 2024; Chen et al., 2024b; Wang et al., 2023b; Liu et al., 2024; Zeng et al., 2023; Khan et al., 2024; Kim et al., 2023; Li et al., 2023b; d; Shi et al., 2024; Li et al., 2023c; Chen et al., 2024c; Chua et al., 2024; Chen et al., 2024a). As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation. Therefore, cost-effective LLM judges are needed to assess these responses efficiently. LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
In many of these tasks, LLM judges have shown a high level of agreement with human evaluators (Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), yet in some tasks, they are less effective, largely due to inherent biases (Shen et al., 2023; Chiang & Lee, 2023b; Zheng et al., 2024a). Even in cases where agreement is high, judgments may still suffer from biases.
When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method. From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment (Qin et al., 2024). By comparative method, it can be either score-based or relation-based (Li et al., 2023d).
For example, pointwise scoring (Kim et al., 2023) lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics. Pairwise/Listwise Scoring (Karpinska et al., 2021; Zhu et al., 2023b; Chen et al., 2024a; Zheng et al., 2024b) prompts the LLM judges to score a pair/list of model-generated answers. Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst). Pairwise comparative assessment (Karpinska et al., 2021; Chen et al., 2024a; Zheng et al., 2024b), on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
Except for pointwise evaluation, all forms of LLM-as-a-Judge suffer—or are susceptible to suffering—from position bias due to the intrinsic nature of position and comparison within the prompt structure.
B.2 Position Bias
Previous studies have discovered and investigated multiple types of biases, such as position bias (Zheng et al., 2024b; Li et al., 2024; Chen et al., 2024a; 2023; Chia et al., 2023; Chiang & Lee, 2023b; Khan et al., 2024; Zhu et al., 2023b; Zhong et al., 2023; Zheng et al., 2024a; Deb & Das, 2023; Zeng et al., 2023; Wang et al., 2023c; Qin et al., 2024; Wang et al., 2023b; Shi et al., 2024; Liusie et al., 2024; Liu et al., 2024; Kim et al., 2023; Karpinska et al., 2021; Hou et al., 2024; Kocmi & Federmann, 2023b; Pezeshkpour & Hruschka, 2023; Shen et al., 2023; Chiang & Lee, 2023a; Kocmi & Federmann, 2023a; Chen et al., 2024b; Li et al., 2023c; d; Raina et al., 2024), verbosity/length bias (Zheng et al., 2024b; Liu et al., 2024; Kim et al., 2023), self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b), selection bias (Zheng et al., 2024a; Pezeshkpour & Hruschka, 2023), and contextual bias (Liu et al., 2024; Zhou et al., 2024).
Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source (Li et al., 2023b; Chen et al., 2024c), proprietary commercial ones (OpenAI, 2023; Anthropic, 2024; Gemini Team, 2024), and fine-tuned models (Huang et al., 2024; Wang et al., 2023c; Zhu et al., 2023b; Li et al., 2023b; Kim et al., 2023).
To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content. For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs. Some other studies also use the term position bias (Ko et al., 2020; Aslanyan & Porwal, 2019; Blunch, 1984; Raghubir & Valenzuela, 2006), but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
Position bias is arguably the most prevalent and impactful bias among all. Chua et al. (2024) notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought (Wei et al., 2022) performance over self-training controls for all biases except position bias. Furthermore, Khan et al. Khan et al. (2024) point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
Moreover, there is ongoing debate over whether selection bias originates from position bias. Pezeshkpour & Hruschka (2023) argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity. Raina et al. (2024) contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
All of these findings underscore the critical importance of addressing the issue of position bias when employing LLM-as-a-Judge.
B.3 Deal with Position Bias
Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts (Zheng et al., 2024b; Raina et al., 2024; Pezeshkpour & Hruschka, 2023; Zhao et al., 2021; Zhu et al., 2023a). Also, LLM judges are vulnerable to attacks (Raina et al., 2024; Chen et al., 2024b; Shi et al., 2024).
Parse data with position bias There are many ways to deal with the position bias. The naive way is to exclude the inconsistent judgments(i.e. if the LLM judge gives a positionally biased judgment on a pair or list of model responses) (Zheng et al., 2024b; Chen et al., 2024b; Wang et al., 2023b; Li et al., 2023b). While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem. Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging (Zheng et al., 2024b; Raina et al., 2024; Li et al., 2023b; Wang et al., 2023b) or regard inconsistency as a half-win or tie for relation-based judging (Zheng et al., 2024b; Li et al., 2023c) after swapping the order of model responses in the prompt. For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B. The latter way of swapping + tie is proposed because intuitively the position bias is more likely to occur when the model responses get evaluated share a similar quality in terms of the evaluating metric. Our quantitative study also verifies this intuition, evidenced by the fact the LLM judge’s positional consistency (the percentage of positionally consistent pairs of judgment) is positively proportional to the answer quality gap between the model responses.
To save the expense of running the experiments more than once using swapped order, in practice, many of these studies also support a random-shuffle option in their code settings such that the baseline model for comparison does not remain in a fixed position.
Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping (Hou et al., 2024), split-and-merge (Li et al., 2023d), and multi-agent discussion (Li et al., 2023c; Khan et al., 2024). However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
Furthermore,Liu et al. (2024) suggests that existing calibration techniques designed to reduce bias, including context decomposition (Li et al., 2023d), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a), and batch calibration (Zhou et al., 2024), are insufficient to align LLM evaluators, even with supervised data.
Thus, position bias is a pervasive, significantly impactful, and challenging problem to solve.
Understand before solving We observe that the existing methods maybe ineffective or not enough satisfactory because there is a lack of understanding of the position bias. Although a variety of studies have researched this type of bias in LLM-as-a-Judge, a comprehensive understanding of what factors affect the position bias remains underexplored. In other words, without clarity on the key factors and their quantitative impact on position bias, the efficacy of current and future methods remains uncertain.
For instance, Li et al. (2023d) proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%. However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful. In other words, if the quality between the two model responses differs considerably, the LLM judges will make little position bias during judgment and hence is easy to calibrate. This example illustrates how lacking a comprehensive understanding of the factors affecting position bias can lead to overestimating or misjudging the effectiveness of methods proposed to resolve it and improve the performance of LLM-as-a-Judge.
Missing analysis for position bias Besides, the positional preference direction aspect of position bias remains underexplored. Positional preference refers to the specific positions an LLM judge favors when position bias is evident. For instance, if an LLM judge consistently favors the responses that appear first in the prompt, it exhibits a preference for the first position. Our study denotes primacy-preferred and recency-preferred to describe biases toward the first and second positions, respectively. We integrate the measurement by formally defining positional fairness score to quantify the extent of a judge model’s preference towards certain positions.
While prior research includes some positional preference results (Zheng et al., 2024b; Wang et al., 2023b; Zhu et al., 2023b; Li et al., 2023c; Liusie et al., 2024), none have conducted a detailed analysis. They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
Therefore, we propose that mitigating the position bias of LLM-as-a-Judge requires a simultaneous improvement in both positional consistency and positional fairness. A positionally consistent (high positional consistency) and positionally fair (nearly equal preference on primacy and recency when position bias occur) LLM judge is undoubtedly preferable. However, a trade-off between consistency and fairness/preference often occurs in practice. Previous studies that have provided results on positional preference have all shown that even state-of-the-art (SOTA) LLM judges struggle with both positional inconsistency and clear positional preferences.
Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions. MLLM-as-a-Judge (Chen et al., 2024a) conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments. In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness. If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
B.4 Summary of Prior Work
To summarize, LLM-as-a-Judge has a large potential in alternating human judges on a wide range of tasks, in particular subjective evaluation, owing to its cost-effectiveness, high agreement with human judgments, reproducibility, and scalability. However, it suffers from various biases, most notably position bias, which is prevalent across different evaluation tasks, models, and judgment types. Not only is it a type of bias that significantly hinders the improvement and promotion of applying LLM-as-a-Judge, but it is also difficult to solve due to its complexity and lack of understanding in the community.
Existing solutions proposed to address this issue may be ineffective or uncertain in efficacy due to the unclear impact and nature of influencing factors. Additionally, whether the position bias is an essential bias or mere random variation has not been validated from prior works. The positional preference side, reflecting the positional fairness of the LLM judge, also requires comprehensive analysis. In a nutshell, a comprehensive understanding of the position bias of LLM-as-a-Judge is crucial to validate existing and future approaches to address this important problem.
Appendix C LLM Agreement Analysis
Besides the exploration of position bias with a broad lens by average $PC$ and $PF$, instance-wise agreement between LLM judges is also insightful. Even two judges with the same $PC$ and $PF$ scores may not reach consensus on each instance. Therefore, this session investigates (1) what percentage of a set of evaluations do two LLM judges agree on each other? (2) how do the choices of all judges on an instance vary?
Figure 5: Disagreement Analysis and Mutual Agreement Heatmap of LLM judges on MTBench and DevBench. The blue bars of Figures (a) and (b) represent the number of instances with certain disagreement values, and the red cumulative curve displays the cumulative count and corresponding percentage over the benchmark dataset. The heat maps of Figures (c)(d)(e) marked higher mutual agreement values with a brighter color.
C.1 Mutual Agreement & Familial Property
We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations. We denote two judges agree on an instance if their judgment choices are identical. Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances. Fig. 5(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively. For MTBench that utilizes the 3-option mode, we also consider the “without tie” agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions. The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig. 5(d).
The heatmaps reveal clear “familial patterns” in the judgment choices of these LLM judges. For instance, the GPT-4, GPT-4-Turbo, and GPT-4o series exhibit high agreement on MTBench, achieving over 70% with ties included and over 85% without. GPT-3.5-Turbo didn’t agree with the GPT-4 series and o1-mini for around 40% of the instances, indicating that they are considerably different in judging capabilities.
For Claude-3 models, similar familial patterns could be observed. Claude-3-Opus highly agrees with Claude-3.5-Sonnet, probably due to their similar capabilities, while it also highly agrees with Claude-3-Sonnet, likely due to their similar model structure within the same series. Interestingly, Claude-3.5-Sonnet and Claude-3-Sonnet do not exhibit a significantly high agreement, indicating that the upgrade from series 3 to 3.5 considerably impacts their judging capabilities.
Gemini models exhibit rather low mutual agreement and “familial property” is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
These patterns suggest that familial similarities, possibly stemming from analogous model sizes, training data, and strategies, influence the positional preferences of these judges. In particular, the LLM judges could be primarily grouped by their capabilities; when judging capabilities are comparable, models within the same family series share a higher mutual agreement than across families.
Identifying such groupings provides valuable insights, as comparisons between judges from different groups, both adept at assessing LLM-generated content, can reveal distinct position biases and enrich our understanding of this phenomenon.
C.2 Disagreement & Benchmark Design Insight
Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful. Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority. By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement. For our study where twelve judges are investigated, the maximum disagreement of an MTBench is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode. On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig. 5(a) and Fig. 5(b), respectively. From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks. These are likely easy-to-evaluate instances, and the reliability of LLM judgments is enhanced by majority voting. In comparison, the instances with the highest disagreement are likely the ones that are difficult to evaluate and where the position bias is most likely to occur. However, luckily, these instances are rare, occupying only 0.03% and 0.07% of MTBench and DevBench respectively. In other words, majority voting of multiple capable LLM judges could be practically useful for over 99% of evaluation instances on both benchmarks.
Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig. 5(a) and Fig. 5(b) exhibit a balanced distribution of instances with varied difficulty. This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
To summarize, the practical implications of the disagreement analysis are three-fold. First, it helps identify the instances that are difficult or trivial to judge, benefiting benchmark designs to control the difficulty of evaluation by managing the number of these instances across the dataset. Second, it assists in filtering out instances where majority voting of LLM evaluators are likely to offer reliable judgments without direct comparison with human-annotated evaluations, enhancing the scalability of LLM judges especially when human evaluations are costly. In other words, if one-shot judgments from only one LLM judge are not enough reliable, multiple capable LLMs and the majoring voting strategy could be employed to make the evaluation more convincing. Last but not least, disagreement analysis provides a convenient way to make the difficulty variance of instances varied across the dataset tangible. Since the difficulty of an evaluation instance is closely related to the quality gap between the two solutions and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
Appendix D More Discussions on Answer Quality Gap
Intuitively, answer pairs with larger quality gaps are easier for LLM judges to evaluate, as one answer is clearly superior. Therefore, we expect LLM judges to exhibit weaker position bias in such cases. This motivates us to quantify the answer quality gap and investigate its impact on position bias. The quantification utilizes the overall win rate ($owr$) of a candidate model (i.e., answer provider) on the other to compute the quality gap. For a set of questions on a certain task and the corresponding solutions by two candidate models, an LLM judge’s evaluations would pick out the candidate winner. For the position consistent judgment instances, we can count the number of a candidate’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively. On the other hand, following the methodology introduced by Zheng et al. (2024b), we consider the inconsistent judgment pairs as a tie for both candidates, whose counting number is denoted as $C_{I}$. This follows the intuition that it is difficult to strictly pick out a better one between two responses with similar quality, hence increasing the likelihood of position bias.
Therefore, considering both cases, we define the overall win rate ($owr$) of a candidate’s solution over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs involving the two candidate solutions in both original and swapped settings. Then the quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, as 0.5 is the exact tie case for the win rate ranging from 0 to 1. We prefer the overall win rate over the consistent win rate (calculated by $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ to mainly address the data sparsity issue. If the proportion of position inconsistent judgments becomes high, the available data becomes sparse and $n_{C}$ becomes small. In the extreme case, where all judgments are inconsistent, $n_{C}$ is 0 and the consistent win rate is not even computable, making the measurement inaccurate and impractical. In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie” and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
Appendix E More Results of Position Bias and Answer Quality Gap Measurement
As shown in Fig. 4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness. In this session, we explore whether the discovery is consistent for each individual judge. Same as Section 2.2, we apply the overall win rate to reflect the answer quality gap for visualization.
Figure 6: Position Consistency and Preference Fairness vs. overall win rate for each judge on MTBench. Figure (a) refers to the relationship investigation of $PC$ and figure (b) for $PF$.
As shown in Fig. 6 (a), the “parabolic shape” is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency” applies to all models. However, Fig. 6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges. For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape”, GPT-4 and GPT-3.5 present nearly linear curves. In other words, while the former models align with the general tendency that a larger answer quality gap improves positional fairness, the latter ones preserve fairness regardless of the answer quality gap. This further demonstrates the necessity to investigate positional fairness in addition to consistency when evaluating a judge model’s position bias.
Appendix F Variable Selection and Tests
F.1 Bidirectional Stepwise Regression with AIC
Bidirectional stepwise regression is a combination of forward selection and backward elimination techniques. It iteratively refines the model by adding or removing predictors based on a statistical criterion—commonly the Akaike Information Criterion (AIC). The objective is to select a model that balances goodness of fit and complexity, aiming for the lowest AIC value.
The AIC is given by:
$\text{AIC}=2k-2\log(L),$ (4)
where $L$ is the likelihood of the model and $k$ is the number of parameters in the model, including the error variance $\sigma^{2}$. For a linear regression model with independent and identically distributed (iid) errors, $N(0,\sigma^{2})$, fitted to $n$ observations, the log-likelihood can be written as:
$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$ (5)
where $\hat{e}_{i}$ is the residual for the $i$th observation, and $\sigma^{2}$ is the variance of the errors. The AIC, in this context, becomes:
$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$ (6)
This form of the AIC balances the goodness of fit (as reflected by the residual sum of squares) and model complexity (as represented by $k$).
The operation of Bidirectional stepwise regression starts with either no predictors (forward selection) or all predictors (backward elimination), where the model iteratively adds or removes variables. Each step evaluates the impact on the AIC score. In forward selection, variables are added one by one, starting from the null model, such that the addition of each variable results in the largest decrease in AIC. In backward elimination, all variables are included in the model initially, and variables are removed one at a time, with the variable whose removal causes the smallest increase in AIC being dropped.
At each iteration, the change in AIC is computed as $\Delta\text{AIC}=\text{AIC}_{\text{new}}-\text{AIC}_{\text{current}}$, where $\text{AIC}_{\text{new}}$ refers to the AIC after adding or removing a variable, and $\text{AIC}_{\text{current}}$ is the AIC of the current model. If $\Delta\text{AIC}<0$, the model is improved by the addition or removal of the variable. The process terminates when neither adding nor removing variables results in a lower AIC, signifying that the most parsimonious model, based on AIC, has been reached.
F.2 Test results
We operated bidirectional stepwise regression on both benchmarks individually and together to identify the factors that are significantly contributing to position bias. Specifically, the variables include lengths (input, output, and prompt), answer quality gap, LLM judges, candidate models, and task categories to predict position consistency and preference fairness respectively. Table 5, 6 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively. Table 7, 8 serves for DevBench, and Table 3, 4 is conducted on the integrated set of both benchmarks. The impact of variables on the model is ranked from highest to lowest, from bottom to top. Removed variables listed as None indicate the full model at this given step.
Through benchmark testing, we verified that LLM judges, task categories, and the answer quality gap significantly contribute to position bias in terms of both position consistency and preference fairness. These findings align with our empirical results, showing that position bias varies notably by judge and task, with the answer quality gap being a key influencing factor. The extent of this impact is reflected by the magnitude of change in AIC when the given variable is removed. It is worth noting that while task output length remains a significant predictor for $PF$ and $PC$ in both benchmarks, the change in AIC magnitude after removing this variable is very minimal. This is consistent across both benchmarks individually and combined. We therefore conclude that, although position bias is influenced by task output length, this dependency is minimal.
Table 3: Final results of stepwise model selection for both benchmarks: Position Consistency
Removed Variables DF Sum of Sq RSS AIC
None 163.75 -18370
Task 20 2.832 166.59 -18319
Candidate 38 4.472 168.23 -18303
Quality gap 1 21.953 185.71 -17703
Judge 13 55.417 219.17 -16846
Table 4: Final results of stepwise model selection for both benchmarks: Preference Fairness
Removed Variables DF Sum of Sq RSS AIC
None 254.28 -16103
Task output length 1 0.836 255.12 -16088
Quality gap 1 11.339 265.62 -15873
Task 21 16.177 270.46 -15817
Judge 13 82.069 336.35 -14641
Table 5: Final results of stepwise model selection for MTBench: Position Consistency
Removed Variables DF Sum of Sq RSS AIC
None 61.974 -13312
Task output length 1 0.0553 62.029 -13311
Candidate 29 1.6474 63.621 -13282
Task 7 1.5304 63.504 -13244
Judge 13 15.3637 77.338 -12594
Quality gap 1 15.6206 77.594 -12559
Table 6: Final results of stepwise model selection for MTBench: Preference Fairness
Removed Variables DF Sum of Sq RSS AIC
None 129.00 -10909.2
Quality gap 1 1.931 130.93 -10861.3
Task 7 9.295 138.29 -10689.4
Judge 13 58.847 187.85 -9672.5
Table 7: Final results of stepwise model selection for DevBench: Position Consistency
Removed Variables DF Sum of Sq RSS AIC
None 55.382 -6940.2
Task output length 1 0.257 55.638 -6933.2
Candidate 9 1.514 56.896 -6905.4
Quality gap 1 13.128 68.510 -6525.3
Judge 13 84.760 140.141 -5146.6
Table 8: Final results of stepwise model selection for DevBench: Preference Fairness
Removed Variables DF Sum of Sq RSS AIC
None 60.104 -6753.9
Task output length 1 0.061 60.165 -6753.9
Candidate 9 0.731 60.834 -6748.2
Task 13 1.305 61.408 -6737.8
Quality gap 1 1.783 61.886 -6698.6
Judge 13 80.875 140.979 -5108.9
Appendix G DevBench
This session includes a similar baseline comparison analysis on DevBench as on MTBench. As shown in Fig.7, position bias is judge-dependent and task-dependent on DevBench as well, as $PC$ and $PF$ vary significantly across judges and tasks. Similarly, although GPT-4 stands as the baseline model with a generally high $PC$ across tasks, certain models achieve comparable or superior performances on certain tasks. For instance, for architecture design evaluations, GPT-4-Turbo, GPT-4o, and Gemini-1.5-pro all surpass GPT-4. Gemini-1.5-pro is especially outstanding, also exceeding GPT-4 in uml class evaluations. However, GPT-4 is still the best-performing model on UML sequence evaluations, with only GPT-3.5-Turbo can achieve comparable performance regarding certain detailed metrics (e.g., interaction complexity). These discoveries, aligning with the findings on MTBench, further necessitate the need to consider the trade-offs between positional consistency and fairness when selecting the optimal judge model for certain tasks.
Figure 7: Baseline comparisons of judges across tasks on DevBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be GPT-4 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
Appendix H Experiment Settings
This session specifies more detailed information about the judges, answer-generating models, tasks, and prompt templates used in this study. We choose to evaluate MTBench and DevBench for the following reasons: (1) all necessary information about the benchmark models, tasks, and questions is publicly available, making modifications convenient (2) they include a wide variety of answer-generating models, tasks, and task questions for a comprehensive evaluation (3) their human evaluations validated the reliability of state-of-the-art judging models (GPT-4 and GPT-4-Turbo) on their evaluation instances, hence model untested by prior work, if reaching high agreement with these validated judges, can be perceived reliable as well.
H.1 Judges, Candidates, and Tasks
Judge
In this study, we choose seven GPT, four Claude, and three Gemini models as the judges. The specific versions for API call are specified as follows: o1-mini-2024-09-12 for o1-mini, gpt-4o-2024-05-13 for GPT-4o, gpt-4-1106-preview for GPT-4-Turbo, gpt-4-0613 for GPT-4, and gpt-3.5-turbo-1106 for GPT-3.5-turbo; claude-3-5-sonnet-20240620, claude-3-opus-20240229, claude-3-sonnet-20240229, and claude-3-haiku-20240307 for Claude series. The other model names and versions are as they are.
Model
The reference (or baseline) answer-generating models are vicuna-13b-v1.3 for MTBench and human for DevBench. They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations. The other models that are compared to the reference models, namely “Model” in our context, are listed as follows.
  • MTBench (30): alpaca-13b, baize-v2-13b, chatglm-6b, claude-instant-v1, claude-v1, dolly-v2-12b, falcon-40b-instruct, fastchat-t5-3b, gpt-3.5-turbo, gpt-4, gpt4all-13b-snoozy, guanaco-33b, guanaco-65b, h2ogpt-oasst-open-llama-13b, koala-13b, llama-13b, mpt-30b-chat, mpt-30b-instruct, mpt-7b-chat, nous-hermes-13b, oasst-sft-4-pythia-12b, oasst-sft-7-llama-30b, palm-2-chat-bison-001, rwkv-4-raven-14b, stablelm-tuned-alpha-7b, tulu-30b, vicuna-33b-v1.3, vicuna-7b-v1.3, wizardlm-13b, wizardlm-30b
  • DevBench (10): codellama-7b-instruct, codellama-13b-instruct, codellama-34b-instruct, deepseek-coder-1.3b-instruct, deepseek-coder-6.7b-instruct, deepseek-coder-33b-instruct, gpt-3.5-turbo-1106, gpt-4-0125-preview, gpt-4-0613, gpt-4-1106-preview
The model names are exactly what MTBench Zheng et al. (2024b) and DevBench Li et al. (2024) used in their studies. That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4). In this study, we directly use the provided answers of these models to the task questions to form answer pairs and queries for the LLM judges.
Task
For tasks, we also follow the original studies of these two benchmarks, except for DevBench we separate the gerenal metrics into detailed ones and considered them as different tasks. In this sense, our study experiments on the following tasks to provide a comprehensive study on the positon bias of LLM-as-a-Judge:
  • MTBench (8): coding, extraction, humanities, math, reasoning, roleplay, stem, and writing.
  • Devbench (14):
    • UML class (4): cohesion_and_decoupling, complexity, practicability, and faithfulness
    • UML sequence (5): cohesion_and_decoupling, interaction_complexity, practicability, uniformity_and_integration, and faithfulness
    • architecture design (5): conformance, design_and_coding, practicability, uniformity_and_integration, and faithfulness
H.2 Prompt Settings
We follow the original prompt settings of MTBench and DevBench in our study of pairwise comparative LLM-as-a-Judge.
Though written differently, these prompts all share same key components:
  • A system prompt explaining the judging task and the role the LLM should be playing.
  • Emphasized “should” and“shouldn’t”s.
  • A prompt structure with placeholders for specific questions and model answers
  • A specified output format for later judgment extraction
  • Chain-of-Thought Wei et al. (2022) prompts requiring the LLM judge to provide reasons for its judgment
The detailed prompt templates are specified below.

### v5

@@
-Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs
+Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative LLM-as-a-Judge
 Abstract
-LLM-as-a-Judge presents a promising alternative to human evaluators across various tasks, but inherent biases, especially position bias — a tendency to favor solutions based on their position in the prompt — have compromised its effectiveness.
-Our study introduces a systematic framework to examine position bias in pairwise comparisons, focusing on repetition stability, position consistency, and preference fairness.
-This research significantly contributes to the field by introducing new concepts for understanding position bias and providing a multi-dimensional framework for evaluations.
-We conducted experiments with 12 LLM judges across MTBench and DevBench, covering 22 tasks and approximately 40 solution-generating models — candidates, resulting in over 100,000 evaluation instances.
-Our findings confirm that position bias in capable LLM judges is not due to random chances, along with notable variations observed across judges and tasks.
-Moreover, position bias is weakly influenced by the length of prompt components but significantly impacted by the quality gap between solutions.
-These insights can help optimize judge model selections, improve benchmark design, and inform future research on debiasing strategies, ultimately enhancing the reliability of LLM judges.
-1 Introduction
-In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications.
+LLM-as-a-Judge offers a promising alternative to human judges across various tasks, yet inherent biases, particularly position bias—a consistent preference for answers based on their position in the prompt—compromise its effectiveness.
+Our study proposes a systematic framework to investigate the position bias of LLM judges in the pairwise comparative scenario in terms of repetition stability, position consistency, and preference fairness.
+In pairwise comparative settings, LLM judges are prompted to choose the better candidate’s answer to the task question, where the task question and a pair of candidate answers are provided.
+Then the prompt with swapped order of candidate answers is given to the same judge again.
+If the judge prefers the same winner, it is denoted position consistent; otherwise, if different winners are selected after swapped settings, position bias is observed.
+To comprehensively understand the position bias of LLM judges, our experiments cover 12 judges across MTBench and DevBench with 22 tasks and around 40 answer-generating models, resulting in over 100,000 evaluation instances.
+Further exploration of the Judge-level, Candidate-level, and Task-level impactful factors on position bias enhances the understanding of this bias.
+Through comprehensive investigations, we validate that the position bias of capable LLM judges is not merely random variations.
+We also observe significant variance of position bias across judges and tasks, highlighting trade-offs of positional consistency, preference fairness, and practicality when selecting task-specific judges.
+Among impactful factors, position bias is largely influenced by the quality gap of given candidate answer pairs for evaluation.
+However, the lengths of the prompt components, including input questions, candidate answers, and the complete prompt, have minimal impact on position bias.
+In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs.
+These insights guide the selection of optimal judge models, enhance benchmark design, and lay the foundation for future research into effective debiasing strategies, ultimately improving the reliability of LLM evaluators for more robust, unbiased evaluation systems.
+Provocations
+Despite advancements in LLM-as-a-Judge and prevalent attention on position bias, the understanding of such bias is still limited and requires a more systematic evaluation.
+Enhancing the understanding of the impactful factors on position bias could inspire more effective mitigation strategies, laying the foundation for more reliable, robust, and unbiased evaluation systems.
+To this end, our study investigates aspects that are underexplored by prior works to offer a systematic evaluation framework for the position bias of LLM judges.
+Repetition Stability
+Validating LLM judges are repetition stable is a primary challenge that is often ignored by previous studies.
+They trusted the one-time generated judgments of the LLMs and conducted further analysis, implicitly assuming that the judgments represent the judge’s capabilities.
+However, the judgment choices and consequently observed position bias may be due to random variations.
+In other words, without verifying that LLM judges are stable in their choices across multiple trials, it is uncertain whether position bias stems from positional shift or is merely randomness.
+To address this gap, our study utilizes repetition stability to validate capable LLMs before further investigations on position bias.
+Preference Fairness
+While position bias is prevalent across most tasks and widely studied by literature, the main focus is on position consistency, as it is the most straightforward measurement of position bias.
+However, the preference directions when position bias occurs are largely underexplored by prior works.
+In this study, we show that previous measurements are either sensitive to the dataset size or overly penalizing highly consistent judges.
+To address this problem, we propose a more granular and scalable measurement that combines the strengths of both methods to assess preference fairness.
+We further stress the trade-off between position consistency, preference fairness, and practicality when choosing the task-specif judges, as position bias largely varies by task and judge.
+Answer Quality Gap
+From intuition, the difficulty of judging a pair of candidate answers (i.e., choosing the winner that better answers the task question) largely depends on the disparities of the answer pairs in quality.
+However, we are the first study to quantify the answer quality gap and investigate its correlation with position bias.
+Our investigations verify that as the answer quality gap enlarges, implying one candidate’s answer is significantly better than the other, the judges become more constant and fair.
+The findings not only benefit the interpretation of other LLM judge biases like verbosity/length bias and self-enhancement bias, but also benefit future benchmark designs for more social goods.
+To summarize, this study proposes new metrics and studies more impacting factors on position bias that are underexplored by prior studies, offering a more comprehensive and systematic framework for understanding and evaluating the position bias of LLM judges.
+We stress the need to have a deeper understanding of the complex bias before moving to solutions.
+More detailed analysis, insights, findings, and contributions can be found in the Appendix.
+[…]
+[appendices]
+Appendix Table of Contents
+[…]
+[appendices]0
+Appendix
+Appendix A Introduction
+In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global attention and stimulating substantial research into their applications.
 Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields.
-While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations (Zeng et al., 2023; Karpinska et al., 2021).
+While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations [1, 2].
 To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks.
-Despite a high level of agreement with human judgments(Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
+Despite a high level of agreement with human judgments[3, 4, 5], inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
 Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig.
 1.
-This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations (Qin et al., 2024), for any score-based and relation-based settings (Li et al., 2023d).
+This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations [6], for any score-based and relation-based settings [7].
 Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias.
-However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks (Khan et al., 2024; Chua et al., 2024).
+However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks [8, 9].
 Moreover, the limited understanding of the bias itself has exacerbated the problem.
@@
 Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge.
-When the positions of candidate model solutions are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same candidate model solution.
+When the positions of model responses are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same response.
 We observe repetition instability if the same LLM judge gives different judgments across repetitions.
@@
 Another key aspect of position bias, preference fairness, has also been under-explored by previous works.
-Most studies have primarily focused on position consistency to measure how frequently position bias occurs across a set of evaluation instances.
+Most studies have primarily focused on positional consistency to measure how frequently position bias occurs across a set of evaluation instances.
 However, the direction of biased preference (i.e., primacy or recency) is also insightful.
 Fig.
-2 illustrates cases where a pair of judgments with swapped prompts become position consistent, inconsistent, primacy-preferred, or recency-preferred.
+2 illustrates cases where a pair of judgments with swapped prompts become positionally consistent, inconsistent, primacy-preferred, or recency-preferred.
 We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations.
-Specifically, judgments are considered to have a fair preference if they are position consistent or if the preference directions are evenly distributed.
-In contrast, an LLM judge is considered to have an unfair preference if it consistently favors candidate model solution that appear either first or last in the prompt.
+Specifically, judgments are considered to have a fair preference if they are positionally consistent or if the preference directions are evenly distributed.
+In contrast, an LLM judge is considered to have an unfair preference if it consistently favors responses that appear either first or last in the prompt.
 Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
-Figure 2: Demonstration of position consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better candidate solution.
+Figure 2: Demonstration of positionally consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better response.
 A judgment is consistent if it selects the same winner across different prompt positions.
-Primacy/recency indicates a preference for the first/last candidate solution, respectively.
-Position inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
+Primacy/recency indicates a preference for the first/last model response, respectively.
+Positionally inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
 To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment.
-We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring (Zheng et al., 2024b; Liusie et al., 2024), making it a practically preferable choice for investigations.
+We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring [3, 10], making it a practically preferable choice for investigations.
 Our contribution can be summarized as following: (1) introducing a systematic framework for evaluating LLM-as-a-Judge’s position bias in terms of Repetition Stability, Position Consistency, and Preference Fairness; (2) identifying Judge-level, Candidate-level, and Task-level factors that affect position bias; and (3) providing insights into future benchmark design and paving the way for effective mitigation strategies for position bias.
-2 Methods & Defintions
+Appendix B Methods & Defintions
 LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments.
 Our study focuses on relation-based assessment, specifically under the pairwise comparison setting, for several reasons.
-First, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields (Qin et al., 2024).
-Second, as noted by Liu et al.
-(2024), LLM and human evaluations are more aligned in the context of pairwise comparison than any of the score-based assessments.
-Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of position consistency(Zheng et al., 2024b; Liusie et al., 2024).
-Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, through some advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024).
+First, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields [6].
+Second, as noted by [11], LLM and human evaluations are more aligned under the context of pairwise comparison than any of the score-based assessments.
+Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of positional consistency[3, 10].
+Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, by some advanced ranking algorithms [6, 11].
 Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
-In the context of pairwise comparative assessments, LLM judges are prompted to choose the better candidate solution to the task question.
+In the context of pairwise comparative assessments, LLM-judges are prompted to choose the better response that answers the question.
 To accommodate the possibility of a tie, various option modes are introduced.
-Two-Option mode restricts judges to choosing the better solution between two options, labeled {A} for the first and {B} for the second as shown in Fig.
+The Two-Option mode restricts judges to choosing the better response between two options, labeled {A} for the first and {B} for the second as shown in Fig.
 1.
-Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither solution is preferable.
-Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie.
-In our experiments, we employed Three-Option mode for MTBench (Zheng et al., 2024b) and Two-Option mode for DevBench (Li et al., 2024), aligning with their respective original frameworks.
-These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM judges.
+The Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither option is preferable.
+The Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie.
+In our experiments, we employ the Three-Option mode for MTBench [3] and the Two-Option mode for DevBench [4], aligning with their respective original frameworks.
+These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM-judges.
 To setup our experiment, as shown in Fig.
-1, the system prompt, option mode, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the winning solution.
+1, the system prompt, option mode, task question, answers/responses from the two candidate models (original prompt) are presented to the LLM-judges to select the winning answer.
 The experiment is conducted in a double-blind setting.
-The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question.
-Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair.
+The identities of the candidate models are hidden from the LLM-judges, and the candidate models are unaware that their responses will be compared to another model when answering the question.
+Then, the prompt with answers in a swapped position (swapped prompt) is given to the same judge again, which results in a choice pair.
 As shown in Fig.
-2, if the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent with a fair preference.
-Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
+2, if the choice pair consistently favors the same answer regardless of the swapped position, it is considered positionally consistent with a fair preference.
+Conversely, if the LLM-judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
 To systematically investigate position bias in pairwise comparative judging scenarios, we evaluate LLM judges using three key metrics: Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$).
@@
 Therefore, for each judge, candidate, and task, we calculate $RS$, $PC$, and $PF$ to thoroughly explore the variance in position bias and its impacting factors.
-2.1 Evaluation Metrics
-We validate the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows.
+B.1 Evaluation Metrics
 Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times.
-It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations.
+It is essential to determine whether the choices of LLM judges, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations.
 We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset.
 This metric is formalized as
-$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,...\}$ (1)
+$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,D\}$
 where $S$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries.
 The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
-Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is swapped.
-It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the judge model prefers the same winning solution.
+Position Consistency ($PC$) quantifies how frequently a judge model selects the same response after the order of responses is swapped.
+It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the model selects the same winner.
 Formally, it is calculated as
-$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}},V=\{(A,B),(B,A),(C,C),...\}$ (2)
+$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}}$
 where $(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})$ denotes the judgment pair for the $j^{th}$ query, $n$ represents the number of prompt pairs, and $V$ is the set of choice pairs that correspond to position consistency.
@@
 This formula provides a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
-Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain solution positions.
+Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain response positions.
 In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency.
-These terms replace the more verbose “preference for the first/second position” used in previous studies, ensuring clarity and generalization for future research.
+These terms replace the more verbose “preference for the first/second position" used in previous studies, ensuring clarity and generalization for future research.
 The examples of such preferences are demonstrated in Fig.
 2.
-Position consistent judgments are fair in preference.
-The measurements become more complex for position inconsistent cases.
+Positionally consistent judgments are fair in preference.
+The measurements become more complex for positionally inconsistent cases.
 Previous studies proposed two common ways to measure the preference fairness.
 A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$).
-The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023b).
+The counts are then normalized by the total number of prompt pairs [3, 5].
 However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
-Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently.
-They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs.
+Alternatively, instead of normalizing over the complete dataset, studies like [12, 10] treat positionally inconsistent evaluation instances independently.
+They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of positionally inconsistent pairs.
 We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$.
-However, this approach is problematic because it overlooks the fact that “position consistent judgments are also position fair,” which can lead to overly penalizing highly consistent LLM-judges.
+However, this approach is problematic because it overlooks the fact that “positionally consistent judgments are also positionally fair," which can lead to overly penalizing highly consistent LLM-judges.
 For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred.
@@
 The $PF$ score is formally calculated by
-$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$ (3)
+$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$
 where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively.
 This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero.
-Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the position consistent judgments.
-Another advantage of this approach is that it quantifies preference fairness with a single score, rather than the two separate scores used in previous studies.
+Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the positionally consistent judgments.
+Another advantage of this approach is that it quantifies positional preference with a single score, rather than the two separate scores used in previous studies.
 The $PF$ score is interpreted as follows:
-$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
+$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Positionally Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
 To summarize, our proposed preference fairness score is superior to prior measurements due to its ability to provide a single, comprehensive score that accounts for all evaluation instances, while maintaining its sensitivity across datasets, even when the number of questions and instances varies for each task.
-2.2 Factors affecting position bias
-Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose solutions to the task-specific questions are evaluated by the judges.
-Factors that have a significant impact on position bias are highlighted with an asterisk (*) by bidirectional stepwise regression tests and marked red by empirical findings.
-‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s solutions to the question, making it both a Candidate-level and Task-level factor.
+B.2 Factors affecting position bias
+Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose responses to the task-specific questions are evaluated by the judges.
+An asterisk (*) highlights the factors with significant impact on position bias that we found through experiments.
+‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s answers to the question, making it both a Candidate-level and Task-level factor.
 ‘Prompt’ encompasses the complete query presented to the judges, including Task Input, Task Output, and system prompts.
@@
 Task Input Length $\times$ $\times$ $\checkmark$
-Task Output Length $\times$ $\checkmark$* $\checkmark$*
+Task Output Length $\times$ $\checkmark$ $\checkmark$
 Prompt Length $\times$ $\checkmark$ $\checkmark$
@@
 By design, our framework allows for the integration and assessment of additional influencing factors.
-Among the influencing factors we identified, we chose “familial property” for Judge-level factors instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments.
+Among the influencing factors we identified, we chose “familial property” instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments.
 The model family grouping reflects similar model sizes and training specifics of the model members.
 The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version.
-However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable” models sharing similar judging capabilities.
+However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable" models sharing similar judging capabilities.
 More details and discussions about the familial property can be found in Appendix Sec.
-C, where we examined the LLM mutual agreements to study the similarities and differences across families.
-Answer quality gap, based on our findings, is highly impactful yet under-explored in prior research.
-We define the quality of a candidate’s answer/solution by how effectively it addresses the question.
-Consequently, the answer quality gap refers to the disparity in quality between the solutions from two candidate models to the same question and is considered the Candidate-level (i.e., solution provider) factor.
-Ideally, when a reliable LLM judge is presented with a question and corresponding solution pairs, it would prefer the higher-quality solution, where the corresponding candidate is denoted as the winner selected by the judge.
-Following this assumption, we measure the answer quality gap by the win rates between two candidates on a set of tasks and questions.
-We regard a tie judgment as ‘half-win’ for both candidates when calculating its win rate over another, depending on option mode.
-If position bias occurs, the winner may be inconsistent when the order of candidate solutions is swapped in the query.
-Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen in both the original and swapped queries (termed “consistent wins”), cases where both responses are deemed equal (termed “consistent ties”), and cases where different winners are selected after the solutions are swapped.
-We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment pairs ($C_{I}$), respectively.
-Following Zheng et al.
-(2024b), we count inconsistent judgment pairs also as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
-Inspired by Zheng et al.
-(2024b), we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs.
+H, where we examined the LLM mutual agreements to study the similarities and differences across families.
+Answer quality gap, based on our findings, is a highly impactful yet under-explored factor in prior research.
+We define the quality of a model’s answer by how effectively it addresses the question.
+Consequently, the answer quality gap refers to the disparity in quality between two candidate model answers to the same question and is considered a Candidate-level (i.e., response provider) factor.
+Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs, it would prefer the higher-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
+Depending on option mode, we regard a tie judgment as‘half-win’ for both candidates when calculating its win rate over another.
+Then, the win rates between two candidates on a set of task questions can reflect the answer quality gap between them on the task.
+However, if position bias is exhibited, the winner may be inconsistent when the order of candidate responses is swapped in the query.
+Therefore, we observe the LLM judgments where the absolute winner is consistently chosen in both original and swapped queries (termed “consistent wins”), both responses are deemed equal (termed “consistent ties”), and different winners are selected after candidate responses are swapped in the query.
+We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent choice pairs ($C_{I}$), respectively.
+Following [3], we count inconsistent choice pairs as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
+Inspired by [3], we define the overall win rate ($owr$) of a model’s response over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of choice pairs.
 Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, where 0.5 is the overall win rate when we only have ties.
-In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
+In contrary to using consistent win rate by [3] (calculated as $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of positionally consistent judgment pairs) to quantify $\delta_{q}$, the adoption of overall win rate incorporates all data points and captures the “comparable quality” scenario, where responses in similar quality might lead to positionally biased judgments, a scenario that the consistent win rate might overlook.
 More discussions can be found in Appendix.
 Sec.
-D.
-3 Experiments
-3.1 Experiment Settings
-In this study, we evaluated twelve up-to-date commercial models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework.
-Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks.
-However, due to their limited performance on our tasks, as shown in Table 2, we focused our investigation solely on more powerful closed-source models.
-We included the results of open-source models for completeness.
-We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks.
+I.
+Appendix C Experiments
+C.1 Experiment Settings
+In this study, we evaluated twelve up-to-date commercial models from the GPT [13], Claude [14], and Gemini [15] series using our systematic framework.
+Three Llama models [16], with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and well-known capability in reasoning tasks.
+However, due to their limited performance on our tasks, as shown in Table 2, we focus our investigation solely on more powerful closed-source models.
+These models are more widely used and reliable, though we still include open-source models to gain more comprehensive insights.
+We adopted the modified MTBench [3] and DevBench [4] datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks.
 We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
 MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task.
-We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges.
+We then paired these candidate models with the baseline candidate for evaluation by the LLM judges.
 The prompt templates we used are identical to those used in the benchmarks.
@@
 Sec.
-H.
+L.
 To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials.
-This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench.
+This resulted in 576 and 432 instances per judge for MTBench and DevBench, respectively.
 The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results.
@@
 In total, approximately 100,000 evaluation instances were analyzed in this study.
-To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks.
-We used variables such as average lengths of input, output, and prompt; answer quality gap; LLM judge series; candidate identities; and task categories to predict $PC$ and $PF$, respectively.
-Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score.
-This process involves both forward selection and backward elimination, with each ”step” testing whether including or excluding a variable improves the model’s AIC value.
-Further details about the process can be found in Appendix.
-Sec.
-F.
-3.2 Empirical Results and Interpretations
+C.2 Empirical Results and Interpretations
 The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2.
@@
 A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not merely random variations.
+This further necessitates the need for investigating and analyzing preference fairness simultaneously with position consistency.
 Fig.
@@
 7 demonstrate that position bias varies by judges and tasks significantly.
+Together with Table 2, it stresses the trade-off between $PC$ and $PF$ when selecting task-specific LLM judges.
 Fig.
 4 further investigates the impact of answer quality gap and lengths of prompt components on position bias.
-These analyses were conducted by considering all judges together on MTBench.
-The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E, and Sec.
-G.
-Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables.
-Similarly, these factors also contribute significantly to Preference Fairness.
-Additionally, we found that average output length is a statistically significant predictor of $PF$.
-This finding is not surprising, as longer outputs are generally perceived as higher quality and more preferred.
-Quantitative results and more discussions can be found in Appendix.
-Sec.
-F.
+These analyses are conducted by considering all judges together on MTBench.
+The detailed analyses for each judge and on DevBench can be accessed from Appendix.Sec.J, and Sec.K.
 Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Positional Fairness ($PF$).
-Top 5 performances are marked in bold.
-“Error” is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format.
-High error rates and low $RS$ are marked red, implying further evaluations (i.e., $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data.
+Top 5 performances are marked in bold.“Error" is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format.
+High error rates and low $RS$ are marked red, implying further evaluations (i.e., $RS$, $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data.
 Therefore, open-source models (e.g., Llama) are not investigated in the paper because of their high error rates under the difficulty and length of our evaluation instances.
 In contrast, up-to-date commercial models generally demonstrated minimal error and high repetition stability, making them more practical and meaningful for investigations.
-Chatbot Arean Leaderboard’s overall ranking of the models is listed.
-Judge MTBench DevBench Arena
+Judge MTBench DevBench
 $RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error
-Caude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 7
-Claude-3-Opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00 17
-Claude-3-Sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00 37
-Claude-3-Haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00 46
-Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 12
-Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96 25
-Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00 66
-o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00 2
-GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 6
-GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00 17
-GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00 52
-GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 94
-Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 99
-Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 109
-Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94 57
+Caude-3.5-sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00
+Claude-3-opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00
+Claude-3-sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00
+Claude-3-haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00
+Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13
+Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96
+Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
+o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00
+GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00
+GPT-4-turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
+GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
+GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00
+Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
+Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
+Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
 Figure 3: Baseline comparisons of judges across tasks on MTBench.
@@
 Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$.
-The baseline model is chosen to be GPT-4 for its generally superior performances.
+The baseline model is chosen to be gpt-4-0613 for its generally superior performances.
 Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
@@
 The irregular and inconclusive patterns imply that position bias is not significantly impacted by the length of prompts and their components.
-4 Main Findings
-Position Bias of Capable Judges are not Mere Random Variations As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal ”Error” rates, generally exhibit $RS$ values above 0.85.
-The most up-to-date and capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks.
-These results confirm that LLM judgments, and the resulting position bias, are not merely random variations.
-This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
-Position Bias Varies by Judge & Task As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig.
+Appendix D Main Findings
+Position Bias of Capable Judges are not Mere Random Variations
+As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal “Error" rates, generally exhibit $RS$ of over 0.85.
+The up-to-date and most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, Gpt-4, and GPT-4o, all achieve near-perfect $RS$ of over 0.95 on both benchmarks.
+These results validate that LLM judgments and the consequently observed position bias are not merely random variations.
+This increases the confidence that one-time generated judgments by these validated LLMs adequately represent their judging capabilities.
+Only with this prerequisite, can we ensure further investigations are reasonable and meaningful.
+Position Bias Varies by Judge & Task
+As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig.
 3 for $PC$ comparisons.
@@
 Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it may be employed as a cost-effective alternative to coding evaluations.
-Additionally, significant variations of $PF$ across tasks are observed.
-The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on the others.
+Similarly, significant variances of $PF$ across tasks are observed.
+The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others.
 Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, roleplay, and writing tasks.
 Even for uniformly recency-preferred judges such as Claude-3’s and Gemini-pro’s, the extent of biased preference, as reflected by $PF$ values, varies by task.
-Moreover, high position consistency does not guarantee fairness.
+Moreover, high position consistency does not guarantee fairness, weighing the trade-offs between position consistency and preference fairness.
 For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively.
 In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
-Therefore, the position bias of LLM judges is judge-dependent and task-dependent.
-This observation is confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$.
+In this sense, sacrificing consistency for more fairness or vice versa can all be practically useful depending on specific needs.
+This trade-off between consistency and fairness is particularly crucial when the most consistent model is not the fairest.
+Therefore, to sum up, the position bias of LLM judges is task-dependent and there is usually a trade-off between position consistency and preference fairness.
 In practice, to achieve the most reliable evaluation results, one needs to select task-specific LLMs whose consistency and fairness are in balance.
-Moreover, as shown in Table 2, a high ranking in the overall reasoning performance of a model does not guarantee its judging reliability.
-Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec.
-G.
-Position Bias Correlates to Answer Quality Gap Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
-In this study, as defined in Section 2.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$).
+The same discoveries can be concluded from DevBench results, as detailed in Appendix.Sec.
+K.
+Position Bias Correlates to Answer Quality Gap
+Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
+In this study, as defined in Section B.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$).
 Since an absolute tie has 0.5 $owr$, $\delta_{p}$ increases as $owr$ extends from 0.5 to 0 or 1.
-Fig.
-4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$.
+Fig.4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$.
 This aligns with our intuition that the answer pairs with larger quality disparities are easier to achieve consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$.
-The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.
-E.
+The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.J.
 Similarly, as shown in Fig.
-1(b), judgments generally become more preference fair as $\delta_{q}$ increases.
+1(b), judgments generally become more positionally fair as $\delta_{q}$ increases.
 However, the extent is not as significant as for $PC$.
-Also, the relationship varies by judge, as some LLMs maintain preference fairness regardless of $\delta_{q}$.
-For example, as shown in Fig.
-6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
-These observations align with the right-arrow shape as demonstrated in Fig.
-4(c), where there is a general trend that judgments become preference fairer as position consistency increases.
-It also justifies the reasonableness of our quantification of preference fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
+Also, the relationship varies by judge, where some LLMs maintain preference fairness regardless of $\delta_{q}$.
+For example, as shown in Fig.6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
+These observations align with the right-arrow shape as demonstrated in Fig.4(c), where there is a general trend that judgments become preference fairer as position consistency increases.
+It also justifies the reasonability of our quantification of positional fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
 Together, we conclude that as the answer quality gap enlarges, judges generally become more position consistent and preference fair according to the regression curves.
@@
 Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
-Our investigation also helps to explain the observed verbosity/length bias (Liu et al., 2024; Kim et al., 2023; Zheng et al., 2024b) and self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b) of LLM judges.
+Our investigation also helps to explain the observed verbosity/length bias [11, 17, 3] and self-enhancement bias [12, 3] of LLM judges.
 From our analysis, LLMs may prefer longer answers (verbosity/length bias) since more capable models usually generate longer solutions with higher qualities.
-Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the solutions generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4).
+Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the answers generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4).
 In both cases, the observed biased preference may be fundamentally due to a prominent answer quality gap instead of length or self-identity.
-Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length).
-Since the tokenization methods vary by models, we apply the length of the prompt component strings for uniformity and scalability.
+Position Bias is Length-Independent
+We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the response length of answer-generating models (task output length), and the length of the entire prompt (prompt length).
+Since the tokenization methods vary by model, we apply the length of the prompt component strings for uniformity and scalability.
 The analysis from Fig.
-4(d)(e)(f) shows that there is very weak relationship between the lengths of prompt components and position bias.
-By stepwise regression, the average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table.
-4.
+4(d)(e)(f) shows that there is no significant and definite relationship between the lengths of prompt components and position bias.
 Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
-Practical Implications for Benchmark Design We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appendix.Sec.
-C to gain further insights.
-In particular, the disagreement analysis helps identify instances where it is easy or difficult for judges to reach a consensus, reflecting the complexity of the judgment task.
-Based on our findings regarding the answer quality gap and LLM agreement/disagreement, this study offers practical implications for designing evaluator benchmarks that control the difficulty of judging tasks.
-Our findings show that the hardest-to-judge instances are those where (1) LLMs frequently disagree with each other, (2) comparable win rates and minimal quality gaps exist between the two candidate models, and (3) significant position bias is exhibited by the majority of judges.
-One qualitative analysis example of a difficult-to-judge instance comes from MTBench question 120, which involved a math problem where both candidate models, Vicuna-13b-v1.3 and WizardLM-30b, provided incorrect solutions.
-The problem was: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find $x$ such that $f(x)=0$.
-Vicuna-13b-v1.3 incorrectly applied the quadratic formula to solve the cubic equation, while WizardLM-30b used an enumeration method but failed to identify the correct answer.
-This resulted in a 4{A}-3{B}-5{C} choice pattern among the twelve judges, illustrating a typical case of significant disagreement.
-We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer.
-Others preferred either correctness (choosing {C} as a tie) or methodological soundness (choosing {B}, favoring WizardLM-30b).
-In the swapped order setting, 9 out of 12 judges displayed position bias.
-This highlights the challenge of judging solutions with minimal quality gaps, leading to greater position bias and subjectivity.
-The key takeaway is that reducing hard-to-evaluate, highly subjective instances can improve benchmark design.
-Likewise, minimizing trivial cases where judges easily agree would enhance the benchmark’s robustness.
-Clearer and more specific evaluation criteria could also help resolve ambiguity in difficult instances.
-5 Conclusion
-In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of LLM judges in pairwise comparative scenarios in terms of repetition stability, position consistency, and preference fairness.
-Through comprehensive evaluations of 12 judges on two benchmarks across 22 tasks with over 100,000 evaluation instances, we observe significant variations of position bias across judges and tasks.
-We also discover that position bias is weakly length-dependent but significantly impacted by the answer quality gap.
-Our findings enhance the understanding of position bias, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
-Appendix Table of Contents
-[…]
-[appendices]0
-Appendix
-Appendix A Contribution & Limitation
+Practical Implications for Benchmark Design
+We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appedix.Sec.
+H to gain more insights.
+Particularly, the disagreement analysis helps to identify the instances where it is easy/difficult for all judges to reach a consensus, reflecting the difficulty of judging them.
+Based on discoveries of the answer quality gap and LLM agreement/disagreement, our study offers practical implications for evaluator benchmark designs in controlling the difficulty of judging tasks.
+From our findings, the hard-to-judge instances are the ones where (1) LLMs mostly disagree with each other (2) comparable win rates and minimal quality gaps are observed for the two candidate models (3) prominent position bias exhibits for a majority of judges.
+Here is a qualitative analysis of a typical difficult-to-judge example.
+This instance (MTBench question 120) involved a math problem where both models, vicuna-13b-v1.3 and wizardlm-30b, provided incorrect answers.
+The problem statement is as follows: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find x such that $f(x)=0$.
+Methodologically, vicuna-13b-v1.3 erroneously applied the quadratic formula to solve the cubic equation, whereas wizardlm-30b employed an enumeration method but failed to identify the correct answer.
+This resulted in a 3{A}-4{B}-5{C} choice pattern among the twelve judges, a typical case of significant disagreement.
+Qualitatively, We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer.
+The remaining ones showed preferences either for correctness (which then chose {C} as a tie) or methodological soundness (which then chose {B}, i.e., preferring wizardlm-30b).
+Notably, in the swapped order setting of the same problem, 8 of the 12 judges encountered position bias.
+This is a typical illustration of the two responses having a minimal quality gap, leading to difficulty in judging, the likelihood of position bias, and prominent subjectivity during evaluations.
+The takeaway message from examples like this is that the number of hard-to-evaluate and highly subjective instances could be reduced for better benchmark designs.
+Similarly, the trivial-to-judge instances where all judges easily reach a consensus could be reduced among the benchmark population.
+One could also specify more detailed and definite evaluation criteria to erase the ambiguity of some hard-to-evaluate instances.
+Appendix E Conclusion
+In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of pairwise comparative LLM-as-a-Judge in terms of repetition stability, position consistency, and preference fairness.
+We focus on position bias for its prevalent occurrences and significant impact on the reliability of LLM judges’ evaluations.
+From prior works, the pairwise comparative assessment offers the most consistent and reliable judgments compared to other types of LLM-as-a-Judge and hence becomes our concentration for practicality.
+The study starts by investigating the repetition stability of LLM judges to validate that the judgments and consequently observed position bias are not mere random variations.
+Then, in addition to the traditional measurement of positional consistency, we propose a systematic way to quantify the preference fairness of LLM judges.
+Our comprehensive evaluation of twelve judge models on two benchmarks across 22 tasks reveals significant variance of position bias by judge and task.
+This stresses the trade-offs between position consistency, preference fairness, and cost-effective practicality when choosing the optimal judges for specific tasks.
+The study on the Judge-level, Candidate-level, and Task-level impactful factors on position bias further enhances the understanding of this bias.
+For example, we are the first the make a direct connection between answer quality gap and position bias, where our discovery reveals that the quality gap between candidate responses to the task question significantly influences the positional consistency and preference fairness of LLM judges.
+However, position bias is length-independent, meaning that the lengths of prompt components have minimal impact on position bias.
+In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs.
+To summarize, insights from this study enhance the understanding of position bias and lay the groundwork for more robust, unbiased evaluation systems.
+Appendix F Contribution & Limitation
 Contributions
 Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge.
-Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through position consistency and preference fairness.
+Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through positional consistency and positional fairness.
 We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components.
@@
 First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation.
-Our focus and investigation on preference fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability.
-Moreover, our emphasized trade-offs between position consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work.
+Our focus and investigation on positional fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability.
+Moreover, our emphasized trade-offs between positional consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work.
 Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results.
-Besides, the LLM agreement/disagreement analysis in Appendix.Sec.C 5 provides an applicable tool to identify the difficulty of judging certain instances.
+Besides, the LLM agreement/disagreement analysis in Appendix.5 provides an applicable tool to identify the difficulty of judging certain instances.
 Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias.
-Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare Chen et al.
-(2024c), instruction following and prompting Zeng et al.
-(2023); Chua et al.
-(2024), multimodal assessment Chen et al.
-(2024a), and recommender systems Hou et al.
-(2024).
+Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare [18], instruction following and prompting [1, 9], multimodal assessment [19], and recommender systems [20].
 Strength
@@
 It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works.
-(2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, position consistency, and preference fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias.
+(2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, positional consistency, and positional fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias.
 (3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
@@
 Then our framework offers better control on the other impactful factors such as the answer quality gap to make these further investigations more accurate.
-Appendix B Related Work
-B.1 LLM-as-a-Judge
+Appendix G Related Work
+G.1 LLM-as-a-Judge
 In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications.
-For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature (Zeng et al., 2023).
-However, it lacks scalability and reproducibility (Karpinska et al., 2021).
-As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks (Zheng et al., 2024b; Li et al., 2024; Chiang & Lee, 2023b; Liusie et al., 2024; Chen et al., 2023; Fu et al., 2023; Wang et al., 2023a; Kocmi & Federmann, 2023b; Ji et al., 2023; Liu et al., 2024; Zeng et al., 2023; Liu et al., 2023; Shen et al., 2023; Kocmi & Federmann, 2023a; Chiang & Lee, 2023a; Karpinska et al., 2021; Dubois et al., 2024; Kim et al., 2023; Liang et al., 2023; Chang et al., 2023; Zhong et al., 2023; Chia et al., 2023; Deb & Das, 2023; Ye et al., 2024; Min et al., 2023; Li et al., 2023d; c; b; Lu et al., 2024), including open-ended story generation (Chiang & Lee, 2023a), adversarial attacks (Chiang & Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi & Federmann, 2023a), and instruction following (Zeng et al., 2023).
-These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles (Zheng et al., 2024b; a; Pezeshkpour & Hruschka, 2023; Chiang & Lee, 2023b; Shen et al., 2023; Liusie et al., 2024; Qin et al., 2024; Chen et al., 2023; Huang et al., 2024; Zhu et al., 2023b; Raina et al., 2024; Chen et al., 2024b; Wang et al., 2023b; Liu et al., 2024; Zeng et al., 2023; Khan et al., 2024; Kim et al., 2023; Li et al., 2023b; d; Shi et al., 2024; Li et al., 2023c; Chen et al., 2024c; Chua et al., 2024; Chen et al., 2024a).
+For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature [1].
+However, it lacks scalability and reproducibility [2].
+As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks [3, 4, 21, 10, 22, 23, 24, 25, 26, 11, 1, 27, 28, 29, 30, 2, 31, 17, 32, 33, 34, 35, 36, 37, 38, 7, 12, 39], including open-ended story generation [30], adversarial attacks [21], summarization [2], machine translation [29], and instruction following [1].
+These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles [3, 40, 41, 21, 28, 10, 6, 22, 42, 5, 43, 44, 45, 11, 1, 8, 17, 39, 7, 46, 12, 18, 9, 19].
 As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation.
@@
 LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
-In many of these tasks, LLM judges have shown a high level of agreement with human evaluators (Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), yet in some tasks, they are less effective, largely due to inherent biases (Shen et al., 2023; Chiang & Lee, 2023b; Zheng et al., 2024a).
+In many of these tasks, LLM judges have shown a high level of agreement with human evaluators [3, 4, 5], yet in some tasks, they are less effective, largely due to inherent biases [28, 21, 40].
 Even in cases where agreement is high, judgments may still suffer from biases.
 When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method.
-From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment (Qin et al., 2024).
-By comparative method, it can be either score-based or relation-based (Li et al., 2023d).
-For example, pointwise scoring (Kim et al., 2023) lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics.
-Pairwise/Listwise Scoring (Karpinska et al., 2021; Zhu et al., 2023b; Chen et al., 2024a; Zheng et al., 2024b) prompts the LLM judges to score a pair/list of model-generated answers.
+From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment [6].
+By comparative method, it can be either score-based or relation-based [7].
+For example, pointwise scoring [17] lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics.
+Pairwise/Listwise Scoring [2, 5, 19, 3] prompts the LLM judges to score a pair/list of model-generated answers.
 Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst).
-Pairwise comparative assessment (Karpinska et al., 2021; Chen et al., 2024a; Zheng et al., 2024b), on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
+Pairwise comparative assessment [2, 19, 3], on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
 Except for pointwise evaluation, all forms of LLM-as-a-Judge suffer—or are susceptible to suffering—from position bias due to the intrinsic nature of position and comparison within the prompt structure.
-B.2 Position Bias
-Previous studies have discovered and investigated multiple types of biases, such as position bias (Zheng et al., 2024b; Li et al., 2024; Chen et al., 2024a; 2023; Chia et al., 2023; Chiang & Lee, 2023b; Khan et al., 2024; Zhu et al., 2023b; Zhong et al., 2023; Zheng et al., 2024a; Deb & Das, 2023; Zeng et al., 2023; Wang et al., 2023c; Qin et al., 2024; Wang et al., 2023b; Shi et al., 2024; Liusie et al., 2024; Liu et al., 2024; Kim et al., 2023; Karpinska et al., 2021; Hou et al., 2024; Kocmi & Federmann, 2023b; Pezeshkpour & Hruschka, 2023; Shen et al., 2023; Chiang & Lee, 2023a; Kocmi & Federmann, 2023a; Chen et al., 2024b; Li et al., 2023c; d; Raina et al., 2024), verbosity/length bias (Zheng et al., 2024b; Liu et al., 2024; Kim et al., 2023), self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b), selection bias (Zheng et al., 2024a; Pezeshkpour & Hruschka, 2023), and contextual bias (Liu et al., 2024; Zhou et al., 2024).
-Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source (Li et al., 2023b; Chen et al., 2024c), proprietary commercial ones (OpenAI, 2023; Anthropic, 2024; Gemini Team, 2024), and fine-tuned models (Huang et al., 2024; Wang et al., 2023c; Zhu et al., 2023b; Li et al., 2023b; Kim et al., 2023).
+G.2 Position Bias
+Previous studies have discovered and investigated multiple types of biases, such as position bias [3, 4, 19, 22, 35, 21, 8, 5, 34, 40, 36, 1, 47, 6, 45, 46, 10, 11, 17, 2, 20, 25, 41, 28, 30, 29, 44, 12, 7, 43], verbosity/length bias [3, 11, 17], self-enhancement bias [12, 3], selection bias [40, 41], and contextual bias [11, 48].
+Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source [39, 18], proprietary commercial ones [13, 14, 15], and fine-tuned models [42, 47, 5, 39, 17].
 To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content.
 For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs.
-Some other studies also use the term position bias (Ko et al., 2020; Aslanyan & Porwal, 2019; Blunch, 1984; Raghubir & Valenzuela, 2006), but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
+Some other studies also use the term position bias [49, 50, 51, 52], but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
 Position bias is arguably the most prevalent and impactful bias among all.
-Chua et al.
-(2024) notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought (Wei et al., 2022) performance over self-training controls for all biases except position bias.
+[9] notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought [53] performance over self-training controls for all biases except position bias.
 Furthermore, Khan et al.
-Khan et al.
-(2024) point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
+[8] point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
 Moreover, there is ongoing debate over whether selection bias originates from position bias.
-Pezeshkpour & Hruschka (2023) argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity.
-Raina et al.
-(2024) contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
+[41] argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity.[43] contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
 All of these findings underscore the critical importance of addressing the issue of position bias when employing LLM-as-a-Judge.
-B.3 Deal with Position Bias
-Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts (Zheng et al., 2024b; Raina et al., 2024; Pezeshkpour & Hruschka, 2023; Zhao et al., 2021; Zhu et al., 2023a).
-Also, LLM judges are vulnerable to attacks (Raina et al., 2024; Chen et al., 2024b; Shi et al., 2024).
+G.3 Deal with Position Bias
+Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts [3, 43, 41, 54, 55]. Also, LLM judges are vulnerable to attacks [43, 44, 46].
 Parse data with position bias There are many ways to deal with the position bias.
 The naive way is to exclude the inconsistent judgments(i.e.
-if the LLM judge gives a positionally biased judgment on a pair or list of model responses) (Zheng et al., 2024b; Chen et al., 2024b; Wang et al., 2023b; Li et al., 2023b).
+if the LLM judge gives a positionally biased judgment on a pair or list of model responses) [3, 44, 45, 39].
 While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem.
 Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
-To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging (Zheng et al., 2024b; Raina et al., 2024; Li et al., 2023b; Wang et al., 2023b) or regard inconsistency as a half-win or tie for relation-based judging (Zheng et al., 2024b; Li et al., 2023c) after swapping the order of model responses in the prompt.
+To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging [3, 43, 39, 45] or regard inconsistency as a half-win or tie for relation-based judging [3, 12] after swapping the order of model responses in the prompt.
 For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B.
@@
 To save the expense of running the experiments more than once using swapped order, in practice, many of these studies also support a random-shuffle option in their code settings such that the baseline model for comparison does not remain in a fixed position.
-Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping (Hou et al., 2024), split-and-merge (Li et al., 2023d), and multi-agent discussion (Li et al., 2023c; Khan et al., 2024).
+Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping [20], split-and-merge [7], and multi-agent discussion [12, 8].
 However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
-Furthermore,Liu et al.
-(2024) suggests that existing calibration techniques designed to reduce bias, including context decomposition (Li et al., 2023d), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a), and batch calibration (Zhou et al., 2024), are insufficient to align LLM evaluators, even with supervised data.
+Furthermore,[11] suggests that existing calibration techniques designed to reduce bias, including context decomposition [7], order permutation [45, 40], ensembling [56], and batch calibration [48], are insufficient to align LLM evaluators, even with supervised data.
 Thus, position bias is a pervasive, significantly impactful, and challenging problem to solve.
@@
 In other words, without clarity on the key factors and their quantitative impact on position bias, the efficacy of current and future methods remains uncertain.
-For instance, Li et al.
-(2023d) proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%.
+For instance, [7] proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%.
 However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful.
@@
 We integrate the measurement by formally defining positional fairness score to quantify the extent of a judge model’s preference towards certain positions.
-While prior research includes some positional preference results (Zheng et al., 2024b; Wang et al., 2023b; Zhu et al., 2023b; Li et al., 2023c; Liusie et al., 2024), none have conducted a detailed analysis.
+While prior research includes some positional preference results [3, 45, 5, 12, 10], none have conducted a detailed analysis.
 They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
@@
 Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions.
-MLLM-as-a-Judge (Chen et al., 2024a) conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments.
+MLLM-as-a-Judge [19] conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments.
 In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness.
 If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
-B.4 Summary of Prior Work
+G.4 Summary of Prior Work
 To summarize, LLM-as-a-Judge has a large potential in alternating human judges on a wide range of tasks, in particular subjective evaluation, owing to its cost-effectiveness, high agreement with human judgments, reproducibility, and scalability.
@@
 In a nutshell, a comprehensive understanding of the position bias of LLM-as-a-Judge is crucial to validate existing and future approaches to address this important problem.
-Appendix C LLM Agreement Analysis
+Appendix H LLM Agreement Analysis
 Besides the exploration of position bias with a broad lens by average $PC$ and $PF$, instance-wise agreement between LLM judges is also insightful.
@@
 (2) how do the choices of all judges on an instance vary?
-Figure 5: Disagreement Analysis and Mutual Agreement Heatmap of LLM judges on MTBench and DevBench.
+Figure 5: Disagreement Analysis (Figures (a)(b)) and Mutual Agreement Heatmap (Figures (c)(d)(e)) of LLM judges on MTBench and DevBench.
 The blue bars of Figures (a) and (b) represent the number of instances with certain disagreement values, and the red cumulative curve displays the cumulative count and corresponding percentage over the benchmark dataset.
 The heat maps of Figures (c)(d)(e) marked higher mutual agreement values with a brighter color.
-C.1 Mutual Agreement & Familial Property
-We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations.
-We denote two judges agree on an instance if their judgment choices are identical.
+H.1 Mutual Agreement & Familial Property
+We compute the LLM judges’ mutual agreement on the instances to explore how “alike" or consistent they are across a set of evaluations.
+We denote two judges agree on an instance if their choices are identical.
 Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances.
@@
 5(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively.
-For MTBench that utilizes the 3-option mode, we also consider the “without tie” agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions.
-The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig.
+For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain response, compared to the case when they prefer different responses.
+The "without tie" agreement heatmap of the twelve judges on MTBench is explored in Fig.
 5(d).
-The heatmaps reveal clear “familial patterns” in the judgment choices of these LLM judges.
+The heatmaps reveal clear “familial patterns” in the choices of these LLM judges.
 For instance, the GPT-4, GPT-4-Turbo, and GPT-4o series exhibit high agreement on MTBench, achieving over 70% with ties included and over 85% without.
-GPT-3.5-Turbo didn’t agree with the GPT-4 series and o1-mini for around 40% of the instances, indicating that they are considerably different in judging capabilities.
+GPT-3.5 models didn’t agree with the GPT-4 series and o1-mini for around 40% of the instances, indicating that they are considerably different in judging capabilities.
 For Claude-3 models, similar familial patterns could be observed.
-Claude-3-Opus highly agrees with Claude-3.5-Sonnet, probably due to their similar capabilities, while it also highly agrees with Claude-3-Sonnet, likely due to their similar model structure within the same series.
+Claude-3-Opus highly agrees with Claude-3.5-Sonnet, probably due to their similar capabilities, while it also highly agrees with Claude-3-Opus, likely due to their similar model structure within the same series.
 Interestingly, Claude-3.5-Sonnet and Claude-3-Sonnet do not exhibit a significantly high agreement, indicating that the upgrade from series 3 to 3.5 considerably impacts their judging capabilities.
-Gemini models exhibit rather low mutual agreement and “familial property” is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
+Gemini models exhibit rather low mutual agreement where "familial property" is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
 These patterns suggest that familial similarities, possibly stemming from analogous model sizes, training data, and strategies, influence the positional preferences of these judges.
@@
 Identifying such groupings provides valuable insights, as comparisons between judges from different groups, both adept at assessing LLM-generated content, can reveal distinct position biases and enrich our understanding of this phenomenon.
-C.2 Disagreement & Benchmark Design Insight
+H.2 Disagreement & Benchmark Design Insight
 Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful.
 Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority.
-By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement.
+By this definition, an instance with all judges reaching a consensus on the better response will have a disagreement of 0; an instance where judgments are widely varied will result in a high disagreement.
 For our study where twelve judges are investigated, the maximum disagreement of an MTBench is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode.
 On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
-The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig.
-5(a) and Fig.
-5(b), respectively.
+The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig.5(a) and (b), respectively.
 From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks.
@@
 In other words, majority voting of multiple capable LLM judges could be practically useful for over 99% of evaluation instances on both benchmarks.
-Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig.
-5(a) and Fig.
-5(b) exhibit a balanced distribution of instances with varied difficulty.
+Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig.5(a)(b) exhibit a balanced distribution of instances with varied difficulty.
 This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
@@
 Last but not least, disagreement analysis provides a convenient way to make the difficulty variance of instances varied across the dataset tangible.
-Since the difficulty of an evaluation instance is closely related to the quality gap between the two solutions and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
-Appendix D More Discussions on Answer Quality Gap
+Since the difficulty of an evaluation instance is closely related to the quality gap between the two answers and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
+Appendix I More Discussions on Answer Quality GAP
 Intuitively, answer pairs with larger quality gaps are easier for LLM judges to evaluate, as one answer is clearly superior.
@@
 The quantification utilizes the overall win rate ($owr$) of a candidate model (i.e., answer provider) on the other to compute the quality gap.
-For a set of questions on a certain task and the corresponding solutions by two candidate models, an LLM judge’s evaluations would pick out the candidate winner.
-For the position consistent judgment instances, we can count the number of a candidate’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively.
-On the other hand, following the methodology introduced by Zheng et al.
-(2024b), we consider the inconsistent judgment pairs as a tie for both candidates, whose counting number is denoted as $C_{I}$.
+For a set of questions on a certain task and the corresponding responses answered by two models, an LLM judge’s evaluations would pick out the winner model.
+For the position consistent judgment instances, we can count the number of a model’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively.
+On the other hand, following the methodology introduced by [3], we consider the inconsistent judgment pairs as a tie for both models, whose counting number is denoted as $C_{I}$.
 This follows the intuition that it is difficult to strictly pick out a better one between two responses with similar quality, hence increasing the likelihood of position bias.
-Therefore, considering both cases, we define the overall win rate ($owr$) of a candidate’s solution over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs involving the two candidate solutions in both original and swapped settings.
+Therefore, considering both cases, we define the overall win rate ($owr$) of a model’s response over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of choice pairs involving the two models in both original and swapped settings.
 Then the quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, as 0.5 is the exact tie case for the win rate ranging from 0 to 1.
@@
 In the extreme case, where all judgments are inconsistent, $n_{C}$ is 0 and the consistent win rate is not even computable, making the measurement inaccurate and impractical.
-In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie” and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
-Appendix E More Results of Position Bias and Answer Quality Gap Measurement
-As shown in Fig.
-4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness.
+In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie" and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
+Appendix J More Results of Position Bias and Answer Quality Gap Measurement
+As shown in Fig.4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness.
 In this session, we explore whether the discovery is consistent for each individual judge.
-Same as Section 2.2, we apply the overall win rate to reflect the answer quality gap for visualization.
-Figure 6: Position Consistency and Preference Fairness vs. overall win rate for each judge on MTBench. Figure (a) refers to the relationship investigation of $PC$ and figure (b) for $PF$.
-As shown in Fig.
-6 (a), the “parabolic shape” is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency” applies to all models.
-However, Fig.
-6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges.
-For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape”, GPT-4 and GPT-3.5 present nearly linear curves.
+Same as Section B.2, we apply the overall win rate to reflect the answer quality gap for visualization.
+Figure 6: Positional consistency and positional fairness vs. overall win rate for each judge on MTBench. Figure (a) refers to the relationship investigation of $PC$ and figure (b) for $PF$.
+As shown in Fig.6 (a), the “parabolic shape" is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency" applies to all models.
+However, Fig.6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges.
+For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape", GPT-4 and GPT-3.5 present nearly linear curves.
 In other words, while the former models align with the general tendency that a larger answer quality gap improves positional fairness, the latter ones preserve fairness regardless of the answer quality gap.
 This further demonstrates the necessity to investigate positional fairness in addition to consistency when evaluating a judge model’s position bias.
-Appendix F Variable Selection and Tests
-F.1 Bidirectional Stepwise Regression with AIC
-Bidirectional stepwise regression is a combination of forward selection and backward elimination techniques.
-It iteratively refines the model by adding or removing predictors based on a statistical criterion—commonly the Akaike Information Criterion (AIC).
-The objective is to select a model that balances goodness of fit and complexity, aiming for the lowest AIC value.
-The AIC is given by:
-$\text{AIC}=2k-2\log(L),$ (4)
-where $L$ is the likelihood of the model and $k$ is the number of parameters in the model, including the error variance $\sigma^{2}$.
-For a linear regression model with independent and identically distributed (iid) errors, $N(0,\sigma^{2})$, fitted to $n$ observations, the log-likelihood can be written as:
-$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$ (5)
-where $\hat{e}_{i}$ is the residual for the $i$th observation, and $\sigma^{2}$ is the variance of the errors. The AIC, in this context, becomes:
-$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$ (6)
-This form of the AIC balances the goodness of fit (as reflected by the residual sum of squares) and model complexity (as represented by $k$).
-The operation of Bidirectional stepwise regression starts with either no predictors (forward selection) or all predictors (backward elimination), where the model iteratively adds or removes variables.
-Each step evaluates the impact on the AIC score.
-In forward selection, variables are added one by one, starting from the null model, such that the addition of each variable results in the largest decrease in AIC.
-In backward elimination, all variables are included in the model initially, and variables are removed one at a time, with the variable whose removal causes the smallest increase in AIC being dropped.
-At each iteration, the change in AIC is computed as $\Delta\text{AIC}=\text{AIC}_{\text{new}}-\text{AIC}_{\text{current}}$, where $\text{AIC}_{\text{new}}$ refers to the AIC after adding or removing a variable, and $\text{AIC}_{\text{current}}$ is the AIC of the current model.
-If $\Delta\text{AIC}<0$, the model is improved by the addition or removal of the variable.
-The process terminates when neither adding nor removing variables results in a lower AIC, signifying that the most parsimonious model, based on AIC, has been reached.
-F.2 Test results
-We operated bidirectional stepwise regression on both benchmarks individually and together to identify the factors that are significantly contributing to position bias.
-Specifically, the variables include lengths (input, output, and prompt), answer quality gap, LLM judges, candidate models, and task categories to predict position consistency and preference fairness respectively.
-Table 5, 6 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively.
-Table 7, 8 serves for DevBench, and Table 3, 4 is conducted on the integrated set of both benchmarks.
-The impact of variables on the model is ranked from highest to lowest, from bottom to top.
-Removed variables listed as None indicate the full model at this given step.
-Through benchmark testing, we verified that LLM judges, task categories, and the answer quality gap significantly contribute to position bias in terms of both position consistency and preference fairness.
-These findings align with our empirical results, showing that position bias varies notably by judge and task, with the answer quality gap being a key influencing factor.
-The extent of this impact is reflected by the magnitude of change in AIC when the given variable is removed.
-It is worth noting that while task output length remains a significant predictor for $PF$ and $PC$ in both benchmarks, the change in AIC magnitude after removing this variable is very minimal.
-This is consistent across both benchmarks individually and combined.
-We therefore conclude that, although position bias is influenced by task output length, this dependency is minimal.
-Table 3: Final results of stepwise model selection for both benchmarks: Position Consistency
-Removed Variables DF Sum of Sq RSS AIC
-None 163.75 -18370
-Task 20 2.832 166.59 -18319
-Candidate 38 4.472 168.23 -18303
-Quality gap 1 21.953 185.71 -17703
-Judge 13 55.417 219.17 -16846
-Table 4: Final results of stepwise model selection for both benchmarks: Preference Fairness
-Removed Variables DF Sum of Sq RSS AIC
-None 254.28 -16103
-Task output length 1 0.836 255.12 -16088
-Quality gap 1 11.339 265.62 -15873
-Task 21 16.177 270.46 -15817
-Judge 13 82.069 336.35 -14641
-Table 5: Final results of stepwise model selection for MTBench: Position Consistency
-Removed Variables DF Sum of Sq RSS AIC
-None 61.974 -13312
-Task output length 1 0.0553 62.029 -13311
-Candidate 29 1.6474 63.621 -13282
-Task 7 1.5304 63.504 -13244
-Judge 13 15.3637 77.338 -12594
-Quality gap 1 15.6206 77.594 -12559
-Table 6: Final results of stepwise model selection for MTBench: Preference Fairness
-Removed Variables DF Sum of Sq RSS AIC
-None 129.00 -10909.2
-Quality gap 1 1.931 130.93 -10861.3
-Task 7 9.295 138.29 -10689.4
-Judge 13 58.847 187.85 -9672.5
-Table 7: Final results of stepwise model selection for DevBench: Position Consistency
-Removed Variables DF Sum of Sq RSS AIC
-None 55.382 -6940.2
-Task output length 1 0.257 55.638 -6933.2
-Candidate 9 1.514 56.896 -6905.4
-Quality gap 1 13.128 68.510 -6525.3
-Judge 13 84.760 140.141 -5146.6
-Table 8: Final results of stepwise model selection for DevBench: Preference Fairness
-Removed Variables DF Sum of Sq RSS AIC
-None 60.104 -6753.9
-Task output length 1 0.061 60.165 -6753.9
-Candidate 9 0.731 60.834 -6748.2
-Task 13 1.305 61.408 -6737.8
-Quality gap 1 1.783 61.886 -6698.6
-Judge 13 80.875 140.979 -5108.9
-Appendix G DevBench
+Appendix K DevBench
 This session includes a similar baseline comparison analysis on DevBench as on MTBench.
@@
 Similarly, although GPT-4 stands as the baseline model with a generally high $PC$ across tasks, certain models achieve comparable or superior performances on certain tasks.
-For instance, for architecture design evaluations, GPT-4-Turbo, GPT-4o, and Gemini-1.5-pro all surpass GPT-4.
+For instance, for architecture design evaluations, gpt-4-1106-preview, GPT-4o, and Gemini-1.5-pro all surpass GPT-4.
 Gemini-1.5-pro is especially outstanding, also exceeding GPT-4 in uml class evaluations.
@@
 Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
-Appendix H Experiment Settings
+Appendix L Experiment Settings
 This session specifies more detailed information about the judges, answer-generating models, tasks, and prompt templates used in this study.
 We choose to evaluate MTBench and DevBench for the following reasons: (1) all necessary information about the benchmark models, tasks, and questions is publicly available, making modifications convenient (2) they include a wide variety of answer-generating models, tasks, and task questions for a comprehensive evaluation (3) their human evaluations validated the reliability of state-of-the-art judging models (GPT-4 and GPT-4-Turbo) on their evaluation instances, hence model untested by prior work, if reaching high agreement with these validated judges, can be perceived reliable as well.
-H.1 Judges, Candidates, and Tasks
+L.1 Judges, Models, and Tasks
 Judge
@@
 They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations.
-The other models that are compared to the reference models, namely “Model” in our context, are listed as follows.
+The other models that are compared to the reference models, namely “Model" in our context, are listed as follows.
   • MTBench (30): alpaca-13b, baize-v2-13b, chatglm-6b, claude-instant-v1, claude-v1, dolly-v2-12b, falcon-40b-instruct, fastchat-t5-3b, gpt-3.5-turbo, gpt-4, gpt4all-13b-snoozy, guanaco-33b, guanaco-65b, h2ogpt-oasst-open-llama-13b, koala-13b, llama-13b, mpt-30b-chat, mpt-30b-instruct, mpt-7b-chat, nous-hermes-13b, oasst-sft-4-pythia-12b, oasst-sft-7-llama-30b, palm-2-chat-bison-001, rwkv-4-raven-14b, stablelm-tuned-alpha-7b, tulu-30b, vicuna-33b-v1.3, vicuna-7b-v1.3, wizardlm-13b, wizardlm-30b
   • DevBench (10): codellama-7b-instruct, codellama-13b-instruct, codellama-34b-instruct, deepseek-coder-1.3b-instruct, deepseek-coder-6.7b-instruct, deepseek-coder-33b-instruct, gpt-3.5-turbo-1106, gpt-4-0125-preview, gpt-4-0613, gpt-4-1106-preview
-The model names are exactly what MTBench Zheng et al.
-(2024b) and DevBench Li et al.
-(2024) used in their studies.
+The model names are exactly what MTBench [3] and DevBench [4] used in their studies.
 That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4).
@@
     • architecture design (5): conformance, design_and_coding, practicability, uniformity_and_integration, and faithfulness
-H.2 Prompt Settings
+L.2 Prompt Settings
 We follow the original prompt settings of MTBench and DevBench in our study of pairwise comparative LLM-as-a-Judge.
@@
   • A system prompt explaining the judging task and the role the LLM should be playing.
-  • Emphasized “should” and“shouldn’t”s.
+  • Emphasized “should" and“shouldn’t"s.
   • A prompt structure with placeholders for specific questions and model answers
   • A specified output format for later judgment extraction
-  • Chain-of-Thought Wei et al. (2022) prompts requiring the LLM judge to provide reasons for its judgment
+  • Chain-of-Thought [53] prompts requiring the LLM judge to provide reasons for its judgment
 The detailed prompt templates are specified below.

### v6

@@
-Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative LLM-as-a-Judge
+Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs
 Abstract
-LLM-as-a-Judge offers a promising alternative to human judges across various tasks, yet inherent biases, particularly position bias—a consistent preference for answers based on their position in the prompt—compromise its effectiveness.
-Our study proposes a systematic framework to investigate the position bias of LLM judges in the pairwise comparative scenario in terms of repetition stability, position consistency, and preference fairness.
-In pairwise comparative settings, LLM judges are prompted to choose the better candidate’s answer to the task question, where the task question and a pair of candidate answers are provided.
-Then the prompt with swapped order of candidate answers is given to the same judge again.
-If the judge prefers the same winner, it is denoted position consistent; otherwise, if different winners are selected after swapped settings, position bias is observed.
-To comprehensively understand the position bias of LLM judges, our experiments cover 12 judges across MTBench and DevBench with 22 tasks and around 40 answer-generating models, resulting in over 100,000 evaluation instances.
-Further exploration of the Judge-level, Candidate-level, and Task-level impactful factors on position bias enhances the understanding of this bias.
-Through comprehensive investigations, we validate that the position bias of capable LLM judges is not merely random variations.
-We also observe significant variance of position bias across judges and tasks, highlighting trade-offs of positional consistency, preference fairness, and practicality when selecting task-specific judges.
-Among impactful factors, position bias is largely influenced by the quality gap of given candidate answer pairs for evaluation.
-However, the lengths of the prompt components, including input questions, candidate answers, and the complete prompt, have minimal impact on position bias.
-In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs.
-These insights guide the selection of optimal judge models, enhance benchmark design, and lay the foundation for future research into effective debiasing strategies, ultimately improving the reliability of LLM evaluators for more robust, unbiased evaluation systems.
-Provocations
-Despite advancements in LLM-as-a-Judge and prevalent attention on position bias, the understanding of such bias is still limited and requires a more systematic evaluation.
-Enhancing the understanding of the impactful factors on position bias could inspire more effective mitigation strategies, laying the foundation for more reliable, robust, and unbiased evaluation systems.
-To this end, our study investigates aspects that are underexplored by prior works to offer a systematic evaluation framework for the position bias of LLM judges.
-Repetition Stability
-Validating LLM judges are repetition stable is a primary challenge that is often ignored by previous studies.
-They trusted the one-time generated judgments of the LLMs and conducted further analysis, implicitly assuming that the judgments represent the judge’s capabilities.
-However, the judgment choices and consequently observed position bias may be due to random variations.
-In other words, without verifying that LLM judges are stable in their choices across multiple trials, it is uncertain whether position bias stems from positional shift or is merely randomness.
-To address this gap, our study utilizes repetition stability to validate capable LLMs before further investigations on position bias.
-Preference Fairness
-While position bias is prevalent across most tasks and widely studied by literature, the main focus is on position consistency, as it is the most straightforward measurement of position bias.
-However, the preference directions when position bias occurs are largely underexplored by prior works.
-In this study, we show that previous measurements are either sensitive to the dataset size or overly penalizing highly consistent judges.
-To address this problem, we propose a more granular and scalable measurement that combines the strengths of both methods to assess preference fairness.
-We further stress the trade-off between position consistency, preference fairness, and practicality when choosing the task-specif judges, as position bias largely varies by task and judge.
-Answer Quality Gap
-From intuition, the difficulty of judging a pair of candidate answers (i.e., choosing the winner that better answers the task question) largely depends on the disparities of the answer pairs in quality.
-However, we are the first study to quantify the answer quality gap and investigate its correlation with position bias.
-Our investigations verify that as the answer quality gap enlarges, implying one candidate’s answer is significantly better than the other, the judges become more constant and fair.
-The findings not only benefit the interpretation of other LLM judge biases like verbosity/length bias and self-enhancement bias, but also benefit future benchmark designs for more social goods.
-To summarize, this study proposes new metrics and studies more impacting factors on position bias that are underexplored by prior studies, offering a more comprehensive and systematic framework for understanding and evaluating the position bias of LLM judges.
-We stress the need to have a deeper understanding of the complex bias before moving to solutions.
-More detailed analysis, insights, findings, and contributions can be found in the Appendix.
-[…]
-[appendices]
-Appendix Table of Contents
-[…]
-[appendices]0
-Appendix
-Appendix A Introduction
-In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global attention and stimulating substantial research into their applications.
+LLM-as-a-Judge presents a promising alternative to human evaluators across various tasks, but inherent biases, especially position bias — a tendency to favor solutions based on their position in the prompt — have compromised its effectiveness.
+Our study introduces a systematic framework to examine position bias in pairwise comparisons, focusing on repetition stability, position consistency, and preference fairness.
+This research significantly contributes to the field by introducing new concepts for understanding position bias and providing a multi-dimensional framework for evaluations.
+We conducted experiments with 12 LLM judges across MTBench and DevBench, covering 22 tasks and approximately 40 solution-generating models — candidates, resulting in over 100,000 evaluation instances.
+Our findings confirm that position bias in capable LLM judges is not due to random chances, along with notable variations observed across judges and tasks.
+Moreover, position bias is weakly influenced by the length of prompt components but significantly impacted by the quality gap between solutions.
+These insights can help optimize judge model selections, improve benchmark design, and inform future research on debiasing strategies, ultimately enhancing the reliability of LLM judges.
+1 Introduction
+In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications.
 Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields.
-While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations [1, 2].
+While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations (Zeng et al., 2023; Karpinska et al., 2021).
 To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks.
-Despite a high level of agreement with human judgments[3, 4, 5], inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
+Despite a high level of agreement with human judgments(Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
 Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig.
 1.
-This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations [6], for any score-based and relation-based settings [7].
+This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations (Qin et al., 2024), for any score-based and relation-based settings (Li et al., 2023d).
 Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias.
-However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks [8, 9].
+However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks (Khan et al., 2024; Chua et al., 2024).
 Moreover, the limited understanding of the bias itself has exacerbated the problem.
@@
 Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge.
-When the positions of model responses are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same response.
+When the positions of candidate model solutions are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same candidate model solution.
 We observe repetition instability if the same LLM judge gives different judgments across repetitions.
@@
 Another key aspect of position bias, preference fairness, has also been under-explored by previous works.
-Most studies have primarily focused on positional consistency to measure how frequently position bias occurs across a set of evaluation instances.
+Most studies have primarily focused on position consistency to measure how frequently position bias occurs across a set of evaluation instances.
 However, the direction of biased preference (i.e., primacy or recency) is also insightful.
 Fig.
-2 illustrates cases where a pair of judgments with swapped prompts become positionally consistent, inconsistent, primacy-preferred, or recency-preferred.
+2 illustrates cases where a pair of judgments with swapped prompts become position consistent, inconsistent, primacy-preferred, or recency-preferred.
 We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations.
-Specifically, judgments are considered to have a fair preference if they are positionally consistent or if the preference directions are evenly distributed.
-In contrast, an LLM judge is considered to have an unfair preference if it consistently favors responses that appear either first or last in the prompt.
+Specifically, judgments are considered to have a fair preference if they are position consistent or if the preference directions are evenly distributed.
+In contrast, an LLM judge is considered to have an unfair preference if it consistently favors candidate model solution that appear either first or last in the prompt.
 Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
-Figure 2: Demonstration of positionally consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better response.
+Figure 2: Demonstration of position consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better candidate solution.
 A judgment is consistent if it selects the same winner across different prompt positions.
-Primacy/recency indicates a preference for the first/last model response, respectively.
-Positionally inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
+Primacy/recency indicates a preference for the first/last candidate solution, respectively.
+Position inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
 To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment.
-We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring [3, 10], making it a practically preferable choice for investigations.
+We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring (Zheng et al., 2024b; Liusie et al., 2024), making it a practically preferable choice for investigations.
 Our contribution can be summarized as following: (1) introducing a systematic framework for evaluating LLM-as-a-Judge’s position bias in terms of Repetition Stability, Position Consistency, and Preference Fairness; (2) identifying Judge-level, Candidate-level, and Task-level factors that affect position bias; and (3) providing insights into future benchmark design and paving the way for effective mitigation strategies for position bias.
-Appendix B Methods & Defintions
+To better understand position bias and the effectiveness of LLMs as evaluators, we have developed a detailed framework specifically for analyzing position bias in scenarios where two options are compared directly, known as pairwise comparative assessment.
+We choose this approach because LLMs tend to perform better in these head-to-head comparisons than in other formats like scoring each option separately (Zheng et al., 2024b; Liusie et al., 2024).
+This makes pairwise comparisons a practical focus for our research.
+Our contributions are threefold: (1) We introduce a systematic framework to analyze how LLMs judge by focusing on Repetition Stability, Position Consistency, and Preference Fairness.
+(2) We explore factors at different levels—Judge, Candidate, and Task—that influence position bias.
+(3) We offer insights that will help in designing better benchmarks and developing strategies to reduce position bias in LLM evaluatio
+2 Methods & Defintions
 LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments.
-Our study focuses on relation-based assessment, specifically under the pairwise comparison setting, for several reasons.
-First, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields [6].
-Second, as noted by [11], LLM and human evaluations are more aligned under the context of pairwise comparison than any of the score-based assessments.
-Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of positional consistency[3, 10].
-Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, by some advanced ranking algorithms [6, 11].
+under OuFirst, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields (Qin et al., 2024).
+Second, as noted by Liu et al.
+(2024), LLM and human evaluations are more aligned in the context of pairwise comparison than any of the score-based assessments.
+Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of position consistency(Zheng et al., 2024b; Liusie et al., 2024).
+Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, through some advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024).
 Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
-In the context of pairwise comparative assessments, LLM-judges are prompted to choose the better response that answers the question.
+In the context of pairwise comparative assessments, LLM judges are prompted to choose the better candidate solution to the task question.
 To accommodate the possibility of a tie, various option modes are introduced.
-The Two-Option mode restricts judges to choosing the better response between two options, labeled {A} for the first and {B} for the second as shown in Fig.
+Two-Option mode restricts judges to choosing the better solution between two options, labeled {A} for the first and {B} for the second as shown in Fig.
 1.
-The Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither option is preferable.
-The Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie.
-In our experiments, we employ the Three-Option mode for MTBench [3] and the Two-Option mode for DevBench [4], aligning with their respective original frameworks.
-These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM-judges.
+Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither solution is preferable.
+Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie.
+In our experiments, we employed Three-Option mode for MTBench (Zheng et al., 2024b) and Two-Option mode for DevBench (Li et al., 2024), aligning with their respective original frameworks.
+These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM judges.
 To set up our experiment, as shown in Fig.
-1, the system prompt, option mode, task question, answers/responses from the two candidate models (original prompt) are presented to the LLM-judges to select the winning answer.
+1, the system prompt, option mode, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the winning solution.
 The experiment is conducted in a double-blind setting.
-The identities of the candidate models are hidden from the LLM-judges, and the candidate models are unaware that their responses will be compared to another model when answering the question.
-Then, the prompt with answers in a swapped position (swapped prompt) is given to the same judge again, which results in a choice pair.
+The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question.
+Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair.
 As shown in Fig.
-2, if the choice pair consistently favors the same answer regardless of the swapped position, it is considered positionally consistent with a fair preference.
-Conversely, if the LLM-judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
+2, if the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent with a fair preference.
+Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
 To systematically investigate position bias in pairwise comparative judging scenarios, we evaluate LLM judges using three key metrics: Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$).
@@
 Therefore, for each judge, candidate, and task, we calculate $RS$, $PC$, and $PF$ to thoroughly explore the variance in position bias and its impacting factors.
-B.1 Evaluation Metrics
+2.1 Evaluation Metrics
+We validate the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows.
 Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times.
-It is essential to determine whether the choices of LLM judges, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations.
+It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations.
 We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset.
 This metric is formalized as
-$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,D\}$
+$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,...\}$ (1)
 where $S$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries.
 The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
-Position Consistency ($PC$) quantifies how frequently a judge model selects the same response after the order of responses is swapped.
-It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the model selects the same winner.
+Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is swapped.
+It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the judge model prefers the same winning solution.
 Formally, it is calculated as
-$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}}$
+$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}},V=\{(A,B),(B,A),(C,C),...\}$ (2)
 where $(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})$ denotes the judgment pair for the $j^{th}$ query, $n$ represents the number of prompt pairs, and $V$ is the set of choice pairs that correspond to position consistency.
@@
 This formula provides a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
-Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain response positions.
+Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain solution positions.
 In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency.
-These terms replace the more verbose “preference for the first/second position" used in previous studies, ensuring clarity and generalization for future research.
+These terms replace the more verbose “preference for the first/second position” used in previous studies, ensuring clarity and generalization for future research.
 The examples of such preferences are demonstrated in Fig.
 2.
-Positionally consistent judgments are fair in preference.
-The measurements become more complex for positionally inconsistent cases.
+Position consistent judgments are fair in preference.
+The measurements become more complex for position inconsistent cases.
 Previous studies proposed two common ways to measure the preference fairness.
 A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$).
-The counts are then normalized by the total number of prompt pairs [3, 5].
+The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023b).
 However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
-Alternatively, instead of normalizing over the complete dataset, studies like [12, 10] treat positionally inconsistent evaluation instances independently.
-They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of positionally inconsistent pairs.
+Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently.
+They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs.
 We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$.
-However, this approach is problematic because it overlooks the fact that “positionally consistent judgments are also positionally fair," which can lead to overly penalizing highly consistent LLM-judges.
+However, this approach is problematic because it overlooks the fact that “position consistent judgments are also position fair,” which can lead to overly penalizing highly consistent LLM-judges.
 For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred.
@@
 The $PF$ score is formally calculated by
-$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$
+$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$ (3)
 where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively.
 This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero.
-Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the positionally consistent judgments.
-Another advantage of this approach is that it quantifies positional preference with a single score, rather than the two separate scores used in previous studies.
+Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the position consistent judgments.
+Another advantage of this approach is that it quantifies preference fairness with a single score, rather than the two separate scores used in previous studies.
 The $PF$ score is interpreted as follows:
-$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Positionally Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
+$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
 To summarize, our proposed preference fairness score is superior to prior measurements due to its ability to provide a single, comprehensive score that accounts for all evaluation instances, while maintaining its sensitivity across datasets, even when the number of questions and instances varies for each task.
-B.2 Factors affecting position bias
-Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose responses to the task-specific questions are evaluated by the judges.
-An asterisk (*) highlights the factors with significant impact on position bias that we found through experiments.
-‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s answers to the question, making it both a Candidate-level and Task-level factor.
+2.2 Factors affecting position bias
+Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose solutions to the task-specific questions are evaluated by the judges.
+Factors that have a significant impact on position bias are highlighted with an asterisk (*) by bidirectional stepwise regression tests and marked red by empirical findings.
+‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s solutions to the question, making it both a Candidate-level and Task-level factor.
 ‘Prompt’ encompasses the complete query presented to the judges, including Task Input, Task Output, and system prompts.
@@
 Task Input Length $\times$ $\times$ $\checkmark$
-Task Output Length $\times$ $\checkmark$ $\checkmark$
+Task Output Length $\times$ $\checkmark$* $\checkmark$*
 Prompt Length $\times$ $\checkmark$ $\checkmark$
@@
 By design, our framework allows for the integration and assessment of additional influencing factors.
-Among the influencing factors we identified, we chose “familial property” instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments.
+Among the influencing factors we identified, we chose “familial property” for Judge-level factors instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments.
 The model family grouping reflects similar model sizes and training specifics of the model members.
 The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version.
-However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable" models sharing similar judging capabilities.
+However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable” models sharing similar judging capabilities.
 More details and discussions about the familial property can be found in Appendix Sec.
-H, where we examined the LLM mutual agreements to study the similarities and differences across families.
-Answer quality gap, based on our findings, is a highly impactful yet under-explored factor in prior research.
-We define the quality of a model’s answer by how effectively it addresses the question.
-Consequently, the answer quality gap refers to the disparity in quality between two candidate model answers to the same question and is considered a Candidate-level (i.e., response provider) factor.
-Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs, it would prefer the higher-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
-Depending on option mode, we regard a tie judgment as‘half-win’ for both candidates when calculating its win rate over another.
-Then, the win rates between two candidates on a set of task questions can reflect the answer quality gap between them on the task.
-However, if position bias is exhibited, the winner may be inconsistent when the order of candidate responses is swapped in the query.
-Therefore, we observe the LLM judgments where the absolute winner is consistently chosen in both original and swapped queries (termed “consistent wins”), both responses are deemed equal (termed “consistent ties”), and different winners are selected after candidate responses are swapped in the query.
-We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent choice pairs ($C_{I}$), respectively.
-Following [3], we count inconsistent choice pairs as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
-Inspired by [3], we define the overall win rate ($owr$) of a model’s response over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of choice pairs.
+C, where we examined the LLM mutual agreements to study the similarities and differences across families.
+Answer quality gap, based on our findings, is highly impactful yet under-explored in prior research.
+We define the quality of a candidate’s answer/solution by how effectively it addresses the question.
+Consequently, the answer quality gap refers to the disparity in quality between the solutions from two candidate models to the same question and is considered the Candidate-level (i.e., solution provider) factor.
+Ideally, when a reliable LLM judge is presented with a question and corresponding solution pairs, it would prefer the higher-quality solution, where the corresponding candidate is denoted as the winner selected by the judge.
+Following this assumption, we measure the answer quality gap by the win rates between two candidates on a set of tasks and questions.
+We regard a tie judgment as ‘half-win’ for both candidates when calculating its win rate over another, depending on option mode.
+If position bias occurs, the winner may be inconsistent when the order of candidate solutions is swapped in the query.
+Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen in both the original and swapped queries (termed “consistent wins”), cases where both responses are deemed equal (termed “consistent ties”), and cases where different winners are selected after the solutions are swapped.
+We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment pairs ($C_{I}$), respectively.
+Following Zheng et al.
+(2024b), we count inconsistent judgment pairs also as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
+Inspired by Zheng et al.
+(2024b), we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs.
 Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, where 0.5 is the overall win rate when we only have ties.
-In contrary to using consistent win rate by [3] (calculated as $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of positionally consistent judgment pairs) to quantify $\delta_{q}$, the adoption of overall win rate incorporates all data points and captures the “comparable quality” scenario, where responses in similar quality might lead to positionally biased judgments, a scenario that the consistent win rate might overlook.
+In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
 More discussions can be found in Appendix.
 Sec.
-I.
-Appendix C Experiments
-C.1 Experiment Settings
-In this study, we evaluated twelve up-to-date commercial models from the GPT [13], Claude [14], and Gemini [15] series using our systematic framework.
-Three Llama models [16], with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and well-known capability in reasoning tasks.
-However, due to their limited performance on our tasks, as shown in Table 2, we focus our investigation solely on more powerful closed-source models.
-These models are more widely used and reliable, though we still include open-source models to gain more comprehensive insights.
-We adopted the modified MTBench [3] and DevBench [4] datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks.
+D.
+3 Experiments
+3.1 Experiment Settings
+In this study, we evaluated twelve up-to-date commercial models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework.
+Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks.
+However, due to their limited performance on our tasks, as shown in Table 2, we focused our investigation solely on more powerful closed-source models.
+We included the results of open-source models for completeness.
+We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks.
 We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
 MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task.
-We then paired these candidate models with the baseline candidate for evaluation by the LLM judges.
+We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges.
 The prompt templates we used are identical to those used in the benchmarks.
@@
 Sec.
-L.
+H.
 To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials.
-This resulted in 576 and 432 instances per judge for MTBench and DevBench, respectively.
+This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench.
 The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results.
@@
 In total, approximately 100,000 evaluation instances were analyzed in this study.
-C.2 Empirical Results and Interpretations
+To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks.
+We used variables such as average lengths of input, output, and prompt; answer quality gap; LLM judge series; candidate identities; and task categories to predict $PC$ and $PF$, respectively.
+Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score.
+This process involves both forward selection and backward elimination, with each ”step” testing whether including or excluding a variable improves the model’s AIC value.
+Further details about the process can be found in Appendix.
+Sec.
+F.
+3.2 Empirical Results and Interpretations
 The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2.
@@
 A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not merely random variations.
-This further necessitates the need for investigating and analyzing preference fairness simultaneously with position consistency.
 Fig.
@@
 7 demonstrate that position bias varies by judges and tasks significantly.
-Together with Table 2, it stresses the trade-off between $PC$ and $PF$ when selecting task-specific LLM judges.
 Fig.
 4 further investigates the impact of answer quality gap and lengths of prompt components on position bias.
-These analyses are conducted by considering all judges together on MTBench.
-The detailed analyses for each judge and on DevBench can be accessed from Appendix.Sec.J, and Sec.K.
-Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Positional Fairness ($PF$).
-Top 5 performances are marked in bold.“Error" is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format.
-High error rates and low $RS$ are marked red, implying further evaluations (i.e., $RS$, $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data.
+These analyses were conducted by considering all judges together on MTBench.
+The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E, and Sec.
+G.
+Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables.
+Similarly, these factors also contribute significantly to Preference Fairness.
+Additionally, we found that average output length is a statistically significant predictor of $PF$.
+This finding is not surprising, as longer outputs are generally perceived as higher quality and more preferred.
+Quantitative results and more discussions can be found in Appendix.
+Sec.
+F.
+Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$).
+Top 5 performances are marked in bold.
+“Error” is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format.
+High error rates and low $RS$ are marked red, implying further evaluations (i.e., $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data.
 Therefore, open-source models (e.g., Llama) are not investigated in the paper because of their high error rates under the difficulty and length of our evaluation instances.
 In contrast, up-to-date commercial models generally demonstrated minimal error and high repetition stability, making them more practical and meaningful for investigations.
-Judge MTBench DevBench
+Chatbot Arean Leaderboard’s overall ranking of the models is listed.
+Judge MTBench DevBench Arena
 $RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error
-Caude-3.5-sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00
-Claude-3-opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00
-Claude-3-sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00
-Claude-3-haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00
-Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13
-Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96
-Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
-o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00
-GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00
-GPT-4-turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
-GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
-GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00
-Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
-Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
-Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
+Caude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 7
+Claude-3-Opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00 17
+Claude-3-Sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00 37
+Claude-3-Haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00 46
+Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 12
+Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96 25
+Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00 66
+o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00 2
+GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 6
+GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00 17
+GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00 52
+GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 94
+Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 99
+Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 109
+Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94 57
 Figure 3: Baseline comparisons of judges across tasks on MTBench.
@@
 Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$.
-The baseline model is chosen to be gpt-4-0613 for its generally superior performances.
+The baseline model is chosen to be GPT-4 for its generally superior performances.
 Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
@@
 overall win rate, hence the answer quality gap, respectively.
-Generally, judgments become more consistent and fair as the quality gap decreases.
+Generally, judgments become more consistent and fair as the quality gap increases.
 Figure (c) demonstrates the relationship between $PC$ and $PF$, where higher consistency is likely to indicate better fairness.
@@
 The irregular and inconclusive patterns imply that position bias is not significantly impacted by the length of prompts and their components.
-Appendix D Main Findings
-Position Bias of Capable Judges are not Mere Random Variations
-As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal “Error" rates, generally exhibit $RS$ of over 0.85.
-The up-to-date and most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, Gpt-4, and GPT-4o, all achieve near-perfect $RS$ of over 0.95 on both benchmarks.
-These results validate that LLM judgments and the consequently observed position bias are not merely random variations.
-This increases the confidence that one-time generated judgments by these validated LLMs adequately represent their judging capabilities.
-Only with this prerequisite, can we ensure further investigations are reasonable and meaningful.
-Position Bias Varies by Judge & Task
-As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig.
+4 Main Findings
+Position Bias of Capable Judges are not Mere Random Variations As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal ”Error” rates, generally exhibit $RS$ values above 0.85.
+The most up-to-date and capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks.
+These results confirm that LLM judgments, and the resulting position bias, are not merely random variations.
+This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
+Position Bias Varies by Judge & Task As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig.
 3 for $PC$ comparisons.
@@
 Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it may be employed as a cost-effective alternative to coding evaluations.
-Similarly, significant variances of $PF$ across tasks are observed.
-The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others.
+Additionally, significant variations of $PF$ across tasks are observed.
+The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on the others.
 Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, role play, and writing tasks.
 Even for uniformly recency-preferred judges such as Claude-3’s and Gemini-pro’s, the extent of biased preference, as reflected by $PF$ values, varies by task.
-Moreover, high position consistency does not guarantee fairness, weighing the trade-offs between position consistency and preference fairness.
+Moreover, high position consistency does not guarantee fairness.
 For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively.
 In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
-In this sense, sacrificing consistency for more fairness or vice versa can all be practically useful depending on specific needs.
-This trade-off between consistency and fairness is particularly crucial when the most consistent model is not the fairest.
-Therefore, to sum up, the position bias of LLM judges is task-dependent and there is usually a trade-off between position consistency and preference fairness.
+Therefore, the position bias of LLM judges is judge-dependent and task-dependent.
+This observation is confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$.
 In practice, to achieve the most reliable evaluation results, one needs to select task-specific LLMs whose consistency and fairness are in balance.
-The same discoveries can be concluded from DevBench results, as detailed in Appendix.Sec.
-K.
-Position Bias Correlates to Answer Quality Gap
-Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
-In this study, as defined in Section B.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$).
+Moreover, as shown in Table 2, a high ranking in the overall reasoning performance of a model does not guarantee its judging reliability.
+Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec.
+G.
+Position Bias Correlates to Answer Quality Gap Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
+In this study, as defined in Section 2.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$).
 Since an absolute tie has 0.5 $owr$, $\delta_{p}$ increases as $owr$ extends from 0.5 to 0 or 1.
-Fig.4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$.
+Fig.
+4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$.
 This aligns with our intuition that the answer pairs with larger quality disparities are easier to achieve consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$.
-The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.J.
+The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.
+E.
 Similarly, as shown in Fig.
-1(b), judgments generally become more positionally fair as $\delta_{q}$ increases.
+1(b), judgments generally become more preference fair as $\delta_{q}$ increases.
 However, the extent is not as significant as for $PC$.
-Also, the relationship varies by judge, where some LLMs maintain preference fairness regardless of $\delta_{q}$.
-For example, as shown in Fig.6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
-These observations align with the right-arrow shape as demonstrated in Fig.4(c), where there is a general trend that judgments become preference fairer as position consistency increases.
-It also justifies the reasonability of our quantification of positional fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
+Also, the relationship varies by judge, as some LLMs maintain preference fairness regardless of $\delta_{q}$.
+For example, as shown in Fig.
+6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
+These observations align with the right-arrow shape as demonstrated in Fig.
+4(c), where there is a general trend that judgments become preference fairer as position consistency increases.
+It also justifies the reasonableness of our quantification of preference fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
 Together, we conclude that as the answer quality gap enlarges, judges generally become more position consistent and preference fair according to the regression curves.
@@
 Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
-Our investigation also helps to explain the observed verbosity/length bias [11, 17, 3] and self-enhancement bias [12, 3] of LLM judges.
+Our investigation also helps to explain the observed verbosity/length bias (Liu et al., 2024; Kim et al., 2023; Zheng et al., 2024b) and self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b) of LLM judges.
 From our analysis, LLMs may prefer longer answers (verbosity/length bias) since more capable models usually generate longer solutions with higher qualities.
-Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the answers generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4).
+Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the solutions generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4).
 In both cases, the observed biased preference may be fundamentally due to a prominent answer quality gap instead of length or self-identity.
-Position Bias is Length-Independent
-We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the response length of answer-generating models (task output length), and the length of the entire prompt (prompt length).
-Since the tokenization methods vary by model, we apply the length of the prompt component strings for uniformity and scalability.
+Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length).
+Since the tokenization methods vary by models, we apply the length of the prompt component strings for uniformity and scalability.
 The analysis from Fig.
-4(d)(e)(f) shows that there is no significant and definite relationship between the lengths of prompt components and position bias.
+4(d)(e)(f) shows that there is very weak relationship between the lengths of prompt components and position bias.
+By stepwise regression, the average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table.
+4.
 Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
-Practical Implications for Benchmark Design
-We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appedix.Sec.
-H to gain more insights.
-Particularly, the disagreement analysis helps to identify the instances where it is easy/difficult for all judges to reach a consensus, reflecting the difficulty of judging them.
-Based on discoveries of the answer quality gap and LLM agreement/disagreement, our study offers practical implications for evaluator benchmark designs in controlling the difficulty of judging tasks.
-From our findings, the hard-to-judge instances are the ones where (1) LLMs mostly disagree with each other (2) comparable win rates and minimal quality gaps are observed for the two candidate models (3) prominent position bias exhibits for a majority of judges.
-Here is a qualitative analysis of a typical difficult-to-judge example.
-This instance (MTBench question 120) involved a math problem where both models, vicuna-13b-v1.3 and wizardlm-30b, provided incorrect answers.
-The problem statement is as follows: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find x such that $f(x)=0$.
-Methodologically, vicuna-13b-v1.3 erroneously applied the quadratic formula to solve the cubic equation, whereas wizardlm-30b employed an enumeration method but failed to identify the correct answer.
-This resulted in a 3{A}-4{B}-5{C} choice pattern among the twelve judges, a typical case of significant disagreement.
-Qualitatively, We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer.
-The remaining ones showed preferences either for correctness (which then chose {C} as a tie) or methodological soundness (which then chose {B}, i.e., preferring wizardlm-30b).
-Notably, in the swapped order setting of the same problem, 8 of the 12 judges encountered position bias.
-This is a typical illustration of the two responses having a minimal quality gap, leading to difficulty in judging, the likelihood of position bias, and prominent subjectivity during evaluations.
-The takeaway message from examples like this is that the number of hard-to-evaluate and highly subjective instances could be reduced for better benchmark designs.
-Similarly, the trivial-to-judge instances where all judges easily reach a consensus could be reduced among the benchmark population.
-One could also specify more detailed and definite evaluation criteria to erase the ambiguity of some hard-to-evaluate instances.
-Appendix E Conclusion
-In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of pairwise comparative LLM-as-a-Judge in terms of repetition stability, position consistency, and preference fairness.
-We focus on position bias for its prevalent occurrences and significant impact on the reliability of LLM judges’ evaluations.
-From prior works, the pairwise comparative assessment offers the most consistent and reliable judgments compared to other types of LLM-as-a-Judge and hence becomes our concentration for practicality.
-The study starts by investigating the repetition stability of LLM judges to validate that the judgments and consequently observed position bias are not mere random variations.
-Then, in addition to the traditional measurement of positional consistency, we propose a systematic way to quantify the preference fairness of LLM judges.
-Our comprehensive evaluation of twelve judge models on two benchmarks across 22 tasks reveals significant variance of position bias by judge and task.
-This stresses the trade-offs between position consistency, preference fairness, and cost-effective practicality when choosing the optimal judges for specific tasks.
-The study on the Judge-level, Candidate-level, and Task-level impactful factors on position bias further enhances the understanding of this bias.
-For example, we are the first the make a direct connection between answer quality gap and position bias, where our discovery reveals that the quality gap between candidate responses to the task question significantly influences the positional consistency and preference fairness of LLM judges.
-However, position bias is length-independent, meaning that the lengths of prompt components have minimal impact on position bias.
-In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs.
-To summarize, insights from this study enhance the understanding of position bias and lay the groundwork for more robust, unbiased evaluation systems.
-Appendix F Contribution & Limitation
+Practical Implications for Benchmark Design We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appendix.Sec.
+C to gain further insights.
+In particular, the disagreement analysis helps identify instances where it is easy or difficult for judges to reach a consensus, reflecting the complexity of the judgment task.
+Based on our findings regarding the answer quality gap and LLM agreement/disagreement, this study offers practical implications for designing evaluator benchmarks that control the difficulty of judging tasks.
+Our findings show that the hardest-to-judge instances are those where (1) LLMs frequently disagree with each other, (2) comparable win rates and minimal quality gaps exist between the two candidate models, and (3) significant position bias is exhibited by the majority of judges.
+One qualitative analysis example of a difficult-to-judge instance comes from MTBench question 120, which involved a math problem where both candidate models, Vicuna-13b-v1.3 and WizardLM-30b, provided incorrect solutions.
+The problem was: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find $x$ such that $f(x)=0$.
+Vicuna-13b-v1.3 incorrectly applied the quadratic formula to solve the cubic equation, while WizardLM-30b used an enumeration method but failed to identify the correct answer.
+This resulted in a 4{A}-3{B}-5{C} choice pattern among the twelve judges, illustrating a typical case of significant disagreement.
+We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer.
+Others preferred either correctness (choosing {C} as a tie) or methodological soundness (choosing {B}, favoring WizardLM-30b).
+In the swapped order setting, 9 out of 12 judges displayed position bias.
+This highlights the challenge of judging solutions with minimal quality gaps, leading to greater position bias and subjectivity.
+The key takeaway is that reducing hard-to-evaluate, highly subjective instances can improve benchmark design.
+Likewise, minimizing trivial cases where judges easily agree would enhance the benchmark’s robustness.
+Clearer and more specific evaluation criteria could also help resolve ambiguity in difficult instances.
+5 Conclusion
+In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of LLM judges in pairwise comparative scenarios in terms of repetition stability, position consistency, and preference fairness.
+Through comprehensive evaluations of 12 judges on two benchmarks across 22 tasks with over 100,000 evaluation instances, we observe significant variations of position bias across judges and tasks.
+We also discover that position bias is weakly length-dependent but significantly impacted by the answer quality gap.
+Our findings enhance the understanding of position bias, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
+Appendix Table of Contents
+[…]
+[appendices]0
+Appendix
+Appendix A Contribution & Limitation
 Contributions
 Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge.
-Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through positional consistency and positional fairness.
+Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through position consistency and preference fairness.
 We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components.
@@
 First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation.
-Our focus and investigation on positional fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability.
-Moreover, our emphasized trade-offs between positional consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work.
+Our focus and investigation on preference fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability.
+Moreover, our emphasized trade-offs between position consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work.
 Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results.
-Besides, the LLM agreement/disagreement analysis in Appendix.5 provides an applicable tool to identify the difficulty of judging certain instances.
+Besides, the LLM agreement/disagreement analysis in Appendix.Sec.C 5 provides an applicable tool to identify the difficulty of judging certain instances.
 Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias.
-Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare [18], instruction following and prompting [1, 9], multimodal assessment [19], and recommender systems [20].
+Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare (Chen et al., 2024c), instruction following and prompting (Zeng et al., 2023; Chua et al., 2024), multimodal assessment (Chen et al., 2024a), and recommender systems (Hou et al., 2024).
 Strength
@@
 It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works.
-(2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, positional consistency, and positional fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias.
+(2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, position consistency, and preference fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias.
 (3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
@@
 Then our framework offers better control on the other impactful factors such as the answer quality gap to make these further investigations more accurate.
-Appendix G Related Work
-G.1 LLM-as-a-Judge
+Appendix B Related Work
+B.1 LLM-as-a-Judge
 In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications.
-For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature [1].
-However, it lacks scalability and reproducibility [2].
-As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks [3, 4, 21, 10, 22, 23, 24, 25, 26, 11, 1, 27, 28, 29, 30, 2, 31, 17, 32, 33, 34, 35, 36, 37, 38, 7, 12, 39], including open-ended story generation [30], adversarial attacks [21], summarization [2], machine translation [29], and instruction following [1].
-These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles [3, 40, 41, 21, 28, 10, 6, 22, 42, 5, 43, 44, 45, 11, 1, 8, 17, 39, 7, 46, 12, 18, 9, 19].
+For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature (Zeng et al., 2023).
+However, it lacks scalability and reproducibility (Karpinska et al., 2021).
+As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks (Zheng et al., 2024b; Li et al., 2024; Chiang & Lee, 2023b; Liusie et al., 2024; Chen et al., 2023; Fu et al., 2023; Wang et al., 2023a; Kocmi & Federmann, 2023b; Ji et al., 2023; Liu et al., 2024; Zeng et al., 2023; Liu et al., 2023; Shen et al., 2023; Kocmi & Federmann, 2023a; Chiang & Lee, 2023a; Karpinska et al., 2021; Dubois et al., 2024; Kim et al., 2023; Liang et al., 2023; Chang et al., 2023; Zhong et al., 2023; Chia et al., 2023; Deb & Das, 2023; Ye et al., 2024; Min et al., 2023; Li et al., 2023d; c; b; Lu et al., 2024), including open-ended story generation (Chiang & Lee, 2023a), adversarial attacks (Chiang & Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi & Federmann, 2023a), and instruction following (Zeng et al., 2023).
+These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles (Zheng et al., 2024b; a; Pezeshkpour & Hruschka, 2023; Chiang & Lee, 2023b; Shen et al., 2023; Liusie et al., 2024; Qin et al., 2024; Chen et al., 2023; Huang et al., 2024; Zhu et al., 2023b; Raina et al., 2024; Chen et al., 2024b; Wang et al., 2023b; Liu et al., 2024; Zeng et al., 2023; Khan et al., 2024; Kim et al., 2023; Li et al., 2023b; d; Shi et al., 2024; Li et al., 2023c; Chen et al., 2024c; Chua et al., 2024; Chen et al., 2024a).
 As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation.
@@
 LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
-In many of these tasks, LLM judges have shown a high level of agreement with human evaluators [3, 4, 5], yet in some tasks, they are less effective, largely due to inherent biases [28, 21, 40].
+In many of these tasks, LLM judges have shown a high level of agreement with human evaluators (Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), yet in some tasks, they are less effective, largely due to inherent biases (Shen et al., 2023; Chiang & Lee, 2023b; Zheng et al., 2024a).
 Even in cases where agreement is high, judgments may still suffer from biases.
 When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method.
-From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment [6].
-By comparative method, it can be either score-based or relation-based [7].
-For example, pointwise scoring [17] lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics.
-Pairwise/Listwise Scoring [2, 5, 19, 3] prompts the LLM judges to score a pair/list of model-generated answers.
+From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment (Qin et al., 2024).
+By comparative method, it can be either score-based or relation-based (Li et al., 2023d).
+For example, pointwise scoring (Kim et al., 2023) lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics.
+Pairwise/Listwise Scoring (Karpinska et al., 2021; Zhu et al., 2023b; Chen et al., 2024a; Zheng et al., 2024b) prompts the LLM judges to score a pair/list of model-generated answers.
 Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst).
-Pairwise comparative assessment [2, 19, 3], on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
+Pairwise comparative assessment (Karpinska et al., 2021; Chen et al., 2024a; Zheng et al., 2024b), on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
 Except for pointwise evaluation, all forms of LLM-as-a-Judge suffer—or are susceptible to suffering—from position bias due to the intrinsic nature of position and comparison within the prompt structure.
-G.2 Position Bias
-Previous studies have discovered and investigated multiple types of biases, such as position bias [3, 4, 19, 22, 35, 21, 8, 5, 34, 40, 36, 1, 47, 6, 45, 46, 10, 11, 17, 2, 20, 25, 41, 28, 30, 29, 44, 12, 7, 43], verbosity/length bias [3, 11, 17], self-enhancement bias [12, 3], selection bias [40, 41], and contextual bias [11, 48].
-Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source [39, 18], proprietary commercial ones [13, 14, 15], and fine-tuned models [42, 47, 5, 39, 17].
+B.2 Position Bias
+Previous studies have discovered and investigated multiple types of biases, such as position bias (Zheng et al., 2024b; Li et al., 2024; Chen et al., 2024a; 2023; Chia et al., 2023; Chiang & Lee, 2023b; Khan et al., 2024; Zhu et al., 2023b; Zhong et al., 2023; Zheng et al., 2024a; Deb & Das, 2023; Zeng et al., 2023; Wang et al., 2023c; Qin et al., 2024; Wang et al., 2023b; Shi et al., 2024; Liusie et al., 2024; Liu et al., 2024; Kim et al., 2023; Karpinska et al., 2021; Hou et al., 2024; Kocmi & Federmann, 2023b; Pezeshkpour & Hruschka, 2023; Shen et al., 2023; Chiang & Lee, 2023a; Kocmi & Federmann, 2023a; Chen et al., 2024b; Li et al., 2023c; d; Raina et al., 2024), verbosity/length bias (Zheng et al., 2024b; Liu et al., 2024; Kim et al., 2023), self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b), selection bias (Zheng et al., 2024a; Pezeshkpour & Hruschka, 2023), and contextual bias (Liu et al., 2024; Zhou et al., 2024).
+Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source (Li et al., 2023b; Chen et al., 2024c), proprietary commercial ones (OpenAI, 2023; Anthropic, 2024; Gemini Team, 2024), and fine-tuned models (Huang et al., 2024; Wang et al., 2023c; Zhu et al., 2023b; Li et al., 2023b; Kim et al., 2023).
 To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content.
 For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs.
-Some other studies also use the term position bias [49, 50, 51, 52], but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
+Some other studies also use the term position bias (Ko et al., 2020; Aslanyan & Porwal, 2019; Blunch, 1984; Raghubir & Valenzuela, 2006), but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
 Position bias is arguably the most prevalent and impactful bias among all.
-[9] notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought [53] performance over self-training controls for all biases except position bias.
+Chua et al.
+(2024) notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought (Wei et al., 2022) performance over self-training controls for all biases except position bias.
 Furthermore, Khan et al.
-[8] point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
+Khan et al.
+(2024) point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
 Moreover, there is ongoing debate over whether selection bias originates from position bias.
-[41] argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity.[43] contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
+Pezeshkpour & Hruschka (2023) argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity.
+Raina et al.
+(2024) contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
 All of these findings underscore the critical importance of addressing the issue of position bias when employing LLM-as-a-Judge.
-G.3 Deal with Position Bias
-Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts [3, 43, 41, 54, 55]. Also, LLM judges are vulnerable to attacks [43, 44, 46].
+B.3 Deal with Position Bias
+Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts (Zheng et al., 2024b; Raina et al., 2024; Pezeshkpour & Hruschka, 2023; Zhao et al., 2021; Zhu et al., 2023a).
+Also, LLM judges are vulnerable to attacks (Raina et al., 2024; Chen et al., 2024b; Shi et al., 2024).
 Parse data with position bias There are many ways to deal with the position bias.
 The naive way is to exclude the inconsistent judgments(i.e.
-if the LLM judge gives a positionally biased judgment on a pair or list of model responses) [3, 44, 45, 39].
+if the LLM judge gives a positionally biased judgment on a pair or list of model responses) (Zheng et al., 2024b; Chen et al., 2024b; Wang et al., 2023b; Li et al., 2023b).
 While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem.
 Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
-To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging [3, 43, 39, 45] or regard inconsistency as a half-win or tie for relation-based judging [3, 12] after swapping the order of model responses in the prompt.
+To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging (Zheng et al., 2024b; Raina et al., 2024; Li et al., 2023b; Wang et al., 2023b) or regard inconsistency as a half-win or tie for relation-based judging (Zheng et al., 2024b; Li et al., 2023c) after swapping the order of model responses in the prompt.
 For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B.
@@
 To save the expense of running the experiments more than once using swapped order, in practice, many of these studies also support a random-shuffle option in their code settings such that the baseline model for comparison does not remain in a fixed position.
-Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping [20], split-and-merge [7], and multi-agent discussion [12, 8].
+Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping (Hou et al., 2024), split-and-merge (Li et al., 2023d), and multi-agent discussion (Li et al., 2023c; Khan et al., 2024).
 However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
-Furthermore,[11] suggests that existing calibration techniques designed to reduce bias, including context decomposition [7], order permutation [45, 40], ensembling [56], and batch calibration [48], are insufficient to align LLM evaluators, even with supervised data.
+Furthermore,Liu et al.
+(2024) suggests that existing calibration techniques designed to reduce bias, including context decomposition (Li et al., 2023d), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a), and batch calibration (Zhou et al., 2024), are insufficient to align LLM evaluators, even with supervised data.
 Thus, position bias is a pervasive, significantly impactful, and challenging problem to solve.
@@
 In other words, without clarity on the key factors and their quantitative impact on position bias, the efficacy of current and future methods remains uncertain.
-For instance, [7] proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%.
+For instance, Li et al.
+(2023d) proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%.
 However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful.
@@
 We integrate the measurement by formally defining positional fairness score to quantify the extent of a judge model’s preference towards certain positions.
-While prior research includes some positional preference results [3, 45, 5, 12, 10], none have conducted a detailed analysis.
+While prior research includes some positional preference results (Zheng et al., 2024b; Wang et al., 2023b; Zhu et al., 2023b; Li et al., 2023c; Liusie et al., 2024), none have conducted a detailed analysis.
 They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
@@
 Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions.
-MLLM-as-a-Judge [19] conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments.
+MLLM-as-a-Judge (Chen et al., 2024a) conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments.
 In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness.
 If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
-G.4 Summary of Prior Work
+B.4 Summary of Prior Work
 To summarize, LLM-as-a-Judge has a large potential in alternating human judges on a wide range of tasks, in particular subjective evaluation, owing to its cost-effectiveness, high agreement with human judgments, reproducibility, and scalability.
@@
 In a nutshell, a comprehensive understanding of the position bias of LLM-as-a-Judge is crucial to validate existing and future approaches to address this important problem.
-Appendix H LLM Agreement Analysis
+Appendix C LLM Agreement Analysis
 Besides the exploration of position bias with a broad lens by average $PC$ and $PF$, instance-wise agreement between LLM judges is also insightful.
@@
 (2) how do the choices of all judges on an instance vary?
-Figure 5: Disagreement Analysis (Figures (a)(b)) and Mutual Agreement Heatmap (Figures (c)(d)(e)) of LLM judges on MTBench and DevBench.
+Figure 5: Disagreement Analysis and Mutual Agreement Heatmap of LLM judges on MTBench and DevBench.
 The blue bars of Figures (a) and (b) represent the number of instances with certain disagreement values, and the red cumulative curve displays the cumulative count and corresponding percentage over the benchmark dataset.
 The heat maps of Figures (c)(d)(e) marked higher mutual agreement values with a brighter color.
-H.1 Mutual Agreement & Familial Property
-We compute the LLM judges’ mutual agreement on the instances to explore how “alike" or consistent they are across a set of evaluations.
-We denote two judges agree on an instance if their choices are identical.
+C.1 Mutual Agreement & Familial Property
+We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations.
+We denote two judges agree on an instance if their judgment choices are identical.
 Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances.
@@
 5(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively.
-For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain response, compared to the case when they prefer different responses.
-The "without tie" agreement heatmap of the twelve judges on MTBench is explored in Fig.
+For MTBench that utilizes the 3-option mode, we also consider the “without tie” agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions.
+The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig.
 5(d).
-The heatmaps reveal clear “familial patterns” in the choices of these LLM judges.
+The heatmaps reveal clear “familial patterns” in the judgment choices of these LLM judges.
 For instance, the GPT-4, GPT-4-Turbo, and GPT-4o series exhibit high agreement on MTBench, achieving over 70% with ties included and over 85% without.
-GPT-3.5 models didn’t agree with the GPT-4 series and o1-mini for around 40% of the instances, indicating that they are considerably different in judging capabilities.
+GPT-3.5-Turbo didn’t agree with the GPT-4 series and o1-mini for around 40% of the instances, indicating that they are considerably different in judging capabilities.
 For Claude-3 models, similar familial patterns could be observed.
-Claude-3-Opus highly agrees with Claude-3.5-Sonnet, probably due to their similar capabilities, while it also highly agrees with Claude-3-Opus, likely due to their similar model structure within the same series.
+Claude-3-Opus highly agrees with Claude-3.5-Sonnet, probably due to their similar capabilities, while it also highly agrees with Claude-3-Sonnet, likely due to their similar model structure within the same series.
 Interestingly, Claude-3.5-Sonnet and Claude-3-Sonnet do not exhibit a significantly high agreement, indicating that the upgrade from series 3 to 3.5 considerably impacts their judging capabilities.
-Gemini models exhibit rather low mutual agreement where "familial property" is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
+Gemini models exhibit rather low mutual agreement and “familial property” is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
 These patterns suggest that familial similarities, possibly stemming from analogous model sizes, training data, and strategies, influence the positional preferences of these judges.
@@
 Identifying such groupings provides valuable insights, as comparisons between judges from different groups, both adept at assessing LLM-generated content, can reveal distinct position biases and enrich our understanding of this phenomenon.
-H.2 Disagreement & Benchmark Design Insight
+C.2 Disagreement & Benchmark Design Insight
 Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful.
 Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority.
-By this definition, an instance with all judges reaching a consensus on the better response will have a disagreement of 0; an instance where judgments are widely varied will result in a high disagreement.
+By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement.
 For our study where twelve judges are investigated, the maximum disagreement of an MTBench is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode.
 On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
-The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig.5(a) and (b), respectively.
+The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig.
+5(a) and Fig.
+5(b), respectively.
 From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks.
@@
 In other words, majority voting of multiple capable LLM judges could be practically useful for over 99% of evaluation instances on both benchmarks.
-Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig.5(a)(b) exhibit a balanced distribution of instances with varied difficulty.
+Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig.
+5(a) and Fig.
+5(b) exhibit a balanced distribution of instances with varied difficulty.
 This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
@@
 Last but not least, disagreement analysis provides a convenient way to make the difficulty variance of instances varied across the dataset tangible.
-Since the difficulty of an evaluation instance is closely related to the quality gap between the two answers and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
-Appendix I More Discussions on Answer Quality GAP
+Since the difficulty of an evaluation instance is closely related to the quality gap between the two solutions and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
+Appendix D More Discussions on Answer Quality Gap
 Intuitively, answer pairs with larger quality gaps are easier for LLM judges to evaluate, as one answer is clearly superior.
@@
 The quantification utilizes the overall win rate ($owr$) of a candidate model (i.e., answer provider) on the other to compute the quality gap.
-For a set of questions on a certain task and the corresponding responses answered by two models, an LLM judge’s evaluations would pick out the winner model.
-For the position consistent judgment instances, we can count the number of a model’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively.
-On the other hand, following the methodology introduced by [3], we consider the inconsistent judgment pairs as a tie for both models, whose counting number is denoted as $C_{I}$.
+For a set of questions on a certain task and the corresponding solutions by two candidate models, an LLM judge’s evaluations would pick out the candidate winner.
+For the position consistent judgment instances, we can count the number of a candidate’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively.
+On the other hand, following the methodology introduced by Zheng et al.
+(2024b), we consider the inconsistent judgment pairs as a tie for both candidates, whose counting number is denoted as $C_{I}$.
 This follows the intuition that it is difficult to strictly pick out a better one between two responses with similar quality, hence increasing the likelihood of position bias.
-Therefore, considering both cases, we define the overall win rate ($owr$) of a model’s response over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of choice pairs involving the two models in both original and swapped settings.
+Therefore, considering both cases, we define the overall win rate ($owr$) of a candidate’s solution over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs involving the two candidate solutions in both original and swapped settings.
 Then the quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, as 0.5 is the exact tie case for the win rate ranging from 0 to 1.
@@
 In the extreme case, where all judgments are inconsistent, $n_{C}$ is 0 and the consistent win rate is not even computable, making the measurement inaccurate and impractical.
-In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie" and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
-Appendix J More Results of Position Bias and Answer Quality Gap Measurement
-As shown in Fig.4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness.
+In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie” and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
+Appendix E More Results of Position Bias and Answer Quality Gap Measurement
+As shown in Fig.
+4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness.
 In this session, we explore whether the discovery is consistent for each individual judge.
-Same as Section B.2, we apply the overall win rate to reflect the answer quality gap for visualization.
-Figure 6: Positional consistency and positional fairness vs. overall win rate for each judge on MTBench. Figure (a) refers to the relationship investigation of $PC$ and figure (b) for $PF$.
-As shown in Fig.6 (a), the “parabolic shape" is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency" applies to all models.
-However, Fig.6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges.
-For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape", GPT-4 and GPT-3.5 present nearly linear curves.
+Same as Section 2.2, we apply the overall win rate to reflect the answer quality gap for visualization.
+Figure 6: Position Consistency and Preference Fairness vs. overall win rate for each judge on MTBench. Figure (a) refers to the relationship investigation of $PC$ and figure (b) for $PF$.
+As shown in Fig.
+6 (a), the “parabolic shape” is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency” applies to all models.
+However, Fig.
+6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges.
+For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape”, GPT-4 and GPT-3.5 present nearly linear curves.
 In other words, while the former models align with the general tendency that a larger answer quality gap improves positional fairness, the latter ones preserve fairness regardless of the answer quality gap.
 This further demonstrates the necessity to investigate positional fairness in addition to consistency when evaluating a judge model’s position bias.
-Appendix K DevBench
+Appendix F Variable Selection and Tests
+F.1 Bidirectional Stepwise Regression with AIC
+Bidirectional stepwise regression is a combination of forward selection and backward elimination techniques.
+It iteratively refines the model by adding or removing predictors based on a statistical criterion—commonly the Akaike Information Criterion (AIC).
+The objective is to select a model that balances goodness of fit and complexity, aiming for the lowest AIC value.
+The AIC is given by:
+$\text{AIC}=2k-2\log(L),$ (4)
+where $L$ is the likelihood of the model and $k$ is the number of parameters in the model, including the error variance $\sigma^{2}$.
+For a linear regression model with independent and identically distributed (iid) errors, $N(0,\sigma^{2})$, fitted to $n$ observations, the log-likelihood can be written as:
+$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$ (5)
+where $\hat{e}_{i}$ is the residual for the $i$th observation, and $\sigma^{2}$ is the variance of the errors. The AIC, in this context, becomes:
+$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$ (6)
+This form of the AIC balances the goodness of fit (as reflected by the residual sum of squares) and model complexity (as represented by $k$).
+The operation of Bidirectional stepwise regression starts with either no predictors (forward selection) or all predictors (backward elimination), where the model iteratively adds or removes variables.
+Each step evaluates the impact on the AIC score.
+In forward selection, variables are added one by one, starting from the null model, such that the addition of each variable results in the largest decrease in AIC.
+In backward elimination, all variables are included in the model initially, and variables are removed one at a time, with the variable whose removal causes the smallest increase in AIC being dropped.
+At each iteration, the change in AIC is computed as $\Delta\text{AIC}=\text{AIC}_{\text{new}}-\text{AIC}_{\text{current}}$, where $\text{AIC}_{\text{new}}$ refers to the AIC after adding or removing a variable, and $\text{AIC}_{\text{current}}$ is the AIC of the current model.
+If $\Delta\text{AIC}<0$, the model is improved by the addition or removal of the variable.
+The process terminates when neither adding nor removing variables results in a lower AIC, signifying that the most parsimonious model, based on AIC, has been reached.
+F.2 Test results
+We operated bidirectional stepwise regression on both benchmarks individually and together to identify the factors that are significantly contributing to position bias.
+Specifically, the variables include lengths (input, output, and prompt), answer quality gap, LLM judges, candidate models, and task categories to predict position consistency and preference fairness respectively.
+Table 5, 6 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively.
+Table 7, 8 serves for DevBench, and Table 3, 4 is conducted on the integrated set of both benchmarks.
+The impact of variables on the model is ranked from highest to lowest, from bottom to top.
+Removed variables listed as None indicate the full model at this given step.
+Through benchmark testing, we verified that LLM judges, task categories, and the answer quality gap significantly contribute to position bias in terms of both position consistency and preference fairness.
+These findings align with our empirical results, showing that position bias varies notably by judge and task, with the answer quality gap being a key influencing factor.
+The extent of this impact is reflected by the magnitude of change in AIC when the given variable is removed.
+It is worth noting that while task output length remains a significant predictor for $PF$ and $PC$ in both benchmarks, the change in AIC magnitude after removing this variable is very minimal.
+This is consistent across both benchmarks individually and combined.
+We therefore conclude that, although position bias is influenced by task output length, this dependency is minimal.
+Table 3: Final results of stepwise model selection for both benchmarks: Position Consistency
+Removed Variables DF Sum of Sq RSS AIC
+None 163.75 -18370
+Task 20 2.832 166.59 -18319
+Candidate 38 4.472 168.23 -18303
+Quality gap 1 21.953 185.71 -17703
+Judge 13 55.417 219.17 -16846
+Table 4: Final results of stepwise model selection for both benchmarks: Preference Fairness
+Removed Variables DF Sum of Sq RSS AIC
+None 254.28 -16103
+Task output length 1 0.836 255.12 -16088
+Quality gap 1 11.339 265.62 -15873
+Task 21 16.177 270.46 -15817
+Judge 13 82.069 336.35 -14641
+Table 5: Final results of stepwise model selection for MTBench: Position Consistency
+Removed Variables DF Sum of Sq RSS AIC
+None 61.974 -13312
+Task output length 1 0.0553 62.029 -13311
+Candidate 29 1.6474 63.621 -13282
+Task 7 1.5304 63.504 -13244
+Judge 13 15.3637 77.338 -12594
+Quality gap 1 15.6206 77.594 -12559
+Table 6: Final results of stepwise model selection for MTBench: Preference Fairness
+Removed Variables DF Sum of Sq RSS AIC
+None 129.00 -10909.2
+Quality gap 1 1.931 130.93 -10861.3
+Task 7 9.295 138.29 -10689.4
+Judge 13 58.847 187.85 -9672.5
+Table 7: Final results of stepwise model selection for DevBench: Position Consistency
+Removed Variables DF Sum of Sq RSS AIC
+None 55.382 -6940.2
+Task output length 1 0.257 55.638 -6933.2
+Candidate 9 1.514 56.896 -6905.4
+Quality gap 1 13.128 68.510 -6525.3
+Judge 13 84.760 140.141 -5146.6
+Table 8: Final results of stepwise model selection for DevBench: Preference Fairness
+Removed Variables DF Sum of Sq RSS AIC
+None 60.104 -6753.9
+Task output length 1 0.061 60.165 -6753.9
+Candidate 9 0.731 60.834 -6748.2
+Task 13 1.305 61.408 -6737.8
+Quality gap 1 1.783 61.886 -6698.6
+Judge 13 80.875 140.979 -5108.9
+Appendix G DevBench
 This session includes a similar baseline comparison analysis on DevBench as on MTBench.
@@
 Similarly, although GPT-4 stands as the baseline model with a generally high $PC$ across tasks, certain models achieve comparable or superior performances on certain tasks.
-For instance, for architecture design evaluations, gpt-4-1106-preview, GPT-4o, and Gemini-1.5-pro all surpass GPT-4.
+For instance, for architecture design evaluations, GPT-4-Turbo, GPT-4o, and Gemini-1.5-pro all surpass GPT-4.
 Gemini-1.5-pro is especially outstanding, also exceeding GPT-4 in uml class evaluations.
@@
 Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
-Appendix L Experiment Settings
+Appendix H Experiment Settings
 This session specifies more detailed information about the judges, answer-generating models, tasks, and prompt templates used in this study.
 We choose to evaluate MTBench and DevBench for the following reasons: (1) all necessary information about the benchmark models, tasks, and questions is publicly available, making modifications convenient (2) they include a wide variety of answer-generating models, tasks, and task questions for a comprehensive evaluation (3) their human evaluations validated the reliability of state-of-the-art judging models (GPT-4 and GPT-4-Turbo) on their evaluation instances, hence model untested by prior work, if reaching high agreement with these validated judges, can be perceived reliable as well.
-L.1 Judges, Models, and Tasks
+H.1 Judges, Candidates, and Tasks
 Judge
@@
 They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations.
-The other models that are compared to the reference models, namely “Model" in our context, are listed as follows.
+The other models that are compared to the reference models, namely “Model” in our context, are listed as follows.
   • MTBench (30): alpaca-13b, baize-v2-13b, chatglm-6b, claude-instant-v1, claude-v1, dolly-v2-12b, falcon-40b-instruct, fastchat-t5-3b, gpt-3.5-turbo, gpt-4, gpt4all-13b-snoozy, guanaco-33b, guanaco-65b, h2ogpt-oasst-open-llama-13b, koala-13b, llama-13b, mpt-30b-chat, mpt-30b-instruct, mpt-7b-chat, nous-hermes-13b, oasst-sft-4-pythia-12b, oasst-sft-7-llama-30b, palm-2-chat-bison-001, rwkv-4-raven-14b, stablelm-tuned-alpha-7b, tulu-30b, vicuna-33b-v1.3, vicuna-7b-v1.3, wizardlm-13b, wizardlm-30b
   • DevBench (10): codellama-7b-instruct, codellama-13b-instruct, codellama-34b-instruct, deepseek-coder-1.3b-instruct, deepseek-coder-6.7b-instruct, deepseek-coder-33b-instruct, gpt-3.5-turbo-1106, gpt-4-0125-preview, gpt-4-0613, gpt-4-1106-preview
-The model names are exactly what MTBench [3] and DevBench [4] used in their studies.
+The model names are exactly what MTBench Zheng et al.
+(2024b) and DevBench Li et al.
+(2024) used in their studies.
 That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4).
@@
     • architecture design (5): conformance, design_and_coding, practicability, uniformity_and_integration, and faithfulness
-L.2 Prompt Settings
+H.2 Prompt Settings
 We follow the original prompt settings of MTBench and DevBench in our study of pairwise comparative LLM-as-a-Judge.
@@
   • A system prompt explaining the judging task and the role the LLM should be playing.
-  • Emphasized “should" and“shouldn’t"s.
+  • Emphasized “should” and“shouldn’t”s.
   • A prompt structure with placeholders for specific questions and model answers
   • A specified output format for later judgment extraction
-  • Chain-of-Thought [53] prompts requiring the LLM judge to provide reasons for its judgment
+  • Chain-of-Thought Wei et al. (2022) prompts requiring the LLM judge to provide reasons for its judgment
 The detailed prompt templates are specified below.
+Appendix I Reproducibility
+Our experiments mainly depend on API. Code will be released upon acceptance.

### v7

@@
-Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs
+Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge
 Abstract
-LLM-as-a-Judge presents a promising alternative to human evaluators across various tasks, but inherent biases, especially position bias — a tendency to favor solutions based on their position in the prompt — have compromised its effectiveness.
-Our study introduces a systematic framework to examine position bias in pairwise comparisons, focusing on repetition stability, position consistency, and preference fairness.
-This research significantly contributes to the field by introducing new concepts for understanding position bias and providing a multi-dimensional framework for evaluations.
-We conducted experiments with 12 LLM judges across MTBench and DevBench, covering 22 tasks and approximately 40 solution-generating models — candidates, resulting in over 100,000 evaluation instances.
-Our findings confirm that position bias in capable LLM judges is not due to random chances, along with notable variations observed across judges and tasks.
-Moreover, position bias is weakly influenced by the length of prompt components but significantly impacted by the quality gap between solutions.
-These insights can help optimize judge model selections, improve benchmark design, and inform future research on debiasing strategies, ultimately enhancing the reliability of LLM judges.
+LLM-as-a-Judge has emerged as a promising alternative to human evaluators across various tasks, yet inherent biases—particularly position bias, the tendency to favor solutions based on their position within the prompt—compromise its reliability.
+This study investigates position bias in LLM judges across pairwise and list-wise comparison settings, introducing three metrics: repetition stability, position consistency, and preference fairness.
+Our experiments, involving 12 LLM judges across MTBench and DevBench with 22 tasks and approximately 40 solution-generating models, result in over 100,000 evaluation instances.
+We identify Judge-Level, Candidate-Level, and Task-Level factors contributing to bias.
+The findings confirm that position bias is not due to random chance and varies significantly across judges and tasks.
+While position bias is weakly influenced by the length of prompt components, it is strongly affected by the quality gap between solutions.
+Our agreement and disagreement analysis among judges further provides insights into the distribution of judging difficulty across the dataset, and highlights the potential for dataset modifications.
+Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge
+Lin Shi, Chiyu Ma, Wenhua Liang, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
 1 Introduction
+Figure 1: Overview of our experiment settings: (a) Position bias is observed when LLM judges consistently favor a specific position rather than evaluating the content, with repeated trials ensuring the deviations are not due to random variations.
+(b) Preference fairness is defined and measured through the distribution of choice pairs to assess the fairness of judgments.
+(c) The settings are extended from pairwise comparisons to list-wise comparisons, involving evaluations of more than two candidate models.
 In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications.
-Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields.
-While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations (Zeng et al., 2023; Karpinska et al., 2021).
+Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields.
+While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations Zeng et al.
+(2023); Karpinska et al.
+(2021).
 To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks.
-Despite a high level of agreement with human judgments(Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
+Despite a high level of agreement with human judgments Zheng et al.
+(2024b); Li et al.
+(2024a); Zhu et al.
+(2023), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
 Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig.
-1.
-This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations (Qin et al., 2024), for any score-based and relation-based settings (Li et al., 2023d).
-Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias.
-However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks (Khan et al., 2024; Chua et al., 2024).
-Moreover, the limited understanding of the bias itself has exacerbated the problem.
-Prior studies on position bias implicitly assume that this bias is not due to random variations.
-In other words, these studies assume that LLM judges produce consistent results across repeated trials.
-However, without validating this assumption, evaluations may become less reliable.
+1 (a).
+This bias has been prevalent in all types of LLM judges Qin et al.
+(2024); Li et al.
+(2023d), raising concerns about their reliability.
+Previous studies Zheng et al.
+(2024a, b); Zeng et al.
+(2023) have identified position bias alongside other biases and assessed its impact.
+Although mitigation strategies have been proposed, they often fall short due to incomplete bias elimination Guo et al.
+(2024), increased complexity Li et al.
+(2024b); Khan et al.
+(2024); Chua et al.
+(2024), the introduction of new biases Ohi et al.
+(2024), or inconsistent effectiveness Gallegos et al.
+(2024).
+These persistent challenges highlight the limitations of current approaches and point to a possible incomplete understanding of position bias.
+Although numerous studies have evaluated position bias in LLM-as-a-Judge settings, the underlying factors contributing to this phenomenon remain insufficiently explored.
+In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge.
+While evaluating Position Consistency, one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability.
+Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences.
+Additionally, the measurement of Repetition Stability ensures that the observed position bias in the given model and tasks is not due to random variations, thus strengthening the reliability of the findings.
+To investigate the underlying factors contributing to position bias, we categorized these factors into three levels: Judge-Level, Candidate-Level, and Task-Level.
+Our experiments are primarily conducted on pairwise comparisons, as LLM judges demonstrate superior performance in this setting.
+We further extend our study to more complicated list-wise comparison settings, involving evaluations of more than two candidate models by LLM judges.
+Our findings reveal several key insights: 1.
+The position bias of capable LLM judges is not a result of random variations.
+2.
+There is a high volatility in the direction of preference, even within the same LLM judge when applied to different tasks.
+3.
+Differences in answer quality among candidate models significantly influence position consistency.
+4.
+Position bias is very weakly correlated with the length of prompts generated by the candidate models.
+Building on these findings, we conduct an agreement analysis among the LLM judges.
+The results reveal that, although measures of position consistency may appear similar in general, judgments on specific instances vary significantly among LLM judges, even when they demonstrate comparable capabilities.
+Instances where numerous LLMs agree are generally easier to judge, whereas instances with disagreements are more challenging to evaluate and more prone to position bias.
+This analysis provides insights into the distribution of judging difficulty across the dataset and highlights the potential for dataset modifications by incorporating more instances that are either easier or more difficult to judge.
+Future work could explore how to measure the likelihood of position bias arise from the datasets by identifying and quantifying such hard-to-judge instances before implementing LLM judges.
+2 Evaluation Settings & Definitions
+We begin by outlining the settings for pairwise and list-wise comparisons employed in our experiments for LLM-as-a-Judge.
+Following this, we define the three metrics used in our evaluation: Position Consistency (PC), Preference Fairness (PF), and Repetition Stability (RS).
+Finally, we provide a detailed description of the factors we found that are related to position bias at the Judge-Level, Candidate-Level, and Task-Level.
+2.1 Pairwise & List-wise Comparison
+Pairwise Comparison: In the context of pairwise comparison, LLM judges are tasked with selecting the better solution provided by two candidate models in response to a given task question.
 As shown in Fig.
-1, the occurrence of position bias can be observed by chance even for the same set of tasks and candidate models.
-Thus, it is unclear whether the observed position bias actually arises from the position of the prompt components or is merely due to random chance.
-Therefore, when exploring the position bias of LLM judges, it is necessary for the judges to be stable across repetitions to yield nontrivial results.
-Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge.
-When the positions of candidate model solutions are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same candidate model solution.
-We observe repetition instability if the same LLM judge gives different judgments across repetitions.
-Only when an LLM judge has high repetition stability can we validate that its position bias is not due to random variations.
-Another key aspect of position bias, preference fairness, has also been under-explored by previous works.
-Most studies have primarily focused on position consistency to measure how frequently position bias occurs across a set of evaluation instances.
-However, the direction of biased preference (i.e., primacy or recency) is also insightful.
-Fig.
-2 illustrates cases where a pair of judgments with swapped prompts become position consistent, inconsistent, primacy-preferred, or recency-preferred.
-We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations.
-Specifically, judgments are considered to have a fair preference if they are position consistent or if the preference directions are evenly distributed.
-In contrast, an LLM judge is considered to have an unfair preference if it consistently favors candidate model solution that appear either first or last in the prompt.
-Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
-Figure 2: Demonstration of position consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better candidate solution.
-A judgment is consistent if it selects the same winner across different prompt positions.
-Primacy/recency indicates a preference for the first/last candidate solution, respectively.
-Position inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
-To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment.
-We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring (Zheng et al., 2024b; Liusie et al., 2024), making it a practically preferable choice for investigations.
-Our contribution can be summarized as following: (1) introducing a systematic framework for evaluating LLM-as-a-Judge’s position bias in terms of Repetition Stability, Position Consistency, and Preference Fairness; (2) identifying Judge-level, Candidate-level, and Task-level factors that affect position bias; and (3) providing insights into future benchmark design and paving the way for effective mitigation strategies for position bias.
-To better understand position bias and the effectiveness of LLMs as evaluators, we have developed a detailed framework specifically for analyzing position bias in scenarios where two options are compared directly, known as pairwise comparative assessment.
-We choose this approach because LLMs tend to perform better in these head-to-head comparisons than in other formats like scoring each option separately (Zheng et al., 2024b; Liusie et al., 2024).
-This makes pairwise comparisons a practical focus for our research.
-Our contributions are threefold: (1) We introduce a systematic framework to analyze how LLMs judge by focusing on Repetition Stability, Position Consistency, and Preference Fairness.
-(2) We explore factors at different levels—Judge, Candidate, and Task—that influence position bias.
-(3) We offer insights that will help in designing better benchmarks and developing strategies to reduce position bias in LLM evaluatio
-2 Methods & Defintions
-LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments.
-under OuFirst, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields (Qin et al., 2024).
-Second, as noted by Liu et al.
-(2024), LLM and human evaluations are more aligned in the context of pairwise comparison than any of the score-based assessments.
-Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of position consistency(Zheng et al., 2024b; Liusie et al., 2024).
-Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, through some advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024).
-Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
-In the context of pairwise comparative assessments, LLM judges are prompted to choose the better candidate solution to the task question.
-To accommodate the possibility of a tie, various option modes are introduced.
-Two-Option mode restricts judges to choosing the better solution between two options, labeled {A} for the first and {B} for the second as shown in Fig.
-1.
-Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither solution is preferable.
-Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie.
-In our experiments, we employed Three-Option mode for MTBench (Zheng et al., 2024b) and Two-Option mode for DevBench (Li et al., 2024), aligning with their respective original frameworks.
-These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM judges.
-To set up our experiment, as shown in Fig.
-1, the system prompt, option mode, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the winning solution.
+1 (a), the system prompt, option choices, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the better solution.
 The experiment is conducted in a double-blind setting.
@@
 Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair.
-As shown in Fig.
-2, if the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent with a fair preference.
-Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
-To systematically investigate position bias in pairwise comparative judging scenarios, we evaluate LLM judges using three key metrics: Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$).
-While $PC$ has been widely adopted in prior research, we introduce $RS$ to validate the reliability of LLM judges and $PF$ to provide a more comprehensive assessment.
-Further analysis examines the impact of judge-level, candidate-level, and task-level factors on position bias.
-Therefore, for each judge, candidate, and task, we calculate $RS$, $PC$, and $PF$ to thoroughly explore the variance in position bias and its impacting factors.
+If the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent.
+Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy (e.g.
+always choose $\{A\}$) or recency (e.g.
+always choose $\{B\}$).
+Example of measuring preference fairness with specific choice pairs is shown in Fig.
+1 (b).
+To accommodate the possibility of ties, various option modes are employed: Two-Option mode restricts judges to choosing between two options, labeled A for the first candidate and B for the second.
+Three-Option mode introduces an additional choice, C, allowing judges to indicate a tie if neither solution is preferable, as illustrated in Fig.
+1 (a).
+These option modes were explicitly specified in the system prompts to ensure clear guidance for the decision-making process of the LLM judges.
+List-wise Comparison: Unlike pairwise settings, where LLM judges select the superior solution from two candidates, list-wise comparative approaches involve evaluating three or more candidates simultaneously, as shown in Fig.
+1 (c).
+For efficiency, we prompt judges to select the best candidate rather than ranking the entire list.
+The “swapped setting" used in pairwise evaluations is generalized to order permutations for list-wise judgments, ensuring that each candidate appears in every possible position exactly once.
+For a list of $p$ candidates, this results in $p$ permutations.
+In the $i^{th}$ permutation, the $i^{th}$ candidate is set to appear in the first position.
+Additionally, an option is provided to account for ties, allowing judges to indicate if there is no certainly superior solution.
 2.2 Evaluation Metrics
-We validate the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows.
+In our study, we first verify if the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness.
+The metrics are introduced as follows:
 Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times.
-It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations.
+It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or by random variations.
 We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset.
 This metric is formalized as
-$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,...\}$ (1)
-where $S$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries.
+$RC$ $=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{k\in S}\left\{|C_{k}^{j}|\right\},$ (1)
+where $S=\{A,B,C,\dots\}$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries.
 The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
-Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is swapped.
-It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the judge model prefers the same winning solution.
+Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is permuted.
+It is calculated as the ratio of consistent evaluation series to the total number of valid evaluations, where a series is deemed consistent if the judge model prefers the same winning solution across permutations.
 Formally, it is calculated as
-$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}},V=\{(A,B),(B,A),(C,C),...\}$ (2)
-where $(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})$ denotes the judgment pair for the $j^{th}$ query, $n$ represents the number of prompt pairs, and $V$ is the set of choice pairs that correspond to position consistency.
-An example of such pairs of choices can be found in Fig.
-2.
-This formula provides a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
-Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain solution positions.
-In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency.
-These terms replace the more verbose “preference for the first/second position” used in previous studies, ensuring clarity and generalization for future research.
+$PC$ $=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{1}^{j},\dots,C_{p}^{j},)\in V\}},$ (2)
+where $V$ is the set of choices that correspond to position consistency, and $(C_{1}^{j},\dots,C_{p}^{j})$ denotes the judgment series for the $j^{th}$ query when there are $p$ candidate solutions in the list, and $n$ represents the number of prompt series.
+An example of such series of choices under pairwise comparison setting can be found in Fig.
+1 (b).
+This formula aims to provide a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
+Preference Fairness ($PF$) measures the extent to which judge models favor certain solution positions.
+In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency.
+These terms replace the more verbose “preference for the first/second candidate model" used in previous studies, ensuring clarity and generalization for future research.
 The examples of such preferences are demonstrated in Fig.
-2.
-Position consistent judgments are fair in preference.
-The measurements become more complex for position inconsistent cases.
+1 (b).
 Previous studies proposed two common ways to measure the preference fairness.
-A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$).
-The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023b).
+One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$).
+The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023).
 However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
@@
 We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$.
-However, this approach is problematic because it overlooks the fact that “position consistent judgments are also position fair,” which can lead to overly penalizing highly consistent LLM-judges.
-For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred.
-However, this outcome should be distinguished from cases where all one hundred pairs are inconsistent and biased toward one position, which would also result in 100% preference unfairness.
+However, this approach overlooks the fact that “position consistent judgments are also preference fair”, which leads to overly penalizing highly consistent LLM-judges.
 To overcome these limitations, we introduce a more granular and scalable measurement that combines the strengths of both methods, to assess preference fairness.
 The $PF$ score is formally calculated by
-$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$ (3)
+$PF$ $=\frac{PF_{\text{raw}}-S^{-}_{\min}}{S^{+}_{\max}-S^{-}_{\min}}\times 2-1,$ (3)
+$PF_{\text{raw}}$ $=(rcn\times irr)-(pcn\times ipr).$
 where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively.
-This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero.
-Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the position consistent judgments.
-Another advantage of this approach is that it quantifies preference fairness with a single score, rather than the two separate scores used in previous studies.
+This min-max scale ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero.
 The $PF$ score is interpreted as follows:
 $PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
-To summarize, our proposed preference fairness score is superior to prior measurements due to its ability to provide a single, comprehensive score that accounts for all evaluation instances, while maintaining its sensitivity across datasets, even when the number of questions and instances varies for each task.
+To extend this metric to list-wise comparisons, we employed a ‘one vs.
+all’ approach, defining primacy preference as favoring the first candidate solution while classifying all others as recency-preferred.
+This straightforward extension of the $PF$ computation maintains consistency with pairwise setups.
+By providing a single and comprehensive metric that applies to all evaluation instances and list-wise settings, our proposed $PF$ score ensures sensitivity across datasets, regardless of variations in the number of questions or instances, offering a significant improvement over previous methods.
 2.3 Factors Affecting Position Bias
-Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose solutions to the task-specific questions are evaluated by the judges.
-Factors that have a significant impact on position bias are highlighted with an asterisk (*) by bidirectional stepwise regression tests and marked red by empirical findings.
-‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s solutions to the question, making it both a Candidate-level and Task-level factor.
-‘Prompt’ encompasses the complete query presented to the judges, including Task Input, Task Output, and system prompts.
 Factor Judge-level Candidate-level Task-level
@@
 Prompt Length $\times$ $\checkmark$ $\checkmark$
+Table 1: Factors influencing position bias.
+Significant factors, identified via bidirectional stepwise regression, are marked with * and highlighted in red based on empirical findings.
+Task Input refers to the question itself, while Task Output denotes the candidate model’s answers, serving as both Candidate-level and Task-level factors.
+Prompt includes the full query presented to judges: Task Input, Task Output, and system prompts.
 To investigate the factors influencing position bias in LLM judges, we categorized these factors into three groups: Judge-level, Candidate-level, and Task-level factors.
 Each group includes specific factors, that we hypothesize, may impact position bias, which we explore through a series of experiments.
-Table 1 lists the five factors analyzed in this study.
-By design, our framework allows for the integration and assessment of additional influencing factors.
-Among the influencing factors we identified, we chose “familial property” for Judge-level factors instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments.
-The model family grouping reflects similar model sizes and training specifics of the model members.
+Table 1 lists the five factors we analyzed in this study.
+Among the influencing factors, we selected “familial property” for Judge-level factors, as it reflects similar model sizes and training specifics, which are often proprietary and not publicly accessible.
 The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version.
-However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable” models sharing similar judging capabilities.
 More details and discussions about the familial property can be found in Appendix Sec.
-C, where we examined the LLM mutual agreements to study the similarities and differences across families.
-Answer quality gap, based on our findings, is highly impactful yet under-explored in prior research.
-We define the quality of a candidate’s answer/solution by how effectively it addresses the question.
-Consequently, the answer quality gap refers to the disparity in quality between the solutions from two candidate models to the same question and is considered the Candidate-level (i.e., solution provider) factor.
-Ideally, when a reliable LLM judge is presented with a question and corresponding solution pairs, it would prefer the higher-quality solution, where the corresponding candidate is denoted as the winner selected by the judge.
-Following this assumption, we measure the answer quality gap by the win rates between two candidates on a set of tasks and questions.
-We regard a tie judgment as ‘half-win’ for both candidates when calculating its win rate over another, depending on option mode.
-If position bias occurs, the winner may be inconsistent when the order of candidate solutions is swapped in the query.
-Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen in both the original and swapped queries (termed “consistent wins”), cases where both responses are deemed equal (termed “consistent ties”), and cases where different winners are selected after the solutions are swapped.
-We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment pairs ($C_{I}$), respectively.
-Following Zheng et al.
-(2024b), we count inconsistent judgment pairs also as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
-Inspired by Zheng et al.
-(2024b), we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs.
-Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, where 0.5 is the overall win rate when we only have ties.
-In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
-More discussions can be found in Appendix.
-Sec.
-D.
-3 Experiments
+C.
+Answer quality gap, based on our findings, is highly influential yet under-explored in prior research.
+We define the quality of a candidate’s solution by how effectively it addresses the question.
+Consequently, the answer quality gap refers to the disparity in quality between the solutions from one candidate model and the others to the same question and hence considered the Candidate-level factor.
+Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
+Following this assumption, we measure the answer quality gap by the win rates of candidates over an expected baseline on a set of tasks and questions.
+However, if position bias occurs, the winner may be inconsistent when the order of candidate solutions is permuted in the query.
+Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen across all permutations (termed “consistent wins”), cases where there is no certain winner (termed “consistent ties”), and cases where different winners are selected after the solutions are permuted in the queries (termed “inconsistent judgment series").
+We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment series ($C_{I}$), respectively.
+Inspired by Zheng et al., we count inconsistent judgment pairs as ties for all candidate models, which is later calculated as a down-scaled win rate depending on the number of candidate models.
+To calculate the win rates of candidate models for all three cases, we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{p}(C_{t}+C_{I})}]$, where we have $p$ candidates in the list and $n$ judgment series.
+Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-1/p|$, where $1/p$ is the expected baseline when all judgments are “ties”.
+In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment series) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
+3 Experiment
 3.1 Experiment Settings
-In this study, we evaluated twelve up-to-date commercial models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework.
+In this study, we evaluated twelve up-to-date close-sourced models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework.
 Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks.
-However, due to their limited performance on our tasks, as shown in Table 2, we focused our investigation solely on more powerful closed-source models.
+However, due to their limited context window size and failure to follow the output format, which led to suboptimal performance on our tasks (as shown in Table 2), we focused our investigation exclusively on more powerful closed-source models.
 We included the results of open-source models for completeness.
-We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks.
-We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
+More discussions can be found in Appendix.Sec.
+D
+We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks.
+For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
 MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task.
 We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges.
-The prompt templates we used are identical to those used in the benchmarks.
+We adopted Two-option mode for MTBench, and Three-option mode for DevBench.
+For list-wise experiments, we randomly sampled 9 models to form three triple-candidate lists and evaluated four representative judges on MTBench.
+The prompt templates we used are identical to those in the benchmarks for pairwise comparisons, with minor modifications to accommodate list-wise evaluations.
 More details about the models, tasks, and prompts can be found in Appendix.
 Sec.
-H.
+G.
 To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials.
@@
 To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets.
-In total, approximately 100,000 evaluation instances were analyzed in this study.
+In total, more than 100,000 evaluation instances were analyzed in this study.
 To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks.
@@
 Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score.
-This process involves both forward selection and backward elimination, with each ”step” testing whether including or excluding a variable improves the model’s AIC value.
+This process involves both forward selection and backward elimination, with each "step" testing whether including or excluding a variable improves the model’s AIC value.
 Further details about the process can be found in Appendix.
@@
 F.
-3.2 Empirical Results and Interpretations
+3.2 Empirical Results
 The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2.
@@
 For $RS$ and $PC$, higher values are preferable.
-A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not merely random variations.
+A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not due to random variations.
 Fig.
-3 and Fig.
-7 demonstrate that position bias varies by judges and tasks significantly.
+2 (a)(b) demonstrate that position bias varies by judges and tasks significantly.
 Fig.
-4 further investigates the impact of answer quality gap and lengths of prompt components on position bias.
+2 (c) explores the correlation between the metrics $PC$ and $PF$.
+Fig.
+(d) to (f) further investigate the impact of the answer quality gap on position bias.
 These analyses were conducted by considering all judges together on MTBench.
-The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E, and Sec.
-G.
+The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E.
 Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables.
@@
 F.
-Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$).
-Top 5 performances are marked in bold.
-“Error” is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format.
-High error rates and low $RS$ are marked red, implying further evaluations (i.e., $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data.
-Therefore, open-source models (e.g., Llama) are not investigated in the paper because of their high error rates under the difficulty and length of our evaluation instances.
-In contrast, up-to-date commercial models generally demonstrated minimal error and high repetition stability, making them more practical and meaningful for investigations.
-Chatbot Arean Leaderboard’s overall ranking of the models is listed.
-Judge MTBench DevBench Arena
-$RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error
-Caude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 7
-Claude-3-Opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00 17
-Claude-3-Sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00 37
-Claude-3-Haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00 46
-Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 12
-Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96 25
-Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00 66
-o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00 2
-GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 6
-GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00 17
-GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00 52
-GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 94
-Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 99
-Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 109
-Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94 57
-Figure 3: Baseline comparisons of judges across tasks on MTBench.
-An asterisk marks the statistical significance by Student’s t-tests.
-Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$.
-The baseline model is chosen to be GPT-4 for its generally superior performances.
-Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
-Figure 4: Impact of answer quality gap and lengths on position bias.
-All figures are based on the integration results of all judge models on MTBench.
-Figures (a) and (b) investigate $PC$ and $PF$ vs.
-overall win rate, hence the answer quality gap, respectively.
-Generally, judgments become more consistent and fair as the quality gap increases.
-Figure (c) demonstrates the relationship between $PC$ and $PF$, where higher consistency is likely to indicate better fairness.
-Figure (d)(e)(f) plots the relationship between lengths of prompt components and $PC$.
-The irregular and inconclusive patterns imply that position bias is not significantly impacted by the length of prompts and their components.
+Judge MTBench Pairwise DevBench Pairwise MTBench List-wise
+$RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error $PC$ $PF$ Error
+Claude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 0.67 $\pm$ 0.19 0.17 $\pm$ 0.19 0.00
+Claude-3-Opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00
+Claude-3-Sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00
+Claude-3-Haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00
+Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 0.55 $\pm$ 0.20 0.33 $\pm$ 0.18 0.00
+Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96
+Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
+GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 0.68 $\pm$ 0.22 0.18 $\pm$ 0.22 0.00
+o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00
+GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
+GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
+GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 0.34 $\pm$ 0.17 -0.05 $\pm$ 0.30 0.12
+Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
+Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
+Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
+Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold.
+Errors arise from judgment failures (e.g., exceeding context window, not following output format).
+High error rates and low $RS$ are marked red, rendering further evaluations invalid due to insufficient data.
+Open-source models (e.g., Llama) are excluded due to high error rates with our lengthy instances (See Appendix.Sec.D for more discussions).
+List-wise evaluation is conducted on four representative judges from each family to validate scalability.
+Figure 2: Judge performances on MTBench.
+Fig.
+(a)(b) are the radar charts for the $PC$ comparison by family, judge, and task Fig.
+(c) to (f) investigate the impact of task output length and answer quality gap on position bias.
+Fig.
+(a)(c)(d)(e) are for pairwise comparative settings, while Fig.
+(b)(f) are obtained under list-wise evaluations.
 4 Main Findings
-Position Bias of Capable Judges are not Mere Random Variations As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal ”Error” rates, generally exhibit $RS$ values above 0.85.
-The most up-to-date and capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks.
-These results confirm that LLM judgments, and the resulting position bias, are not merely random variations.
+Position Bias of Capable Judges are not Mere Random Variations: As shown in Table 2, the capable judges on the benchmark tasks, supported by minimal "Error" rates, generally exhibit $RS$ values above 0.85.
+The most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks.
+These results confirm that LLM judgments, and the resulting position bias, are not due to random variations.
 This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
-Position Bias Varies by Judge & Task As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig.
-3 for $PC$ comparisons.
-However, certain models achieve comparable or superior performances than GPT-4 on certain tasks.
+Position Bias Varies by Judge & Task: As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig.
+2 (a) for $PC$ comparisons.
+Moreover, certain models achieve comparable or superior performances than GPT-4 on certain tasks.
 For example, according to Fig.
-3, for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges.
-Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it may be employed as a cost-effective alternative to coding evaluations.
+2 (a), for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges.
+Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it could be employed as a cost-effective alternative to coding evaluations.
 Additionally, significant variations of $PF$ across tasks are observed.
-The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on the others.
+The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others.
 Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, role play, and writing tasks.
-Even for uniformly recency-preferred judges such as Claude-3’s and Gemini-pro’s, the extent of biased preference, as reflected by $PF$ values, varies by task.
-Moreover, high position consistency does not guarantee fairness.
+Even for judges that are recency-preferred across all tasks (e.g., Claude-3’s and Gemini-pro’s), the extent of biased preference, as reflected by $PF$ values, varies by task.
+Moreover, a high position consistency does not guarantee fairness.
 For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively.
 In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
+In list-wise comparisons, similar variations by judge and task in position bias were observed, as shown in Fig.
+2 (b).
+Furthermore, Table 2 highlights that more capable models, such as GPT-4o and Claude-3.5-Sonnet, maintain high consistency when transitioning from pairwise to list-wise evaluations, while less capable models, such as GPT-3.5-Turbo, exhibit greater sensitivity to the increased number of candidates in list-wise tasks.
 Therefore, the position bias of LLM judges is judge-dependent and task-dependent.
-This observation is confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$.
-In practice, to achieve the most reliable evaluation results, one needs to select task-specific LLMs whose consistency and fairness are in balance.
-Moreover, as shown in Table 2, a high ranking in the overall reasoning performance of a model does not guarantee its judging reliability.
+This observation is further confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$.
+In practice, to achieve the most reliable evaluations, one needs to select task-specific LLMs whose consistency and fairness are in balance.
 Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec.
-G.
-Position Bias Correlates to Answer Quality Gap Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
-In this study, as defined in Section 2.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$).
-Since an absolute tie has 0.5 $owr$, $\delta_{p}$ increases as $owr$ extends from 0.5 to 0 or 1.
+E.2.
+Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
+In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate ($owr$).
+Therefore, $\delta_{q}$ increases as $owr$ extends from baseline to 0 or 1.
 Fig.
-4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$.
-This aligns with our intuition that the answer pairs with larger quality disparities are easier to achieve consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$.
+2 (e) and (f) exhibit significant parabolic shapes, indicating that $PC$ is positively proportional to $\delta_{q}$.
+This aligns with our intuition that the answers with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$.
 The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.
@@
 Similarly, as shown in Fig.
-1(b), judgments generally become more preference fair as $\delta_{q}$ increases.
+2 (d), judgments generally become more preference fair as $\delta_{q}$ increases.
 However, the extent is not as significant as for $PC$.
 Also, the relationship varies by judge, as some LLMs maintain preference fairness regardless of $\delta_{q}$.
-For example, as shown in Fig.
+For example, as shown in Appendix.Fig.
 6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
 These observations align with the right-arrow shape as demonstrated in Fig.
-4(c), where there is a general trend that judgments become preference fairer as position consistency increases.
+2 (c), where there is a general trend that judgments become preference fairer as position consistency increases.
 It also justifies the reasonableness of our quantification of preference fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
@@
 Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
-Our investigation also helps to explain the observed verbosity/length bias (Liu et al., 2024; Kim et al., 2023; Zheng et al., 2024b) and self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b) of LLM judges.
-From our analysis, LLMs may prefer longer answers (verbosity/length bias) since more capable models usually generate longer solutions with higher qualities.
-Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the solutions generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4).
-In both cases, the observed biased preference may be fundamentally due to a prominent answer quality gap instead of length or self-identity.
 Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length).
-Since the tokenization methods vary by models, we apply the length of the prompt component strings for uniformity and scalability.
-The analysis from Fig.
-4(d)(e)(f) shows that there is very weak relationship between the lengths of prompt components and position bias.
-By stepwise regression, the average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table.
+By stepwise regression, we discovered that average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table.
 4.
-Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
-Practical Implications for Benchmark Design We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appendix.Sec.
-C to gain further insights.
-In particular, the disagreement analysis helps identify instances where it is easy or difficult for judges to reach a consensus, reflecting the complexity of the judgment task.
-Based on our findings regarding the answer quality gap and LLM agreement/disagreement, this study offers practical implications for designing evaluator benchmarks that control the difficulty of judging tasks.
-Our findings show that the hardest-to-judge instances are those where (1) LLMs frequently disagree with each other, (2) comparable win rates and minimal quality gaps exist between the two candidate models, and (3) significant position bias is exhibited by the majority of judges.
-One qualitative analysis example of a difficult-to-judge instance comes from MTBench question 120, which involved a math problem where both candidate models, Vicuna-13b-v1.3 and WizardLM-30b, provided incorrect solutions.
-The problem was: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find $x$ such that $f(x)=0$.
-Vicuna-13b-v1.3 incorrectly applied the quadratic formula to solve the cubic equation, while WizardLM-30b used an enumeration method but failed to identify the correct answer.
-This resulted in a 4{A}-3{B}-5{C} choice pattern among the twelve judges, illustrating a typical case of significant disagreement.
-We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer.
-Others preferred either correctness (choosing {C} as a tie) or methodological soundness (choosing {B}, favoring WizardLM-30b).
-In the swapped order setting, 9 out of 12 judges displayed position bias.
-This highlights the challenge of judging solutions with minimal quality gaps, leading to greater position bias and subjectivity.
-The key takeaway is that reducing hard-to-evaluate, highly subjective instances can improve benchmark design.
-Likewise, minimizing trivial cases where judges easily agree would enhance the benchmark’s robustness.
-Clearer and more specific evaluation criteria could also help resolve ambiguity in difficult instances.
+In other words, there is a very weak relationship between the lengths of prompt components and position bias.
+LLM Agreement Analysis: We complement our investigation of position bias with an agreement and disagreement analysis among LLM judges.
+Rather than focusing exclusively on overall consistency or fairness, we examine how judges converge and diverge in their assessments of individual instances.
+Agreement analysis quantifies the percentage of instances where two judges mutually agree on the outcome.
+Disagreement analysis counts the number of choices deviating from the mode for each instance among all judges.
+This further complies a “distribution of disagreement” across the dataset.
+Our findings reveal that, despite exhibiting similar overall $PC$ and $PF$ scores, judges vary significantly in their judgments on individual instance.
+Disagreement analysis, in particular, highlights instances where consensus is either easily or difficultly achieved, reflecting the inherent complexity of the judgment task.
+For example, as shown in Fig.
+3, more than half of the dataset can be considered relatively easy to judge, as over 75% of all 12 judges agree with each other on these instances.
+Conversely, fewer than 4% of instances represent the likely hard-to-judge cases where a majority of judges fail to reach consensus.
+Figure 3: Distribution of disagreement on MTBench. The y-axis indicates the proportion of the dataset where the level of disagreement among judges does not exceed a specific threshold.
+Based on our observations of answer quality gaps and LLM agreement/disagreement patterns, this study offers practical insights for designing evaluator benchmarks that account for the varying difficulty levels of judgment tasks.
+Specifically, the most challenging instances to evaluate are characterized by: (1) frequent disagreements among LLM judges, (2) closely matched win rates and minimal quality gaps among candidate models, and (3) significant position bias exhibited by the majority of judges.
+Further discussions and analyses can be found in Appendix Sec.
+C.
 5 Conclusion
-In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of LLM judges in pairwise comparative scenarios in terms of repetition stability, position consistency, and preference fairness.
-Through comprehensive evaluations of 12 judges on two benchmarks across 22 tasks with over 100,000 evaluation instances, we observe significant variations of position bias across judges and tasks.
-We also discover that position bias is weakly length-dependent but significantly impacted by the answer quality gap.
-Our findings enhance the understanding of position bias, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
+In conclusion, this paper provides an in-depth analysis of position bias in LLM judges, a critical challenge in automated evaluation.
+Using metrics such as repetition stability, position consistency, and preference fairness, we identify significant variations in position bias across judges and tasks, consistent across pairwise and list-wise comparison settings.
+Our findings show that position bias is weakly influenced by prompt length but strongly impacted by the quality gap between solutions.
+Furthermore, agreement and disagreement analysis highlights variability in judgment reliability, providing valuable insights for refining benchmarks.
+This study enhances understanding of position bias and contributes to the development of fairer and more reliable LLM evaluation frameworks.
+6 Limitations
+Despite proposing scalable metrics and investigating key factors influencing position bias, our study has several limitations.
+First, due to computational constraints, we evaluated only twelve commercial LLM judges for pairwise settings and four for list-wise settings, limited to 3-candidate lists, using two benchmarks.
+Open-source models, such as Llama 7B-13B, struggled to provide reliable judgments, restricting their inclusion.
+Additionally, while we tested original benchmark prompt templates, exploring variations in prompting techniques could provide further insights.
+Future work could expand this study to include more models, tasks, prompting strategies, and larger list-wise candidate pools to validate and generalize our findings.
+Second, data accessibility limitations prevented a direct analysis of Judge-level factors like architecture and parameter size.
+Instead, we approximated these factors by grouping models by family properties.
+Open-source models with accessible architectural details could allow deeper analysis of these influences.
+Additionally, our analyses rely on completed judgments, such as measuring the answer quality gap and task difficulty after evaluations.
+Future work could explore methods to estimate or control these factors pre-judgment, reducing computational costs and enabling proactive mitigation strategies.
+Lastly, our focus was on understanding position bias rather than mitigating it.
+While our findings provide a foundation for effective mitigation, further research is needed to address issues like maintaining consistency and fairness when answer quality gaps are minimal, where position bias is most pronounced.
+Multivariate analyses exploring interactions between factors like prompt length, task complexity, and answer quality gaps could also yield deeper insights and enhance mitigation approaches.
 Appendix Table of Contents
@@
 [appendices]0
-Appendix
-Appendix A Contribution & Limitation
-Contributions
-Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge.
-Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through position consistency and preference fairness.
-We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components.
-Particularly, our study is the first to establish a direct connection between quantified answer quality disparities and position bias, highlighting the urgent need for a comprehensive understanding of how position bias affects LLM judges.
-Additionally, the mutual agreement and disagreement analysis of LLM judges are insightful for identifying the difficulty of evaluation instances and beneficial for benchmark assessment and designs.
-All findings from our study enhance the understanding of the position bias of LLM-as-a-Judge, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
-Social Benefits Our comprehensive understanding of position bias in LLM-as-a-Judge provides multiple benefits to the community.
-First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation.
-Our focus and investigation on preference fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability.
-Moreover, our emphasized trade-offs between position consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work.
-Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results.
-Besides, the LLM agreement/disagreement analysis in Appendix.Sec.C 5 provides an applicable tool to identify the difficulty of judging certain instances.
-Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias.
-Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare (Chen et al., 2024c), instruction following and prompting (Zeng et al., 2023; Chua et al., 2024), multimodal assessment (Chen et al., 2024a), and recommender systems (Hou et al., 2024).
-Strength
-Our framework of systematically understanding position bias has the following strengths: (1) scalability and ease of use - our proposed framework is scalable to various judges, models, tasks, and judging types.
-It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works.
-(2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, position consistency, and preference fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias.
-(3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
-Limitations
-Due to computational limits, our study can be potentially extended to provide more comprehensive insights.
-(1) We only study the pairwise comparative assessment of LLM-as-a-Judge for its best consistency, ease of use, and scalability, but more types of judging can be explored, such as pairwise scoring and listwise ranking.
-(2) We only study 12 up-to-date commercial LLM judges for systematic investigations.
-Llama 7b-13b models that we choose as representatives of open-source models of similar sizes, face extreme difficulty in judging the instances explored in this study.
-Due to limited resources and considering practicality, we did not experiment with open-source models with larger sizes.
-However, further exploration may include analysis of these state-of-the-art open-source and fine-tuned judge models using our proposed framework.
-(3) We only study position bias on two benchmarks, MTBench and DevBench, which is still limited despite 22 tasks and 40 answer-generating models.
-Also, the number of questions for each task is limited.
-Further studies can be extended to more benchmarks to have a broader understanding of position bias across benchmarks, tasks, and questions.
-(4) Due to data accessibility, we are unable to investigate Judge-level factors in terms of model parameter sizes and architectures, which we expect will have a significant impact on position bias.
-Alternatively, we consider the family property to account for Judge-level factors by grouping models with similar properties (e.g., approximate parameter size) together.
-We also investigate variances of position bias across individual judges.
-However, future exploration on these model properties, especially those available on open-source models, will provide more insights.
-Future Work
-Besides addressing the limitations, the following future work, built on findings from our study, could provide more insights: (1)prompt setting - in this study, we apply the exact default prompt settings of MTBench and DevBench.
-However, the positional order and prompt style of not just the model-generated responses but also system prompt components (e.g., agent role assignment, mode of judging, direct mention of not making biases) may also have an impact on the extent to which LLM judges make position bias.
-(2) open-source insights: Further investigations of open-source and fine-tuned models on easier-to-evaluate benchmark may provide more insights into the training stages or model structures that are affecting or causing the position bias.
-Then our framework offers better control on the other impactful factors such as the answer quality gap to make these further investigations more accurate.
+Appendix A Reproducibility
+Our experiments were conducted primarily using API access, with a total cost of approximately 3,000 USD.
+The code repository for reproducing our results is available at: https://anonymous.4open.science/r/Position-Bias-Analyzer-Demo-F7E3
 Appendix B Related Work
 B.1 LLM-as-a-Judge
-In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications.
-For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature (Zeng et al., 2023).
-However, it lacks scalability and reproducibility (Karpinska et al., 2021).
-As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks (Zheng et al., 2024b; Li et al., 2024; Chiang & Lee, 2023b; Liusie et al., 2024; Chen et al., 2023; Fu et al., 2023; Wang et al., 2023a; Kocmi & Federmann, 2023b; Ji et al., 2023; Liu et al., 2024; Zeng et al., 2023; Liu et al., 2023; Shen et al., 2023; Kocmi & Federmann, 2023a; Chiang & Lee, 2023a; Karpinska et al., 2021; Dubois et al., 2024; Kim et al., 2023; Liang et al., 2023; Chang et al., 2023; Zhong et al., 2023; Chia et al., 2023; Deb & Das, 2023; Ye et al., 2024; Min et al., 2023; Li et al., 2023d; c; b; Lu et al., 2024), including open-ended story generation (Chiang & Lee, 2023a), adversarial attacks (Chiang & Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi & Federmann, 2023a), and instruction following (Zeng et al., 2023).
-These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles (Zheng et al., 2024b; a; Pezeshkpour & Hruschka, 2023; Chiang & Lee, 2023b; Shen et al., 2023; Liusie et al., 2024; Qin et al., 2024; Chen et al., 2023; Huang et al., 2024; Zhu et al., 2023b; Raina et al., 2024; Chen et al., 2024b; Wang et al., 2023b; Liu et al., 2024; Zeng et al., 2023; Khan et al., 2024; Kim et al., 2023; Li et al., 2023b; d; Shi et al., 2024; Li et al., 2023c; Chen et al., 2024c; Chua et al., 2024; Chen et al., 2024a).
-As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation.
-Therefore, cost-effective LLM judges are needed to assess these responses efficiently.
-LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
-In many of these tasks, LLM judges have shown a high level of agreement with human evaluators (Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), yet in some tasks, they are less effective, largely due to inherent biases (Shen et al., 2023; Chiang & Lee, 2023b; Zheng et al., 2024a).
-Even in cases where agreement is high, judgments may still suffer from biases.
-When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method.
-From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment (Qin et al., 2024).
-By comparative method, it can be either score-based or relation-based (Li et al., 2023d).
-For example, pointwise scoring (Kim et al., 2023) lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics.
-Pairwise/Listwise Scoring (Karpinska et al., 2021; Zhu et al., 2023b; Chen et al., 2024a; Zheng et al., 2024b) prompts the LLM judges to score a pair/list of model-generated answers.
-Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst).
-Pairwise comparative assessment (Karpinska et al., 2021; Chen et al., 2024a; Zheng et al., 2024b), on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
-Except for pointwise evaluation, all forms of LLM-as-a-Judge suffer—or are susceptible to suffering—from position bias due to the intrinsic nature of position and comparison within the prompt structure.
+Large Language Models (LLMs) have become a transformative tool in automating evaluative tasks, offering scalability and reproducibility advantages over human assessments.
+The methodology of using LLMs as evaluators ("LLM-as-a-Judge") has been widely used for tasks such as open-ended story generation (Chiang and Lee, 2023a), adversarial attacks (Chiang and Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi and Federmann, 2023), and instruction following (Zeng et al., 2023), where models are tasked with scoring or ranking outputs.
+Despite their potential, inherent biases—particularly position bias—pose significant challenges to their reliability and fairness, even in the most effective pairwise comparative settings.
 B.2 Position Bias
-Previous studies have discovered and investigated multiple types of biases, such as position bias (Zheng et al., 2024b; Li et al., 2024; Chen et al., 2024a; 2023; Chia et al., 2023; Chiang & Lee, 2023b; Khan et al., 2024; Zhu et al., 2023b; Zhong et al., 2023; Zheng et al., 2024a; Deb & Das, 2023; Zeng et al., 2023; Wang et al., 2023c; Qin et al., 2024; Wang et al., 2023b; Shi et al., 2024; Liusie et al., 2024; Liu et al., 2024; Kim et al., 2023; Karpinska et al., 2021; Hou et al., 2024; Kocmi & Federmann, 2023b; Pezeshkpour & Hruschka, 2023; Shen et al., 2023; Chiang & Lee, 2023a; Kocmi & Federmann, 2023a; Chen et al., 2024b; Li et al., 2023c; d; Raina et al., 2024), verbosity/length bias (Zheng et al., 2024b; Liu et al., 2024; Kim et al., 2023), self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b), selection bias (Zheng et al., 2024a; Pezeshkpour & Hruschka, 2023), and contextual bias (Liu et al., 2024; Zhou et al., 2024).
-Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source (Li et al., 2023b; Chen et al., 2024c), proprietary commercial ones (OpenAI, 2023; Anthropic, 2024; Gemini Team, 2024), and fine-tuned models (Huang et al., 2024; Wang et al., 2023c; Zhu et al., 2023b; Li et al., 2023b; Kim et al., 2023).
-To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content.
-For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs.
-Some other studies also use the term position bias (Ko et al., 2020; Aslanyan & Porwal, 2019; Blunch, 1984; Raghubir & Valenzuela, 2006), but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
-Position bias is arguably the most prevalent and impactful bias among all.
-Chua et al.
-(2024) notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought (Wei et al., 2022) performance over self-training controls for all biases except position bias.
-Furthermore, Khan et al.
-Khan et al.
-(2024) point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
-Moreover, there is ongoing debate over whether selection bias originates from position bias.
-Pezeshkpour & Hruschka (2023) argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity.
-Raina et al.
-(2024) contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
-All of these findings underscore the critical importance of addressing the issue of position bias when employing LLM-as-a-Judge.
-B.3 Deal with Position Bias
-Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts (Zheng et al., 2024b; Raina et al., 2024; Pezeshkpour & Hruschka, 2023; Zhao et al., 2021; Zhu et al., 2023a).
-Also, LLM judges are vulnerable to attacks (Raina et al., 2024; Chen et al., 2024b; Shi et al., 2024).
-Parse data with position bias There are many ways to deal with the position bias.
-The naive way is to exclude the inconsistent judgments(i.e.
-if the LLM judge gives a positionally biased judgment on a pair or list of model responses) (Zheng et al., 2024b; Chen et al., 2024b; Wang et al., 2023b; Li et al., 2023b).
-While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem.
-Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
-To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging (Zheng et al., 2024b; Raina et al., 2024; Li et al., 2023b; Wang et al., 2023b) or regard inconsistency as a half-win or tie for relation-based judging (Zheng et al., 2024b; Li et al., 2023c) after swapping the order of model responses in the prompt.
-For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B.
-The latter way of swapping + tie is proposed because intuitively the position bias is more likely to occur when the model responses get evaluated share a similar quality in terms of the evaluating metric.
-Our quantitative study also verifies this intuition, evidenced by the fact the LLM judge’s positional consistency (the percentage of positionally consistent pairs of judgment) is positively proportional to the answer quality gap between the model responses.
-To save the expense of running the experiments more than once using swapped order, in practice, many of these studies also support a random-shuffle option in their code settings such that the baseline model for comparison does not remain in a fixed position.
-Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping (Hou et al., 2024), split-and-merge (Li et al., 2023d), and multi-agent discussion (Li et al., 2023c; Khan et al., 2024).
-However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
-Furthermore,Liu et al.
-(2024) suggests that existing calibration techniques designed to reduce bias, including context decomposition (Li et al., 2023d), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a), and batch calibration (Zhou et al., 2024), are insufficient to align LLM evaluators, even with supervised data.
-Thus, position bias is a pervasive, significantly impactful, and challenging problem to solve.
-Understand before solving We observe that the existing methods maybe ineffective or not enough satisfactory because there is a lack of understanding of the position bias.
-Although a variety of studies have researched this type of bias in LLM-as-a-Judge, a comprehensive understanding of what factors affect the position bias remains underexplored.
-In other words, without clarity on the key factors and their quantitative impact on position bias, the efficacy of current and future methods remains uncertain.
-For instance, Li et al.
-(2023d) proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%.
-However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful.
-In other words, if the quality between the two model responses differs considerably, the LLM judges will make little position bias during judgment and hence is easy to calibrate.
-This example illustrates how lacking a comprehensive understanding of the factors affecting position bias can lead to overestimating or misjudging the effectiveness of methods proposed to resolve it and improve the performance of LLM-as-a-Judge.
-Missing analysis for position bias Besides, the positional preference direction aspect of position bias remains underexplored.
-Positional preference refers to the specific positions an LLM judge favors when position bias is evident.
-For instance, if an LLM judge consistently favors the responses that appear first in the prompt, it exhibits a preference for the first position.
-Our study denotes primacy-preferred and recency-preferred to describe biases toward the first and second positions, respectively.
-We integrate the measurement by formally defining positional fairness score to quantify the extent of a judge model’s preference towards certain positions.
-While prior research includes some positional preference results (Zheng et al., 2024b; Wang et al., 2023b; Zhu et al., 2023b; Li et al., 2023c; Liusie et al., 2024), none have conducted a detailed analysis.
-They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
-Therefore, we propose that mitigating the position bias of LLM-as-a-Judge requires a simultaneous improvement in both positional consistency and positional fairness.
-A positionally consistent (high positional consistency) and positionally fair (nearly equal preference on primacy and recency when position bias occur) LLM judge is undoubtedly preferable.
-However, a trade-off between consistency and fairness/preference often occurs in practice.
-Previous studies that have provided results on positional preference have all shown that even state-of-the-art (SOTA) LLM judges struggle with both positional inconsistency and clear positional preferences.
-Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions.
-MLLM-as-a-Judge (Chen et al., 2024a) conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments.
-In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness.
-If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
-B.4 Summary of Prior Work
-To summarize, LLM-as-a-Judge has a large potential in alternating human judges on a wide range of tasks, in particular subjective evaluation, owing to its cost-effectiveness, high agreement with human judgments, reproducibility, and scalability.
-However, it suffers from various biases, most notably position bias, which is prevalent across different evaluation tasks, models, and judgment types.
-Not only is it a type of bias that significantly hinders the improvement and promotion of applying LLM-as-a-Judge, but it is also difficult to solve due to its complexity and lack of understanding in the community.
-Existing solutions proposed to address this issue may be ineffective or uncertain in efficacy due to the unclear impact and nature of influencing factors.
-Additionally, whether the position bias is an essential bias or mere random variation has not been validated from prior works.
-The positional preference side, reflecting the positional fairness of the LLM judge, also requires comprehensive analysis.
-In a nutshell, a comprehensive understanding of the position bias of LLM-as-a-Judge is crucial to validate existing and future approaches to address this important problem.
+However, position bias is a complex problem that is challenging to deal with or solve.
+The naive way is to exclude the position inconsistent judgments, which does not solve the fundamental issue and would likely result in data sparsity when position bias is frequently exhibited.
+Therefore, researchers proposed ‘inconsistency-as-a-tie’ for both models in pairwise comparative settings to consider all judgments for further analysis (Zheng et al., 2024b; Li et al., 2023c).
+This approach, while practically useful for evaluations, does not mitigate position bias.
+Given the significance of position bias, researchers have developed more sophisticated approaches to mitigate it, including bootstrapping (Hou et al., 2024), split-and-merge techniques (Li et al., 2023d), and multi-agent discussions (Li et al., 2023c; Khan et al., 2024).
+While these methods demonstrate potential, they are often costly, time-consuming, or insufficient.
+For example, multi-agent discussions and iterative reviews require extensive computational resources and human-like deliberation, while simpler techniques, such as order permutation (Wang et al., 2023; Zheng et al., 2024a) and batch calibration (Zhou et al., 2024), have limited effectiveness in addressing bias holistically.
+Furthermore, calibration methods like ensembling (Li et al., 2023a) or context decomposition (Li et al., 2023d) are often constrained by the inadequacy of supervised data to align LLM evaluators effectively (Liu et al., 2024).
+This highlights that addressing this crucial bias is complex and challenging.
+The limitations of existing approaches underscore a critical gap in our understanding of position bias itself.
+A lack of clarity on the factors influencing this bias and their quantitative impacts hampers the development of effective solutions.
+For instance, while the PORTIA approach (Li et al., 2023d) demonstrates significant improvements in consistency for GPT-4 (from 93.44% to 97.03%), this success may be overestimated due to the selection of tasks with large answer quality gaps.
+As our study reveals, such gaps inherently reduce the likelihood of position bias, making these cases easier to calibrate.
+This example underscores the necessity of first comprehensively understanding the underlying factors contributing to position bias before proposing or refining mitigation strategies.
+Without this foundational understanding, the efficacy of current and future methods remains uncertain and potentially misjudged.
+B.3 Research Gaps
+To enhance the understanding of position bias, our study builds on and extends prior work through the following dimensions.
+Repetition Stability: Previous studies on LLM-as-a-Judge often implicitly assume that single-instance judgments are sufficient to reflect the evaluating capabilities of judge models, thereby attributing observed biases entirely to systematic tendencies rather than random variations.
+These studies typically proceed directly to analyses on position bias or others without such verification.
+However, if judgments vary by chance or lack consistency across repetitions, the reliability and accuracy of downstream analyses are compromised.
+To avoid this issue, Chen et al.
+(2024) repeated evaluations multiple times and adopted the mode judgment as the representative, an approach that, while effective, is computationally expensive.
+Addressing this gap, our work introduces a method or metric for systematically evaluating the stability of LLM judgments across repetitions.
+This ensures that observed biases are systematic rather than artifacts of variability.
+By performing this validation on a sample dataset for cost-efficiency, researchers can establish a robust foundation for subsequent evaluations and analyses, ensuring the reliability of their conclusions.
+Preference Fairness: Prior studies on position bias often focus on the proportion of position consistent judgments, referred to as “position consistency”.
+Only a few have explored the preference directions of LLM evaluators in addition to consistency.
+However, traditional measurements either overly penalize highly consistent judges (Li et al., 2023c; Liusie et al., 2024) or rely on metrics whose sensitivity is highly dependent on dataset size (Zheng et al., 2024b; Zhu et al., 2023).
+To address these limitations, we proposed a refined metric, preference fairness ($PF$), which integrates the strengths of both approaches.
+This single, comprehensive metric ensures sensitivity across datasets, remains independent of the number of questions and tasks, and represents a significant improvement over previous metrics.
+Answer Quality Gap: Intuitively, candidate solutions with significant quality disparities are easier to judge, likely resulting in reduced position bias.
+This suggests that the quality gap among answers is a potentially significant factor influencing position bias.
+However, this factor has been largely overlooked in prior works, both in evaluations and in the development of mitigation strategies, underscoring our limited understanding of position bias.
+While Wang et al.
+(2023) explored the answer quality gap in score-based LLM-as-a-Judge evaluations using judgment scores, quantifying this variable in comparative settings (pairwise or list-wise) remains an open research problem.
+To address this gap, we propose the use of win rates (Zheng et al., 2024b) of candidate solutions against an expected baseline to measure the answer quality gap.
+This approach offers a practical quantitative metric for subsequent analysis and deepens our understanding of the factors driving position bias.
+List-wise Evaluations: Although LLM-as-a-Judge has been categorized into pointwise, pairwise, and list-wise paradigms (Qin et al., 2024), most research has focused on pairwise comparative assessments due to their demonstrated effectiveness (Zheng et al., 2024b; Liusie et al., 2024).
+However, this leaves a significant gap in measuring and understanding the position bias in list-wise settings, where LLMs are prompted to evaluate three or more candidate solutions simultaneously.
+With advancements in model capabilities, list-wise evaluations present a potentially more cost-effective alternative to pairwise approaches, requiring fewer comparisons while maintaining consistency and fairness.
+Our work bridges this gap by extending position bias assessments from pairwise to list-wise settings, enabling a more comprehensive evaluation of LLM-as-a-Judge across varied configurations.
+These identified gaps - spanning from the validation of judgments to the evaluation of list-wise settings - highlight the need for a more comprehensive framework to evaluate and understand position bias.
+By addressing these dimensions collectively, our work provides a foundation for a deeper understanding of the consistency, fairness, and reliability of LLM evaluators.
 Appendix C LLM Agreement Analysis
@@
 Fig.
-5(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively.
-For MTBench that utilizes the 3-option mode, we also consider the “without tie” agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions.
+4(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively.
+For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions.
 The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig.
-5(d).
+4(d).
 The heatmaps reveal clear “familial patterns” in the judgment choices of these LLM judges.
@@
 Interestingly, Claude-3.5-Sonnet and Claude-3-Sonnet do not exhibit a significantly high agreement, indicating that the upgrade from series 3 to 3.5 considerably impacts their judging capabilities.
-Gemini models exhibit rather low mutual agreement and “familial property” is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
+Gemini models exhibit rather low mutual agreement and “familial property" is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
 These patterns suggest that familial similarities, possibly stemming from analogous model sizes, training data, and strategies, influence the positional preferences of these judges.
@@
 By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement.
-For our study where twelve judges are investigated, the maximum disagreement of an MTBench is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode.
+For our study where twelve judges are investigated, the maximum disagreement of an MTBench instance is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode.
 On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
 The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig.
-5(a) and Fig.
-5(b), respectively.
+4(a) and Fig.
+4(b), respectively.
 From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks.
@@
 Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig.
-5(a) and Fig.
-5(b) exhibit a balanced distribution of instances with varied difficulty.
+4(a) and Fig.
+4(b) exhibit a balanced distribution of instances with varied difficulty.
 This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
@@
 Since the difficulty of an evaluation instance is closely related to the quality gap between the two solutions and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
-Appendix D More Discussions on Answer Quality Gap
-Intuitively, answer pairs with larger quality gaps are easier for LLM judges to evaluate, as one answer is clearly superior.
-Therefore, we expect LLM judges to exhibit weaker position bias in such cases.
-This motivates us to quantify the answer quality gap and investigate its impact on position bias.
-The quantification utilizes the overall win rate ($owr$) of a candidate model (i.e., answer provider) on the other to compute the quality gap.
-For a set of questions on a certain task and the corresponding solutions by two candidate models, an LLM judge’s evaluations would pick out the candidate winner.
-For the position consistent judgment instances, we can count the number of a candidate’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively.
-On the other hand, following the methodology introduced by Zheng et al.
-(2024b), we consider the inconsistent judgment pairs as a tie for both candidates, whose counting number is denoted as $C_{I}$.
-This follows the intuition that it is difficult to strictly pick out a better one between two responses with similar quality, hence increasing the likelihood of position bias.
-Therefore, considering both cases, we define the overall win rate ($owr$) of a candidate’s solution over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs involving the two candidate solutions in both original and swapped settings.
-Then the quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, as 0.5 is the exact tie case for the win rate ranging from 0 to 1.
-We prefer the overall win rate over the consistent win rate (calculated by $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ to mainly address the data sparsity issue.
-If the proportion of position inconsistent judgments becomes high, the available data becomes sparse and $n_{C}$ becomes small.
-In the extreme case, where all judgments are inconsistent, $n_{C}$ is 0 and the consistent win rate is not even computable, making the measurement inaccurate and impractical.
-In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie” and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
+Appendix D Open-source Model Discussion
+Due to computational constraints, we conducted limited experiments on open-source models, specifically Llama 7B-13B, as representatives of smaller-scale models.
+However, as shown in Table 2, these models struggled to perform effectively on our evaluation tasks.
+A primary issue was their limited context window, which proved insufficient for benchmarks like DevBench, where questions often include extensive instructions, such as full product requirement documents.
+This context limitation caused frequent truncation of either the prompt or the candidate responses, leading to incomplete or erroneous outputs and significantly higher error rates.
+Figure 5: Llama-13B output example: Despite providing analysis and a choice, the output format does not match the specified “[[choice]]" pattern, causing extraction failures.
+Another key factor affecting performance was mismatched output formatting.
+MTBench requires responses in the format [[choice]] (e.g., [[A]]), but Llama-13B, as shown in Fig.
+5, frequently produced outputs such as [A], which could not be parsed correctly by the regular expression used for choice extraction.
+Interestingly, Llama-7B outperformed Llama-13B on MTBench, likely due to its simpler, shorter responses that inadvertently aligned better with the expected format.
+These errors illustrate the challenges these models faced in meeting benchmark-specific requirements, highlighting nuances in their evaluation results.
+Considering practicality and effectiveness, we only focus on more capable commercial judge models.
 Appendix E More Results of Position Bias and Answer Quality Gap Measurement
+E.1 MTBench
 As shown in Fig.
-4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness.
+2(c), considering all judges together, a larger answer quality gap generally leads to better position consistency and preference fairness.
 In this session, we explore whether the discovery is consistent for each individual judge.
@@
 As shown in Fig.
-6 (a), the “parabolic shape” is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency” applies to all models.
+6 (a), the “parabolic shape" is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher position consistency applies to all models.
 However, Fig.
-6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges.
-For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape”, GPT-4 and GPT-3.5 present nearly linear curves.
-In other words, while the former models align with the general tendency that a larger answer quality gap improves positional fairness, the latter ones preserve fairness regardless of the answer quality gap.
-This further demonstrates the necessity to investigate positional fairness in addition to consistency when evaluating a judge model’s position bias.
+6 (b) reveals that preference fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges.
+For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape", GPT-4 and GPT-3.5 present nearly linear curves.
+In other words, while the former models align with the general tendency that a larger answer quality gap improves preference fairness, the latter ones preserve fairness regardless of the answer quality gap.
+This further demonstrates the necessity to investigate preference fairness in addition to consistency when evaluating a judge model’s position bias.
+E.2 DevBench
+This session includes a similar baseline comparison analysis on DevBench as on MTBench.
+As shown in Fig.7, position bias is judge-dependent and task-dependent on DevBench as well, as $PC$ and $PF$ vary significantly across judges and tasks.
+Similarly, although GPT-4 stands as the baseline model with a generally high $PC$ across tasks, certain models achieve comparable or superior performances on certain tasks.
+For instance, for architecture design evaluations, GPT-4-Turbo, GPT-4o, and Gemini-1.5-pro all surpass GPT-4.
+Gemini-1.5-pro is especially outstanding, also exceeding GPT-4 in uml class evaluations.
+However, GPT-4 is still the best-performing model on UML sequence evaluations, with only GPT-3.5-Turbo can achieve comparable performance regarding certain detailed metrics (e.g., interaction complexity).
+These discoveries, aligning with the findings on MTBench, further necessitate the need to consider the trade-offs between positional consistency and fairness when selecting the optimal judge model for certain tasks.
+Figure 7: Baseline comparisons of judges across tasks on DevBench.
+An asterisk marks the statistical significance by Student’s t-tests.
+Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$.
+The baseline model is chosen to be GPT-4 for its generally superior performances.
+Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
 Appendix F Variable Selection and Tests
@@
 For a linear regression model with independent and identically distributed (iid) errors, $N(0,\sigma^{2})$, fitted to $n$ observations, the log-likelihood can be written as:
-$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$ (5)
+$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$
 where $\hat{e}_{i}$ is the residual for the $i$th observation, and $\sigma^{2}$ is the variance of the errors. The AIC, in this context, becomes:
-$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$ (6)
+$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$
 This form of the AIC balances the goodness of fit (as reflected by the residual sum of squares) and model complexity (as represented by $k$).
@@
 We therefore conclude that, although position bias is influenced by task output length, this dependency is minimal.
-Table 3: Final results of stepwise model selection for both benchmarks: Position Consistency
 Removed Variables DF Sum of Sq RSS AIC
@@
 Judge 13 55.417 219.17 -16846
-Table 4: Final results of stepwise model selection for both benchmarks: Preference Fairness
+Table 3: Final results of stepwise model selection for both benchmarks: Position Consistency
 Removed Variables DF Sum of Sq RSS AIC
@@
 Judge 13 82.069 336.35 -14641
-Table 5: Final results of stepwise model selection for MTBench: Position Consistency
+Table 4: Final results of stepwise model selection for both benchmarks: Preference Fairness
 Removed Variables DF Sum of Sq RSS AIC
@@
 Quality gap 1 15.6206 77.594 -12559
-Table 6: Final results of stepwise model selection for MTBench: Preference Fairness
+Table 5: Final results of stepwise model selection for MTBench: Position Consistency
 Removed Variables DF Sum of Sq RSS AIC
@@
 Judge 13 58.847 187.85 -9672.5
-Table 7: Final results of stepwise model selection for DevBench: Position Consistency
+Table 6: Final results of stepwise model selection for MTBench: Preference Fairness
 Removed Variables DF Sum of Sq RSS AIC
@@
 Judge 13 84.760 140.141 -5146.6
-Table 8: Final results of stepwise model selection for DevBench: Preference Fairness
+Table 7: Final results of stepwise model selection for DevBench: Position Consistency
 Removed Variables DF Sum of Sq RSS AIC
@@
 Judge 13 80.875 140.979 -5108.9
-Appendix G DevBench
-This session includes a similar baseline comparison analysis on DevBench as on MTBench.
-As shown in Fig.7, position bias is judge-dependent and task-dependent on DevBench as well, as $PC$ and $PF$ vary significantly across judges and tasks.
-Similarly, although GPT-4 stands as the baseline model with a generally high $PC$ across tasks, certain models achieve comparable or superior performances on certain tasks.
-For instance, for architecture design evaluations, GPT-4-Turbo, GPT-4o, and Gemini-1.5-pro all surpass GPT-4.
-Gemini-1.5-pro is especially outstanding, also exceeding GPT-4 in uml class evaluations.
-However, GPT-4 is still the best-performing model on UML sequence evaluations, with only GPT-3.5-Turbo can achieve comparable performance regarding certain detailed metrics (e.g., interaction complexity).
-These discoveries, aligning with the findings on MTBench, further necessitate the need to consider the trade-offs between positional consistency and fairness when selecting the optimal judge model for certain tasks.
-Figure 7: Baseline comparisons of judges across tasks on DevBench.
-An asterisk marks the statistical significance by Student’s t-tests.
-Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$.
-The baseline model is chosen to be GPT-4 for its generally superior performances.
-Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
-Appendix H Experiment Settings
+Table 8: Final results of stepwise model selection for DevBench: Preference Fairness
+Appendix G Experiment Settings
 This session specifies more detailed information about the judges, answer-generating models, tasks, and prompt templates used in this study.
 We choose to evaluate MTBench and DevBench for the following reasons: (1) all necessary information about the benchmark models, tasks, and questions is publicly available, making modifications convenient (2) they include a wide variety of answer-generating models, tasks, and task questions for a comprehensive evaluation (3) their human evaluations validated the reliability of state-of-the-art judging models (GPT-4 and GPT-4-Turbo) on their evaluation instances, hence model untested by prior work, if reaching high agreement with these validated judges, can be perceived reliable as well.
-H.1 Judges, Candidates, and Tasks
+G.1 Judges, Candidates, and Tasks
 Judge
@@
 They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations.
-The other models that are compared to the reference models, namely “Model” in our context, are listed as follows.
+The other models that are compared to the reference models, namely “Model" in our context, are listed as follows.
   • MTBench (30): alpaca-13b, baize-v2-13b, chatglm-6b, claude-instant-v1, claude-v1, dolly-v2-12b, falcon-40b-instruct, fastchat-t5-3b, gpt-3.5-turbo, gpt-4, gpt4all-13b-snoozy, guanaco-33b, guanaco-65b, h2ogpt-oasst-open-llama-13b, koala-13b, llama-13b, mpt-30b-chat, mpt-30b-instruct, mpt-7b-chat, nous-hermes-13b, oasst-sft-4-pythia-12b, oasst-sft-7-llama-30b, palm-2-chat-bison-001, rwkv-4-raven-14b, stablelm-tuned-alpha-7b, tulu-30b, vicuna-33b-v1.3, vicuna-7b-v1.3, wizardlm-13b, wizardlm-30b
@@
 (2024b) and DevBench Li et al.
-(2024) used in their studies.
+(2024a) used in their studies.
 That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4).
@@
     • architecture design (5): conformance, design_and_coding, practicability, uniformity_and_integration, and faithfulness
-H.2 Prompt Settings
+G.2 Prompt Settings
 We follow the original prompt settings of MTBench and DevBench in our study of pairwise comparative LLM-as-a-Judge.
@@
   • A system prompt explaining the judging task and the role the LLM should be playing.
-  • Emphasized “should” and“shouldn’t”s.
+  • Emphasized “should" and“shouldn’t"s.
   • A prompt structure with placeholders for specific questions and model answers
@@
 The detailed prompt templates are specified below.
-Appendix I Reproducibility
-Our experiments mainly depend on API. Code will be released upon acceptance.

### v8

@@
 LLM-as-a-Judge has emerged as a promising alternative to human evaluators across various tasks, yet inherent biases—particularly position bias, the tendency to favor solutions based on their position within the prompt—compromise its reliability.
-This study investigates position bias in LLM judges across pairwise and list-wise comparison settings, introducing three metrics: repetition stability, position consistency, and preference fairness.
-Our experiments, involving 12 LLM judges across MTBench and DevBench with 22 tasks and approximately 40 solution-generating models, result in over 100,000 evaluation instances.
+This exploratory study evaluates position bias in LLM judges across pairwise and list-wise comparison settings, introducing three metrics: repetition stability, position consistency, and preference fairness.
+Our experiments, involving 15 LLM judges across MTBench and DevBench with 22 tasks and approximately 40 solution-generating models, result in over 150,000 evaluation instances.
 We identify Judge-Level, Candidate-Level, and Task-Level factors contributing to bias.
@@
 Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge
-Lin Shi, Chiyu Ma, Wenhua Liang, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
+Lin Shi, Chiyu Ma, Wenhua Liang, Xingjian Diao, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, xingjian.diao.gr, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
 1 Introduction
@@
 Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields.
-While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations Zeng et al.
+While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability in extensive evaluations Zeng et al.
 (2023); Karpinska et al.
@@
 (2024), the introduction of new biases Ohi et al.
-(2024), or inconsistent effectiveness Gallegos et al.
+(2024), inconsistent effectiveness Gallegos et al.
+(2024), or impracticality for closed-source models, as mechanistic solutions require modifications to model internals Wang et al.
+(2025); Yu et al.
 (2025).
-These persistent challenges highlight the limitations of current approaches and point to a possible incomplete understanding of position bias.
-Although numerous studies have evaluated position bias in LLM-as-a-Judge settings, the underlying factors contributing to this phenomenon remain insufficiently explored.
+From an exploratory perspectives, despite empirical studies evaluating position bias in pairwise LLM-as-a-Judge settings Zheng et al.
+(2024b); Wang et al.
+(2023), the underlying factors contributing to this phenomenon and extended measurements in list-wise comparative paradigms remain incompletely explored.
 In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge.
-While evaluating Position Consistency, one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability.
+While evaluating Position Consistency Zheng et al.
+(2024b), one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability.
 Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences.
@@
 Finally, we provide a detailed description of the factors we found that are related to position bias at the Judge-Level, Candidate-Level, and Task-Level.
+Our exploratory study was conducted post hoc, meaning the LLM judgments were collected first, and the factors influencing position bias were then identified and analyzed.
 2.1 Pairwise & List-wise Comparison
@@
 The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question.
-Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair.
+Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same LLM judge again, which results in a judgment pair.
 If the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent.
@@
 1 (b).
-To accommodate the possibility of ties, various option modes are employed: Two-Option mode restricts judges to choosing between two options, labeled A for the first candidate and B for the second.
-Three-Option mode introduces an additional choice, C, allowing judges to indicate a tie if neither solution is preferable, as illustrated in Fig.
+To accommodate the possibility of ties, various option modes are employed: Two-Option mode restricts LLM judges to choosing between two options, labeled A for the first candidate and B for the second.
+Three-Option mode introduces an additional choice, C, allowing LLM judges to indicate a tie if neither solution is preferable, as illustrated in Fig.
 1 (a).
@@
 1 (c).
-For efficiency, we prompt judges to select the best candidate rather than ranking the entire list.
+For efficiency, we prompt LLM judges to select the best candidate rather than ranking the entire list.
 The “swapped setting" used in pairwise evaluations is generalized to order permutations for list-wise judgments, ensuring that each candidate appears in every possible position exactly once.
@@
 In the $i^{th}$ permutation, the $i^{th}$ candidate is set to appear in the first position.
-Additionally, an option is provided to account for ties, allowing judges to indicate if there is no certainly superior solution.
+Additionally, an option is provided to account for ties, allowing LLM judges to indicate if there is no certainly superior solution.
 2.2 Evaluation Metrics
-In our study, we first verify if the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness.
+In our study, we first verify whether capable LLM judges exhibit high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness.
 The metrics are introduced as follows:
@@
 The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
-Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is permuted.
-It is calculated as the ratio of consistent evaluation series to the total number of valid evaluations, where a series is deemed consistent if the judge model prefers the same winning solution across permutations.
+Position Consistency ($PC$) quantifies how frequently LLM judges prefer the same solution after the order of solutions is permuted.
+It is calculated as the ratio of consistent evaluation series to the total number of valid evaluations, where a series is deemed consistent if the LLM judge prefers the same winning solution across permutations.
 Formally, it is calculated as
@@
 1 (b).
-This formula aims to provide a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
-Preference Fairness ($PF$) measures the extent to which judge models favor certain solution positions.
+This formula aims to provide a direct measure of a LLM judge’s position bias and has been widely used in previous studies for its simplicity.
+Preference Fairness ($PF$) measures the extent to which LLM judges favor certain solution positions.
 In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency.
-These terms replace the more verbose “preference for the first/second candidate model" used in previous studies, ensuring clarity and generalization for future research.
+These terms replace the more verbose “preference/bias for the first/second candidate model" used in previous studies Zheng et al.
+(2024b), ensuring clarity and generalization for future research.
 The examples of such preferences are demonstrated in Fig.
@@
 2.3 Factors Affecting Position Bias
-Factor Judge-level Candidate-level Task-level
-Familial Property $\checkmark$* $\times$ $\times$
-Answer Quality Gap $\times$ $\checkmark$* $\times$
-Task Input Length $\times$ $\times$ $\checkmark$
-Task Output Length $\times$ $\checkmark$* $\checkmark$*
-Prompt Length $\times$ $\checkmark$ $\checkmark$
-Table 1: Factors influencing position bias.
-Significant factors, identified via bidirectional stepwise regression, are marked with * and highlighted in red based on empirical findings.
-Task Input refers to the question itself, while Task Output denotes the candidate model’s answers, serving as both Candidate-level and Task-level factors.
-Prompt includes the full query presented to judges: Task Input, Task Output, and system prompts.
 To investigate the factors influencing position bias in LLM judges, we categorized these factors into three groups: Judge-level, Candidate-level, and Task-level factors.
@@
 Table 1 lists the five factors we analyzed in this study.
-Among the influencing factors, we selected “familial property” for Judge-level factors, as it reflects similar model sizes and training specifics, which are often proprietary and not publicly accessible.
-The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version.
+Among the influencing factors, we selected “familial property” for Judge-level factors, as it reflects similar model sizes or training specifics, which are often proprietary and not publicly accessible for closed-source capable models.
+The familial categories of the models used in our studies are (1) GPT, (2) Claude, (3) Gemini, and (4) Llama allowing for straightforward grouping by company and version.
 More details and discussions about the familial property can be found in Appendix Sec.
 C.
-Answer quality gap, based on our findings, is highly influential yet under-explored in prior research.
+Answer quality gap: While prior studies Wang et al.
+(2023) explored quality disparities using "score gaps" in score-based LLM-as-a-Judge, this factor remains under-explored in comparative settings, which we address by introducing "answer quality gap" for both pairwise and list-wise evaluations.
 We define the quality of a candidate’s solution by how effectively it addresses the question.
 Consequently, the answer quality gap refers to the disparity in quality between the solutions from one candidate model and the others to the same question and hence considered the Candidate-level factor.
-Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
+Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the LLM judge.
 Following this assumption, we measure the answer quality gap by the win rates of candidates over an expected baseline on a set of tasks and questions.
@@
 3.1 Experiment Settings
-In this study, we evaluated twelve up-to-date close-sourced models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework.
-Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks.
-However, due to their limited context window size and failure to follow the output format, which led to suboptimal performance on our tasks (as shown in Table 2), we focused our investigation exclusively on more powerful closed-source models.
-We included the results of open-source models for completeness.
-More discussions can be found in Appendix.Sec.
-D
-We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks.
-For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
+In this study, we evaluated position bias of 15 models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), Gemini (Gemini Team, 2024), and Llama (Touvron et al., 2023) series using our framework.
+For datasets, we adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) due to their demonstrated high human-LLM agreement and the validated reliability of state-of-the-art LLM judges on the evaluation tasks.
+For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of solutions to the given questions.
 MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task.
@@
 We adopted Two-option mode for MTBench, and Three-option mode for DevBench.
-For list-wise experiments, we randomly sampled 9 models to form three triple-candidate lists and evaluated four representative judges on MTBench.
+For list-wise experiments, we randomly sampled 9 models to form three triple-candidate lists and evaluated four representative LLM judges on MTBench.
 The prompt templates we used are identical to those in the benchmarks for pairwise comparisons, with minor modifications to accommodate list-wise evaluations.
@@
 Sec.
-G.
-To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials.
+F.
+To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each LLM judge to evaluate across 3 repetitive trials.
 This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench.
-The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results.
+The temperature hyperparameter was set to 1 for all LLM judges to generate nontrivial results.
 To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets.
@@
 Sec.
-F.
+E.
+Factor Judge-level Candidate-level Task-level
+Familial Property $\checkmark$* $\times$ $\times$
+Answer Quality Gap $\times$ $\checkmark$* $\times$
+Task Input Length $\times$ $\times$ $\checkmark$
+Task Output Length $\times$ $\checkmark$* $\checkmark$*
+Prompt Length $\times$ $\checkmark$ $\checkmark$
+Table 1: Factors influencing position bias.
+Significant factors, identified via bidirectional stepwise regression, are marked with * and highlighted in red based on empirical findings on both MTBench and DevBench results.
+Task Input refers to the question itself, while Task Output denotes the candidate model’s answers, serving as both Candidate-level and Task-level factors.
+Prompt includes the full query presented to LLM judges: Task Input, Task Output, and system prompts.
 3.2 Empirical Results
-The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2.
+The evaluation results of 12 close-source and 3 open-source LLM judges in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2.
 For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks.
@@
 These analyses were conducted by considering all judges together on MTBench.
-The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E.
-Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables.
+More analyses can be found in Appendix.
+Sec.
+D.
+Through bidirectional stepwise regression, as shown in Table 1, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables.
 Similarly, these factors also contribute significantly to Preference Fairness.
@@
 Sec.
-F.
+E.
 Judge MTBench Pairwise DevBench Pairwise MTBench List-wise
@@
 Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 0.55 $\pm$ 0.20 0.33 $\pm$ 0.18 0.00
+Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
 Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96
-Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
+GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
+GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
 GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 0.68 $\pm$ 0.22 0.18 $\pm$ 0.22 0.00
+GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 0.34 $\pm$ 0.17 -0.05 $\pm$ 0.30 0.12
 o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00
-GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
-GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
-GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 0.34 $\pm$ 0.17 -0.05 $\pm$ 0.30 0.12
-Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
-Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
-Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
+Llama-3.3-70B 0.96 $\pm$ 0.06 0.80 $\pm$ 0.16 -0.05 0.00 0.99 $\pm$ 0.01 0.89 $\pm$ 0.12 -0.03 0.00
+Llama-3.1-405B 0.93 $\pm$ 0.10 0.77 $\pm$ 0.16 0.10 0.02 0.94 $\pm$ 0.10 0.79 $\pm$ 0.15 0.01 0.00
+Llama-3.1-8B 0.75 $\pm$ 0.32 0.69 $\pm$ 0.23 -0.03 0.25 0.79 $\pm$ 0.36 0.47 $\pm$ 0.18 0.25 0.00
 Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold.
@@
 High error rates and low $RS$ are marked red, rendering further evaluations invalid due to insufficient data.
-Open-source models (e.g., Llama) are excluded due to high error rates with our lengthy instances (See Appendix.Sec.D for more discussions).
-List-wise evaluation is conducted on four representative judges from each family to validate scalability.
+List-wise evaluation is conducted on four representative judges to validate scalability.
+4 Main Findings
 Figure 2: Judge performances on MTBench.
 Fig.
-(a)(b) are the radar charts for the $PC$ comparison by family, judge, and task Fig.
-(c) to (f) investigate the impact of task output length and answer quality gap on position bias.
+(a)(b) are the radar charts for the $PC$ comparison by family, judge, and task.
 Fig.
-(a)(c)(d)(e) are for pairwise comparative settings, while Fig.
-(b)(f) are obtained under list-wise evaluations.
-4 Main Findings
+(c) leverages linear regression to explore the general relationship between $PC$ and $PF$.
+Fig.
+(d) to (f) investigate the impact of answer quality gap on position bias using overall win rates.
+Fig.
+(a) to (e) are for pairwise comparative settings, while Fig.
+(f) are obtained under list-wise evaluations.
 Position Bias of Capable Judges are not Mere Random Variations: As shown in Table 2, the capable judges on the benchmark tasks, supported by minimal "Error" rates, generally exhibit $RS$ values above 0.85.
-The most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks.
-These results confirm that LLM judgments, and the resulting position bias, are not due to random variations.
+The most capable models, such as Claude-3.5-Sonnet, GPT-4, and Llama-3.3-70B, all achieve near-perfect $RS$ scores over 0.95 on both benchmarks.
+These results confirm that judgments from capable LLM judges, and the resulting position bias, are not due to random variations.
 This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
-Position Bias Varies by Judge & Task: As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig.
-2 (a) for $PC$ comparisons.
-Moreover, certain models achieve comparable or superior performances than GPT-4 on certain tasks.
-For example, according to Fig.
-2 (a), for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges.
-Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it could be employed as a cost-effective alternative to coding evaluations.
-Additionally, significant variations of $PF$ across tasks are observed.
-The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others.
-Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, role play, and writing tasks.
-Even for judges that are recency-preferred across all tasks (e.g., Claude-3’s and Gemini-pro’s), the extent of biased preference, as reflected by $PF$ values, varies by task.
-Moreover, a high position consistency does not guarantee fairness.
-For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively.
-In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
-In list-wise comparisons, similar variations by judge and task in position bias were observed, as shown in Fig.
-2 (b).
+Position Bias Varies by Judge & Task: As shown in Fig.
+2(a), position bias among LLM judges varies significantly across different judges and tasks.
+For instance, GPT-4o demonstrates higher position consistency when evaluating coding tasks but performs less consistently on other tasks compared to GPT-4.
+Similarly, Gemini-1.5-pro, while achieving higher $PC$ than other Gemini models in most tasks, exhibits comparable consistency when judging extraction tasks.
+Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec.
+D.2.
+Variations in preference fairness are also evident.
+As shown in Table 2, GPT-4 and GPT-3.5-Turbo display different preference biases across datasets and tasks—being recency-preferred on MTBench but primacy-preferred on DevBench.
+Likewise, Claude-3.5-Sonnet, which is nearly preference-fair on MTBench ($PF=0.01$), exhibits a strong recency-preferred position bias on DevBench ($PF=0.22$).
+While higher position consistency generally correlates with improved preference fairness (as demonstrated by the regression curve in Fig.
+2(c)), consistency alone does not guarantee fairness.
+Certain LLM judges, despite achieving high $PC$, still exhibit significant and varied preference directions across different tasks, underscoring the need to evaluate both consistency and fairness when assessing LLM judges.
+In list-wise comparisons, similar variations in position bias were observed across judges and tasks.
 Furthermore, Table 2 highlights that more capable models, such as GPT-4o and Claude-3.5-Sonnet, maintain high consistency when transitioning from pairwise to list-wise evaluations, while less capable models, such as GPT-3.5-Turbo, exhibit greater sensitivity to the increased number of candidates in list-wise tasks.
-Therefore, the position bias of LLM judges is judge-dependent and task-dependent.
+Therefore, the position bias of LLM judges is both judge-dependent and task-dependent.
 This observation is further confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$.
-In practice, to achieve the most reliable evaluations, one needs to select task-specific LLMs whose consistency and fairness are in balance.
-Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec.
-E.2.
+In practice, when evaluating LLM judge’s reliability or selecting suitable LLM judges, considering the balance between consistency and fairness, as well as accounting for task-specific variations, may be beneficial.
 Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
@@
 2 (e) and (f) exhibit significant parabolic shapes, indicating that $PC$ is positively proportional to $\delta_{q}$.
-This aligns with our intuition that the answers with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$.
+This aligns with our intuition that the answer pairs or series with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$.
 The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.
-E.
+D.
 Similarly, as shown in Fig.
@@
 For example, as shown in Appendix.Fig.
-6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
+5, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
 These observations align with the right-arrow shape as demonstrated in Fig.
@@
 However, exceptions are common, as shown by the individual scatter points of these figures.
-This indicates that though the answer quality gap significantly influences the position bias of LLMs, other factors also play important roles.
+This indicates that though the answer quality gap significantly influences the position bias of LLM judges, other factors also play important roles.
 Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
@@
 LLM Agreement Analysis: We complement our investigation of position bias with an agreement and disagreement analysis among LLM judges.
-Rather than focusing exclusively on overall consistency or fairness, we examine how judges converge and diverge in their assessments of individual instances.
-Agreement analysis quantifies the percentage of instances where two judges mutually agree on the outcome.
+Rather than focusing exclusively on overall consistency or fairness, we examine how LLM judges converge and diverge in their assessments of individual instances.
+Agreement analysis quantifies the percentage of instances where two LLM judges mutually agree on the outcome.
 Disagreement analysis counts the number of choices deviating from the mode for each instance among all judges.
@@
 For example, as shown in Fig.
-3, more than half of the dataset can be considered relatively easy to judge, as over 75% of all 12 judges agree with each other on these instances.
-Conversely, fewer than 4% of instances represent the likely hard-to-judge cases where a majority of judges fail to reach consensus.
-Figure 3: Distribution of disagreement on MTBench. The y-axis indicates the proportion of the dataset where the level of disagreement among judges does not exceed a specific threshold.
+3, more than half of the dataset can be considered relatively easy to judge, as over 80% of all 15 LLM judges agree with each other on these instances (disagreement $\leq$ 3).
+Conversely, fewer than 2% of instances represent the likely especially-hard-to-judge cases where a majority of LLM judges fail to reach consensus (disagreement $\geq$ 8).
+Figure 3: Distribution of disagreement on MTBench. The y-axis indicates the proportion of the dataset where the level of disagreement among LLM judges does not exceed a specific threshold.
 Based on our observations of answer quality gaps and LLM agreement/disagreement patterns, this study offers practical insights for designing evaluator benchmarks that account for the varying difficulty levels of judgment tasks.
-Specifically, the most challenging instances to evaluate are characterized by: (1) frequent disagreements among LLM judges, (2) closely matched win rates and minimal quality gaps among candidate models, and (3) significant position bias exhibited by the majority of judges.
+Specifically, the most challenging instances to evaluate are characterized by: (1) frequent disagreements among LLM judges, (2) closely matched win rates and minimal quality gaps among candidate models, and (3) significant position bias exhibited by the majority of LLM judges.
 Further discussions and analyses can be found in Appendix Sec.
@@
 Despite proposing scalable metrics and investigating key factors influencing position bias, our study has several limitations.
-First, due to computational constraints, we evaluated only twelve commercial LLM judges for pairwise settings and four for list-wise settings, limited to 3-candidate lists, using two benchmarks.
-Open-source models, such as Llama 7B-13B, struggled to provide reliable judgments, restricting their inclusion.
-Additionally, while we tested original benchmark prompt templates, exploring variations in prompting techniques could provide further insights.
-Future work could expand this study to include more models, tasks, prompting strategies, and larger list-wise candidate pools to validate and generalize our findings.
-Second, data accessibility limitations prevented a direct analysis of Judge-level factors like architecture and parameter size.
+First, we evaluated only 12 commercial closed-source LLM judges for pairwise settings and 4 for list-wise paradigms across two benchmarks, limiting the list-wise comparisons to three-candidate lists.
+Among open-source models, we assessed only three Llama 3.1+ models of varying sizes, as earlier versions or other smaller models lacked sufficient context window lengths, making them unsuitable given the length of evaluation instances in our study.
+Additionally, while we used original benchmark prompt templates, exploring alternative prompting techniques could offer further insights.
+Future work could expand on this by incorporating more models, tasks, prompting strategies, and larger list-wise candidate pools to enhance the generalizability of our findings.
+Second, data accessibility limitations prevented a direct analysis of Judge-level factors like architecture and parameter size of closed-source models.
 Instead, we approximated these factors by grouping models by family properties.
-Open-source models with accessible architectural details could allow deeper analysis of these influences.
-Additionally, our analyses rely on completed judgments, such as measuring the answer quality gap and task difficulty after evaluations.
+While open-source models offer accessible architectural details for deeper analysis, our assessment of only three Llama models may not provide sufficient evidence for broader conclusions.
+Additionally, our analyses were conducted post hoc, relying on completed judgments before analysis.
 Future work could explore methods to estimate or control these factors pre-judgment, reducing computational costs and enabling proactive mitigation strategies.
-Lastly, our focus was on understanding position bias rather than mitigating it.
+Lastly, our focus was on evaluating and understanding position bias rather than mitigating it.
 While our findings provide a foundation for effective mitigation, further research is needed to address issues like maintaining consistency and fairness when answer quality gaps are minimal, where position bias is most pronounced.
@@
 Our experiments were conducted primarily using API access, with a total cost of approximately 3,000 USD.
-The code repository for reproducing our results is available at: https://anonymous.4open.science/r/Position-Bias-Analyzer-Demo-F7E3
+The code repository for reproducing our results is available:https://anonymous.4open.science/r/Position-Bias-Analyzer-Demo-F7E3
 Appendix B Related Work
@@
 (2) how do the choices of all judges on an instance vary?
-Figure 4: Disagreement Analysis and Mutual Agreement Heatmap of LLM judges on MTBench and DevBench.
-The blue bars of Figures (a) and (b) represent the number of instances with certain disagreement values, and the red cumulative curve displays the cumulative count and corresponding percentage over the benchmark dataset.
-The heat maps of Figures (c)(d)(e) marked higher mutual agreement values with a brighter color.
+Figure 4: Figures (a) to (c) are mutual agreement heatmap of LLM judges on MTBench and DevBench, where (b) is the agreement computation excluding the tie option {C}.
+Higher mutual agreement between two LLM judges is marked with brighter color.
+Figure (d), like Figure 3, is the distribution of disagreement on DevBench.
 C.1 Mutual Agreement & Familial Property
@@
 We denote two judges agree on an instance if their judgment choices are identical.
-Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances.
-Fig.
-4(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively.
+Then the mutual agreement between two LLM judges on a benchmark is defined as the proportion of their agreed instances.
+Figs.
+4(a) and (c) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively.
 For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions.
 The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig.
-4(d).
+4(b).
 The heatmaps reveal clear “familial patterns” in the judgment choices of these LLM judges.
@@
 Gemini models exhibit rather low mutual agreement and “familial property" is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
+Llama models demonstrate a high agreement among capable family members (Llama-3.3-70B and Llama-405B) and with GPT series.
+However, significantly smaller and less capable model like Llama-3.1-8B does not strong agree with them.
 These patterns suggest that familial similarities, possibly stemming from analogous model sizes, training data, and strategies, influence the positional preferences of these judges.
@@
 By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement.
-For our study where twelve judges are investigated, the maximum disagreement of an MTBench instance is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode.
-On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
+For our study where 15 judges are investigated, the maximum disagreement of an MTBench instance is 10, accounting for the 5{A}-5{B}-5{C} choice pattern by 3-option mode.
+On the other hand, for DevBench instances, since Gemini-1.5-flash is excluded due to insufficient data (as shown by high error rates in Tabel 2), the maximum possible disagreement for the remaining 14 judges is 7, representing the 7{A}-7{B} judgment distribution for the 2-option mode.
 The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig.
-4(a) and Fig.
-4(b), respectively.
+3 and Fig.
+4(d), respectively.
 From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks.
@@
 In comparison, the instances with the highest disagreement are likely the ones that are difficult to evaluate and where the position bias is most likely to occur.
-However, luckily, these instances are rare, occupying only 0.03% and 0.07% of MTBench and DevBench respectively.
-In other words, majority voting of multiple capable LLM judges could be practically useful for over 99% of evaluation instances on both benchmarks.
+However, luckily, these instances are rare, occupying only less than 5% for both benchmarks respectively.
+In other words, majority voting of multiple capable LLM judges could be practically useful for over 95% of evaluation instances on both benchmarks.
 Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig.
-4(a) and Fig.
-4(b) exhibit a balanced distribution of instances with varied difficulty.
+3 and Fig.
+4(d) exhibit a balanced distribution of instances with varied difficulty.
 This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
@@
 Since the difficulty of an evaluation instance is closely related to the quality gap between the two solutions and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
-Appendix D Open-source Model Discussion
-Due to computational constraints, we conducted limited experiments on open-source models, specifically Llama 7B-13B, as representatives of smaller-scale models.
-However, as shown in Table 2, these models struggled to perform effectively on our evaluation tasks.
-A primary issue was their limited context window, which proved insufficient for benchmarks like DevBench, where questions often include extensive instructions, such as full product requirement documents.
-This context limitation caused frequent truncation of either the prompt or the candidate responses, leading to incomplete or erroneous outputs and significantly higher error rates.
-Figure 5: Llama-13B output example: Despite providing analysis and a choice, the output format does not match the specified “[[choice]]" pattern, causing extraction failures.
-Another key factor affecting performance was mismatched output formatting.
-MTBench requires responses in the format [[choice]] (e.g., [[A]]), but Llama-13B, as shown in Fig.
-5, frequently produced outputs such as [A], which could not be parsed correctly by the regular expression used for choice extraction.
-Interestingly, Llama-7B outperformed Llama-13B on MTBench, likely due to its simpler, shorter responses that inadvertently aligned better with the expected format.
-These errors illustrate the challenges these models faced in meeting benchmark-specific requirements, highlighting nuances in their evaluation results.
-Considering practicality and effectiveness, we only focus on more capable commercial judge models.
-Appendix E More Results of Position Bias and Answer Quality Gap Measurement
-E.1 MTBench
+Appendix D More Results of Position Bias and Answer Quality Gap Measurement
+D.1 MTBench
 As shown in Fig.
@@
 This further demonstrates the necessity to investigate preference fairness in addition to consistency when evaluating a judge model’s position bias.
-E.2 DevBench
+D.2 DevBench
 This session includes a similar baseline comparison analysis on DevBench as on MTBench.
@@
 These discoveries, aligning with the findings on MTBench, further necessitate the need to consider the trade-offs between positional consistency and fairness when selecting the optimal judge model for certain tasks.
-Figure 7: Baseline comparisons of judges across tasks on DevBench.
+Figure 6: Baseline comparisons of position bias of LLM judges across tasks on DevBench.
 An asterisk marks the statistical significance by Student’s t-tests.
-Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$.
-The baseline model is chosen to be GPT-4 for its generally superior performances.
-Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
-Appendix F Variable Selection and Tests
-F.1 Bidirectional Stepwise Regression with AIC
+Figure (a) selects GPT-4 as the baseline, where asterisk demonstrates signficantly better or worse $PC$ of other models compared to it.
+Figure (b) utilizes an absolute $PF$ baseline of 0 and depicts preference fairness performances of LLM judges across tasks.
+Similar to findings on MTBench, position bias is significantly judge-dependent and task-dependent on DevBench as well.
+Appendix E Variable Selection and Tests
+E.1 Bidirectional Stepwise Regression with AIC
 Bidirectional stepwise regression is a combination of forward selection and backward elimination techniques.
@@
 The process terminates when neither adding nor removing variables results in a lower AIC, signifying that the most parsimonious model, based on AIC, has been reached.
-F.2 Test results
+E.2 Test results
 We operated bidirectional stepwise regression on both benchmarks individually and together to identify the factors that are significantly contributing to position bias.
@@
 Table 8: Final results of stepwise model selection for DevBench: Preference Fairness
-Appendix G Experiment Settings
+Appendix F Experiment Settings
 This session specifies more detailed information about the judges, answer-generating models, tasks, and prompt templates used in this study.
 We choose to evaluate MTBench and DevBench for the following reasons: (1) all necessary information about the benchmark models, tasks, and questions is publicly available, making modifications convenient (2) they include a wide variety of answer-generating models, tasks, and task questions for a comprehensive evaluation (3) their human evaluations validated the reliability of state-of-the-art judging models (GPT-4 and GPT-4-Turbo) on their evaluation instances, hence model untested by prior work, if reaching high agreement with these validated judges, can be perceived reliable as well.
-G.1 Judges, Candidates, and Tasks
+F.1 Judges, Candidates, and Tasks
 Judge
@@
     • architecture design (5): conformance, design_and_coding, practicability, uniformity_and_integration, and faithfulness
-G.2 Prompt Settings
+F.2 Prompt Settings
 We follow the original prompt settings of MTBench and DevBench in our study of pairwise comparative LLM-as-a-Judge.

### v9

@@
 Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge
-Lin Shi, Chiyu Ma, Wenhua Liang, Xingjian Diao, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, xingjian.diao.gr, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
+Lin Shi, Chiyu Ma, Wenhua Liang, Xingjian Diao, Weicheng Ma, Soroush Vosoughi Dartmouth College lin.shi.26@dartmouth.edu
 1 Introduction
@@
 Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields.
-While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability in extensive evaluations Zeng et al.
-(2023); Karpinska et al.
-(2021).
+While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability in extensive evaluations zeng2023evaluating; karpinska2021perils.
 To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks.
-Despite a high level of agreement with human judgments Zheng et al.
-(2024b); Li et al.
-(2024a); Zhu et al.
-(2023), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
-Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig.
+Despite a high level of agreement with human judgments zheng2024judging; li2024devbench; zhu2023judgelm, inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
+Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than evaluating content objectively, as shown in Fig.
 1 (a).
-This bias has been prevalent in all types of LLM judges Qin et al.
-(2024); Li et al.
-(2023d), raising concerns about their reliability.
-Previous studies Zheng et al.
-(2024a, b); Zeng et al.
-(2023) have identified position bias alongside other biases and assessed its impact.
-Although mitigation strategies have been proposed, they often fall short due to incomplete bias elimination Guo et al.
-(2024), increased complexity Li et al.
-(2024b); Khan et al.
-(2024); Chua et al.
-(2024), the introduction of new biases Ohi et al.
-(2024), inconsistent effectiveness Gallegos et al.
-(2024), or impracticality for closed-source models, as mechanistic solutions require modifications to model internals Wang et al.
-(2025); Yu et al.
-(2025).
-From an exploratory perspectives, despite empirical studies evaluating position bias in pairwise LLM-as-a-Judge settings Zheng et al.
-(2024b); Wang et al.
-(2023), the underlying factors contributing to this phenomenon and extended measurements in list-wise comparative paradigms remain incompletely explored.
+This bias has been observed across various types of LLM judges qin2024large; li2023split, raising concerns about their reliability.
+Prior studies have identified position bias alongside other biases and assessed its impact zheng2024large; zheng2024judging; zeng2023evaluating, but these investigations remain largely preliminary and lack a focused, systematic exploration.
+Although mitigation strategies have been proposed, they often suffer from incomplete bias removal guo2024bias, added complexity li2024split; khan2024debating; chua2024biasaugmented, the introduction of new biases ohi2024likelihood, inconsistent effectiveness gallegos2024bias, or impracticality for closed-source models requiring access to model internals wang2025eliminating; yu2025mitigate.
+Moreover, most empirical studies primarily examine position bias in pairwise settings wang2023large; zheng2024judging, leaving its underlying factors and behavior in more complex list-wise paradigms underexplored.
+These gaps underscore the need for a more systematic, in-depth exploratory analysis of position bias to better understand its origins, manifestations, and implications.
 In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge.
-While evaluating Position Consistency Zheng et al.
-(2024b), one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability.
+While evaluating with Position Consistency zheng2024judging, we further introduce two novel metrics: Preference Fairness and Repetition Stability.
 Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences.
@@
 In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency.
-These terms replace the more verbose “preference/bias for the first/second candidate model" used in previous studies Zheng et al.
-(2024b), ensuring clarity and generalization for future research.
+These terms replace the more verbose “preference/bias for the first/second candidate model" used in previous studies zheng2024judging, ensuring clarity and generalization for future research.
 The examples of such preferences are demonstrated in Fig.
@@
 Previous studies proposed two common ways to measure the preference fairness.
-One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$).
-The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023).
+One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number (pcnpcn) and recency-count-number (rcnrcn).
+The counts are then normalized by the total number of prompt pairs (zheng2024judging; zhu2023judgelm).
 However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
-Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently.
+Alternatively, instead of normalizing over the complete dataset, studies like (li2023prd; liusie2024llm) treat position inconsistent evaluation instances independently.
 They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs.
-We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$.
+We denote these as inconsistent primacy rates (ipripr) and inconsistent recency rates (irrirr), where $ipr+irr=1$.
 However, this approach overlooks the fact that “position consistent judgments are also preference fair”, which leads to overly penalizing highly consistent LLM-judges.
@@
 The $PF$ score is interpreted as follows:
-$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
+[…]
 To extend this metric to list-wise comparisons, we employed a ‘one vs.
@@
 C.
-Answer quality gap: While prior studies Wang et al.
-(2023) explored quality disparities using "score gaps" in score-based LLM-as-a-Judge, this factor remains under-explored in comparative settings, which we address by introducing "answer quality gap" for both pairwise and list-wise evaluations.
+Answer quality gap: While prior studies wang2023large explored quality disparities using "score gaps" in score-based LLM-as-a-Judge, this factor remains under-explored in comparative settings, which we address by introducing "answer quality gap" for both pairwise and list-wise evaluations.
 We define the quality of a candidate’s solution by how effectively it addresses the question.
@@
 We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment series ($C_{I}$), respectively.
-Inspired by Zheng et al., we count inconsistent judgment pairs as ties for all candidate models, which is later calculated as a down-scaled win rate depending on the number of candidate models.
-To calculate the win rates of candidate models for all three cases, we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{p}(C_{t}+C_{I})}]$, where we have $p$ candidates in the list and $n$ judgment series.
+Inspired by zheng2024judging, we count inconsistent judgment pairs as ties for all candidate models, which is later calculated as a down-scaled win rate depending on the number of candidate models.
+To calculate the win rates of candidate models for all three cases, we define the overall win rate (owrowr) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{p}(C_{t}+C_{I})}]$, where we have $p$ candidates in the list and $n$ judgment series.
 Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-1/p|$, where $1/p$ is the expected baseline when all judgments are “ties”.
-In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment series) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
+In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment series) to quantify $\delta_{q}$ (zheng2024judging; li2023generative; raina2024llmasajudge), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
 3 Experiment
 3.1 Experiment Settings
-In this study, we evaluated position bias of 15 models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), Gemini (Gemini Team, 2024), and Llama (Touvron et al., 2023) series using our framework.
-For datasets, we adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) due to their demonstrated high human-LLM agreement and the validated reliability of state-of-the-art LLM judges on the evaluation tasks.
+In this study, we evaluated position bias of 15 models from the GPT (openai2024gpt4), Claude (anthropic2024claude3), Gemini (geminiteam2024gemini), and Llama (Llama2023) series using our framework.
+For datasets, we adopted the modified MTBench (zheng2024judging) and DevBench (li2024devbench) due to their demonstrated high human-LLM agreement and the validated reliability of state-of-the-art LLM judges on the evaluation tasks.
 For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of solutions to the given questions.
@@
 The evaluation results of 12 close-source and 3 open-source LLM judges in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2.
-For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks.
+For each judge, we calculate its average $RS$, PCPC, and $PF$ across all candidates and tasks.
 For $RS$ and $PC$, higher values are preferable.
@@
 Judge MTBench Pairwise DevBench Pairwise MTBench List-wise
-$RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error $PC$ $PF$ Error
+RSRS PCPC PFPF Error RSRS PCPC PFPF Error PCPC PFPF Error
 Claude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 0.67 $\pm$ 0.19 0.17 $\pm$ 0.19 0.00
@@
 Llama-3.1-8B 0.75 $\pm$ 0.32 0.69 $\pm$ 0.23 -0.03 0.25 0.79 $\pm$ 0.36 0.47 $\pm$ 0.18 0.25 0.00
-Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold.
+Table 2: Evaluation results for Repetition Stability (RSRS), Position Consistency (PCPC), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold.
 Errors arise from judgment failures (e.g., exceeding context window, not following output format).
-High error rates and low $RS$ are marked red, rendering further evaluations invalid due to insufficient data.
+High error rates and low RSRS are marked red, rendering further evaluations invalid due to insufficient data.
 List-wise evaluation is conducted on four representative judges to validate scalability.
@@
 Fig.
-(a)(b) are the radar charts for the $PC$ comparison by family, judge, and task.
+(a)(b) are the radar charts for the PCPC comparison by family, judge, and task.
 Fig.
-(c) leverages linear regression to explore the general relationship between $PC$ and $PF$.
+(c) leverages linear regression to explore the general relationship between PCPC and $PF$.
 Fig.
@@
 Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality.
-In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate ($owr$).
-Therefore, $\delta_{q}$ increases as $owr$ extends from baseline to 0 or 1.
+In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate (owrowr).
+Therefore, $\delta_{q}$ increases as owrowr extends from baseline to 0 or 1.
 Fig.
@@
 By stepwise regression, we discovered that average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table.
-4.
+5.
 In other words, there is a very weak relationship between the lengths of prompt components and position bias.
@@
 Appendix Table of Contents
-[…]
-[appendices]0
 Appendix A Reproducibility
@@
 B.1 LLM-as-a-Judge
-Large Language Models (LLMs) have become a transformative tool in automating evaluative tasks, offering scalability and reproducibility advantages over human assessments.
-The methodology of using LLMs as evaluators ("LLM-as-a-Judge") has been widely used for tasks such as open-ended story generation (Chiang and Lee, 2023a), adversarial attacks (Chiang and Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi and Federmann, 2023), and instruction following (Zeng et al., 2023), where models are tasked with scoring or ranking outputs.
-Despite their potential, inherent biases—particularly position bias—pose significant challenges to their reliability and fairness, even in the most effective pairwise comparative settings.
+Large Language Models (LLMs) have become a transformative tool in automating evaluative tasks, offering scalability and reproducibility advantages over human assessments (zheng2024judging).
+The methodology of using LLMs as evaluators ("LLM-as-a-Judge") has been widely used for tasks such as open-ended story generation (chiang-lee-2023-large), adversarial attacks (chiang2023can), summarization (karpinska2021perils), machine translation (kocmi-federmann-2023-large), and instruction following (zeng2023evaluating), where models are tasked with scoring or ranking outputs.
+LLM-as-a-Judge paradigms range from pointwise, pairwise, and list-wise ranking approaches (qin2024large; zhu2024starlingb) to score-based and relation-based (comparative or ranking) settings (li2024split).
+Furthermore, explorations into multi-agent frameworks (li2023prd; chan2023chatevalbetterllmbasedevaluators) and meta-thinking or meta-judge systems (wu2024metarewardinglanguagemodelsselfimproving; li2025leveragingllmsmetajudgesmultiagent) have emerged to enhance reliability of LLM evaluations.
+Despite their potential, inherent biases—particularly position bias—pose significant challenges to their reliability and fairness (zheng2024judging; ye2024justiceprejudicequantifyingbiases; ma2025judging), even in the most effective pairwise comparative settings.
 B.2 Position Bias
-However, position bias is a complex problem that is challenging to deal with or solve.
-The naive way is to exclude the position inconsistent judgments, which does not solve the fundamental issue and would likely result in data sparsity when position bias is frequently exhibited.
-Therefore, researchers proposed ‘inconsistency-as-a-tie’ for both models in pairwise comparative settings to consider all judgments for further analysis (Zheng et al., 2024b; Li et al., 2023c).
+Position bias, the tendency of LLMs to favor certain positions within the prompt irrespective of actual content, is a pervasive issue observed across various domains and tasks (zheng2024judging; qin2024large; li2024devbench; li2024split).
+In the context of LLM-as-a-Judge, it typically refers to the position inconsistent scenarios where a LLM judge prefers different candidate solutions to a given question when the order of solutions in the prompt is permuted.
+To deal with such bias, the naive way is to exclude the position inconsistent judgments, which does not solve the fundamental issue and would likely result in data sparsity when position bias is frequently exhibited.
+Then, researchers proposed ‘inconsistency-as-a-tie’ for both candidate models in pairwise comparative settings to consider all judgments for further analysis (zheng2024judging; li2023prd).
 This approach, while practically useful for evaluations, does not mitigate position bias.
-Given the significance of position bias, researchers have developed more sophisticated approaches to mitigate it, including bootstrapping (Hou et al., 2024), split-and-merge techniques (Li et al., 2023d), and multi-agent discussions (Li et al., 2023c; Khan et al., 2024).
-While these methods demonstrate potential, they are often costly, time-consuming, or insufficient.
-For example, multi-agent discussions and iterative reviews require extensive computational resources and human-like deliberation, while simpler techniques, such as order permutation (Wang et al., 2023; Zheng et al., 2024a) and batch calibration (Zhou et al., 2024), have limited effectiveness in addressing bias holistically.
-Furthermore, calibration methods like ensembling (Li et al., 2023a) or context decomposition (Li et al., 2023d) are often constrained by the inadequacy of supervised data to align LLM evaluators effectively (Liu et al., 2024).
-This highlights that addressing this crucial bias is complex and challenging.
-The limitations of existing approaches underscore a critical gap in our understanding of position bias itself.
-A lack of clarity on the factors influencing this bias and their quantitative impacts hampers the development of effective solutions.
-For instance, while the PORTIA approach (Li et al., 2023d) demonstrates significant improvements in consistency for GPT-4 (from 93.44% to 97.03%), this success may be overestimated due to the selection of tasks with large answer quality gaps.
-As our study reveals, such gaps inherently reduce the likelihood of position bias, making these cases easier to calibrate.
-This example underscores the necessity of first comprehensively understanding the underlying factors contributing to position bias before proposing or refining mitigation strategies.
-Without this foundational understanding, the efficacy of current and future methods remains uncertain and potentially misjudged.
-B.3 Research Gaps
-To enhance the understanding of position bias, our study builds on and extends prior work through the following dimensions.
-Repetition Stability: Previous studies on LLM-as-a-Judge often implicitly assume that single-instance judgments are sufficient to reflect the evaluating capabilities of judge models, thereby attributing observed biases entirely to systematic tendencies rather than random variations.
-These studies typically proceed directly to analyses on position bias or others without such verification.
-However, if judgments vary by chance or lack consistency across repetitions, the reliability and accuracy of downstream analyses are compromised.
-To avoid this issue, Chen et al.
-(2024) repeated evaluations multiple times and adopted the mode judgment as the representative, an approach that, while effective, is computationally expensive.
-Addressing this gap, our work introduces a method or metric for systematically evaluating the stability of LLM judgments across repetitions.
-This ensures that observed biases are systematic rather than artifacts of variability.
-By performing this validation on a sample dataset for cost-efficiency, researchers can establish a robust foundation for subsequent evaluations and analyses, ensuring the reliability of their conclusions.
-Preference Fairness: Prior studies on position bias often focus on the proportion of position consistent judgments, referred to as “position consistency”.
-Only a few have explored the preference directions of LLM evaluators in addition to consistency.
-However, traditional measurements either overly penalize highly consistent judges (Li et al., 2023c; Liusie et al., 2024) or rely on metrics whose sensitivity is highly dependent on dataset size (Zheng et al., 2024b; Zhu et al., 2023).
-To address these limitations, we proposed a refined metric, preference fairness ($PF$), which integrates the strengths of both approaches.
-This single, comprehensive metric ensures sensitivity across datasets, remains independent of the number of questions and tasks, and represents a significant improvement over previous metrics.
-Answer Quality Gap: Intuitively, candidate solutions with significant quality disparities are easier to judge, likely resulting in reduced position bias.
-This suggests that the quality gap among answers is a potentially significant factor influencing position bias.
-However, this factor has been largely overlooked in prior works, both in evaluations and in the development of mitigation strategies, underscoring our limited understanding of position bias.
-While Wang et al.
-(2023) explored the answer quality gap in score-based LLM-as-a-Judge evaluations using judgment scores, quantifying this variable in comparative settings (pairwise or list-wise) remains an open research problem.
-To address this gap, we propose the use of win rates (Zheng et al., 2024b) of candidate solutions against an expected baseline to measure the answer quality gap.
-This approach offers a practical quantitative metric for subsequent analysis and deepens our understanding of the factors driving position bias.
-List-wise Evaluations: Although LLM-as-a-Judge has been categorized into pointwise, pairwise, and list-wise paradigms (Qin et al., 2024), most research has focused on pairwise comparative assessments due to their demonstrated effectiveness (Zheng et al., 2024b; Liusie et al., 2024).
-However, this leaves a significant gap in measuring and understanding the position bias in list-wise settings, where LLMs are prompted to evaluate three or more candidate solutions simultaneously.
-With advancements in model capabilities, list-wise evaluations present a potentially more cost-effective alternative to pairwise approaches, requiring fewer comparisons while maintaining consistency and fairness.
-Our work bridges this gap by extending position bias assessments from pairwise to list-wise settings, enabling a more comprehensive evaluation of LLM-as-a-Judge across varied configurations.
+Given the significance of position bias, researchers have developed debiasing strategies, including bootstrapping (hou2024large), split-and-merge techniques (li2023split), and multi-agent discussions (li2023prd; khan2024debating).
+While these methods demonstrate potential, they are often costly, time-consuming, or insufficient to eliminate position bias.
+More recently, mechanistic approaches have been explored to address position bias by modifying model internals like positional embeddings and causal attention mechanisms (yu2025mitigate; wang2025eliminating).
+These methods, while effective in eliminating position bias, lack applicability to state-of-the-art closed-source models, which are frequently used in practice but model internals are inaccessible.
+Therefore, position bias remains a complex and challenging issue to fully resolve.
+B.3 Evaluating Position Bias Evaluation
+Parallel to developing debiasing strategies, a crucial line of research focuses on comprehensively assessing position bias and identifying its influencing factors, thereby fostering a deeper understanding of this phenomenon.
+The evaluation of position bias typically concentrates on two main aspects: the consistency of judgments when the order of candidate solutions in the prompt is changed, and the preference direction (e.g., primacy or recency bias) when inconsistencies arise.
+Metrics for these aspects in pairwise comparisons are relatively established; however, their extension to, and rigorous measurement in, list-wise settings remain less explored.
+While prior works (qin2024large; zhu2024starlingb) introduced list-wise ranking settings, they often become computationally expensive due to exponentially increasing number of pairs in the list or encounter significant degradation of performance due to ranking complexity.
+Moreover, the traditional preference direction measurement of position bias becomes increasingly complicated in list-wise paradigms (i.e., preferring the $i^{th}$ candidate solution for all i’s).
+However, practically, “choosing the best out of an n-item list” is a simpler but common need compared to exhaustive ranking of the entire list, which is not yet fully explored by prior works.
+A unified single metric for effectively measuring preference direction is therefore necessary to facilitate position bias evaluation in a more comprehensive manner.
+B.4 Factors influencing Position Bias
+Position bias may be influenced by several factors, including judge attributes, task types, candidate solution lengths and qualities, and more.
+For example, wang2023large discovered that judgment conflict rate (i.e., frequency where position bias exhibits across the dataset) is negatively correlated with the score gap between the two candidate responses.
+However, quantifying and understanding the impact of this “answer quality gap" in comparative settings (both pairwise and list-wise) is an area that requires further investigation.
+Besides, prior works often observe position bias and conduct subsequent analyses based on single-instances judgments.
+However, observed biases might stem from random variations rather than systematic tendencies, hence compromising the reliability of downstream bias analyses.
+To mitigate this, researchers have employed multiple repetitions and used the mode judgment, an approach that, while robust, can be computationally expensive (chen2024mllmasajudge).
+In this sense, validating that position bias emerges from positional information rather than randomness becomes a crucial yet insufficiently-explored area.
 These identified gaps - spanning from the validation of judgments to the evaluation of list-wise settings - highlight the need for a more comprehensive framework to evaluate and understand position bias.
@@
 Appendix D More Results of Position Bias and Answer Quality Gap Measurement
+Table 3 presents additional evaluation results for more open-source models, reinforcing our earlier observations.
+The scores consistently show that position bias varies across judges and tasks, while high repetition stability confirms that such biases are systematic rather than random.
+These extended results further validate our evaluation findings.
+Model MTBench Results DevBench Results
+PCPC PFPF RSRS Error PCPC PFPF RSRS Error
+GPT-OSS-20B 0.79 $\pm$ 0.16 -0.04 $\pm$ 0.13 0.89 $\pm$ 0.17 0.06 0.84 $\pm$ 0.16 -0.04 $\pm$ 0.14 0.81 $\pm$ 0.25 0.08
+Gemma-3n-E4B 0.68 $\pm$ 0.17 -0.07 $\pm$ 0.18 0.90 $\pm$ 0.17 0.00 0.41 $\pm$ 0.24 0.40 $\pm$ 0.35 0.93 $\pm$ 0.12 0.08
+Llama-4-Scout-17B 0.79 $\pm$ 0.16 0.05 $\pm$ 0.15 0.95 $\pm$ 0.08 0.00 0.12 $\pm$ 0.18 0.88 $\pm$ 0.18 0.98 $\pm$ 0.03 0.00
+Mistral-Small-24B 0.72 $\pm$ 0.19 -0.02 $\pm$ 0.17 0.87 $\pm$ 0.20 0.05 0.57 $\pm$ 0.18 0.27 $\pm$ 0.19 0.82 $\pm$ 0.31 0.02
+Qwen2.5-72B 0.77 $\pm$ 0.17 -0.01 $\pm$ 0.16 0.94 $\pm$ 0.09 0.00 0.81 $\pm$ 0.15 0.11 $\pm$ 0.16 0.93 $\pm$ 0.13 0.00
+Qwen2.5-Coder-32B 0.74 $\pm$ 0.18 0.01 $\pm$ 0.17 0.91 $\pm$ 0.14 0.00 0.81 $\pm$ 0.16 0.09 $\pm$ 0.17 0.92 $\pm$ 0.13 0.03
+Qwen2.5-7B 0.55 $\pm$ 0.18 -0.21 $\pm$ 0.23 0.87 $\pm$ 0.22 0.00 0.54 $\pm$ 0.21 -0.45 $\pm$ 0.21 0.93 $\pm$ 0.12 0.00
+Table 3: More evaluation results for Position Consistency (PCPC), Preference Fairness (PFPF), and Repetition Stability ($RS$) across MTBench and DevBench datasets.
 D.1 MTBench
@@
 An asterisk marks the statistical significance by Student’s t-tests.
-Figure (a) selects GPT-4 as the baseline, where asterisk demonstrates signficantly better or worse $PC$ of other models compared to it.
-Figure (b) utilizes an absolute $PF$ baseline of 0 and depicts preference fairness performances of LLM judges across tasks.
+Figure (a) selects GPT-4 as the baseline, where asterisk demonstrates signficantly better or worse PCPC of other models compared to it.
+Figure (b) utilizes an absolute PFPF baseline of 0 and depicts preference fairness performances of LLM judges across tasks.
 Similar to findings on MTBench, position bias is significantly judge-dependent and task-dependent on DevBench as well.
@@
 Specifically, the variables include lengths (input, output, and prompt), answer quality gap, LLM judges, candidate models, and task categories to predict position consistency and preference fairness respectively.
-Table 5, 6 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively.
-Table 7, 8 serves for DevBench, and Table 3, 4 is conducted on the integrated set of both benchmarks.
+Table 6, 7 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively.
+Table 8, 9 serves for DevBench, and Table 4, 5 is conducted on the integrated set of both benchmarks.
 The impact of variables on the model is ranked from highest to lowest, from bottom to top.
@@
   • DevBench (10): codellama-7b-instruct, codellama-13b-instruct, codellama-34b-instruct, deepseek-coder-1.3b-instruct, deepseek-coder-6.7b-instruct, deepseek-coder-33b-instruct, gpt-3.5-turbo-1106, gpt-4-0125-preview, gpt-4-0613, gpt-4-1106-preview
-The model names are exactly what MTBench Zheng et al.
-(2024b) and DevBench Li et al.
-(2024a) used in their studies.
+The model names are exactly what MTBench zheng2024judging and DevBench li2024devbench used in their studies.
 That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4).
@@
   • A specified output format for later judgment extraction
-  • Chain-of-Thought Wei et al. (2022) prompts requiring the LLM judge to provide reasons for its judgment
+  • Chain-of-Thought wei2022chain prompts requiring the LLM judge to provide reasons for its judgment
 The detailed prompt templates are specified below.